### v2

# Triple DES AES Advanced Encryption Standard

3DES
Triple DES
AES
Advanced Encryption Standard
AI
Artificial Intelligence
ANF
Algebraic Normal Form
ALU
Arithmetic Logic Unit
ANN
Artificial Neural Network
API
Application Programming Interface
ABI
Application Binary Interface
APIC
Advanced Programmable Interrupt Controller
ARX
Addition, Rotate, XOR
ASLR
Address Space Layout Randomization
ASK
Amplitude-Shift Keying
ASIC
Application Specific Integrated Circuit
ASID
Address Space Identifier
BL
Bootloader Enable
BOR
Brown-Out Reset
BPSK
Binary Phase Shift Keying
BTB
Branch Target Buffer
CBC
Cipher Block Chaining
CBS
Critical Bootloader Section
CGM
Continuous Glucose Monitoring System
CMOS
Complementary Metal Oxide Semiconductor
CNN
Convolutional Neural Network
COPACOBANA
Cost-Optimized Parallel Code Breaker and Analyzer
CPA
Correlation Power Analysis
CPU
Central Processing Unit
CTR
Counter (mode of operation)
CRC
Cyclic Redundancy Check
CRP
Code Readout Protection
CVE
Common Vulnerabilities and Exposures
DES
Data Encryption Standard
DC
Direct Current
DDR
Double Data Rate RAM
DDS
Digital Direct Synthesis
DFT
Discrete Fourier Transform
DMA
Direct Memory Access
DNN
Deep Neural Network
DoS
Denial-of-Service
DFA
Differential Fault Analysis
DPA
Differential Power Analysis
DRAM
Dynamic Random-Access Memory
DRM
Digital Rights Management
DSO
Digital Storage Oscilloscope
DSP
Digital Signal Processing
DST
Digital Signature Transponder
DUT
Device Under Test
DVFS
Dynamic Voltage and Frequency Scaling
ECB
Electronic Code Book
ECC
Elliptic Curve Cryptography
ECU
Electronic Control Unit
EDE
Encrypt-Decrypt-Encrypt (mode of operation)
EEPROM
Electrically Erasable Programmable Read-Only Memory
EL
Exception Level
EM
Electro-Magnetic
EPID
Enhanced Privacy ID
FFT
Fast Fourier Transform
FCC
Federal Communications Commission
FIR
Finite Impulse Response
FIVR
Fully Integrated Voltage Regulator
FPGA
Field Programmable Gate Array
FSK
Frequency Shift Keying
GMSK
Gaussian Minimum Shift Keying
GPC
Granule Protection Check
GPT
Granule Protection Table
GPIO
General Purpose I/O
GPU
Graphics Processing Unit
GPGPU
General Purpose GPU
GVT
Graphics Virtualization Technology
HD
Hamming Distance
HDL
Hardware Description Language
HE
Homomorphic Encryption
HF
High Frequency
HMAC
Hash-based Message Authentication Code
HW
Hamming Weight
MPS
Multi-Process Service
IC
Integrated Circuit
ID
Identifier
ISM
Industrial, Scientific, and Medical (frequencies)
IIR
Infinite Impulse Response
IP
Intellectual Property
IoT
Internet of Things
ISA
Instruction Set Architecture
IV
Initialization Vector
JTAG
Joint Test Action Group
LLC
Last-Level Cache
LF
Low Frequency
LFSR
Linear Feedback Shift Register
LQI
Link Quality Indicator
LSB
Least Significant Bit
LSByte
Least Significant Byte
LUT
Look-Up Table
MAC
Message Authentication Code
MKTME
Multi-Key Total Memory Encryption
MDS
Microarchitectural Data Sampling
MEE
Memory Encryption Engine
MF
Medium Frequency
ML
Machine Learning
MMIO
Memory-Mapped I/O
MMU
Memory Management Unit
MITM
Man-In-The-Middle
MPE
Memory Protection Engine
MIG
Multi-Instance GPU
MPC
Multi-Party Computation
MSB
Most Significant Bit
MSByte
Most Significant Byte
MSK
Minimum Shift Keying
MSR
Model Specific Register
[…]
Microcontroller
NLFSR
Non-Linear Feedback Shift Register
NLF
Non-Linear Function
NFC
Near Field Communication
NN
Neural Network
NPU
Neural Processing Unit
NRZ
Non-Return-to-Zero (encoding)
NVM
Non-Volatile Memory
NW
Normal World
SW
Secure World
OOK
On-Off-Keying
OP
Operational Amplifier
OTP
One-Time Password
ORAM
Oblivious RAM
PC
Personal Computer
PCB
Printed Circuit Board
PCIe
Peripheral Component Interconnect Express
PHT
Page History Table
PhD
Patiently hoping for a Degree
PKE
Passive Keyless Entry
PKES
Passive Keyless Entry and Start
PKI
Public Key Infrastructure
PLL
Phase Locked Loop
PMBus
Power Management Bus
PMIC
Power Management Integrated Circuit
PoC
Proof-of-Concept
POR
Power-On Reset
PPC
Pulse Pause Coding
PRNG
Pseudo-Random Number Generator
PSK
Phase Shift Keying
PWM
Pulse Width Modulation
RAPL
Running Average Power Limit
RDP
Read-out Protection
REE
Rich Execution Environment
RMM
Realm Management Monitor
RMP
Reverse Map Table
RF
Radio Frequency
RFID
Radio Frequency IDentification
RKE
Remote Keyless Entry
RNG
Random Number Generator
ROM
Read Only Memory
ROP
Return-Oriented Programming
RSA
Rivest Shamir and Adleman
RSB
Return Stack Buffer
RTL
Register Transfer Level
SCA
Side-Channel Analysis
SDK
Software Development Kit
SDR
Software-Defined Radio
SGX
Software Guard Extensions
SEV
Secure Encrypted Virtualization
SNR
Signal to Noise Ratio
SNP
Secure Nested Paging
SHA
Secure Hash Algorithm
SHA-1
Secure Hash Algorithm 1
SHA-256
Secure Hash Algorithm 2 (256-bit version)
SMA
SubMiniature version A (connector)
SMBus
System Management Bus
I2C
Inter-Integrated Circuit
SPA
Simple Power Analysis
SPI
Serial Peripheral Interface
SPOF
Single Point of Failure
SoC
System on Chip
SVID
Serial Voltage Identification
SWD
Serial Wire Debug
TCB
Trusted Computing Base
TEE
Trusted Execution Environment
TLB
Translation Lookaside Buffer
TMTO
Time-Memory Tradeoff
TMDTO
Time-Memory-Data Tradeoff
TA
Trusted Application
TrustZone
TrustZone
RSSI
Received Signal Strength Indicator
SHF
Superhigh Frequency
SMC
Secure Monitor Call
UART
Universal Asynchronous Receiver Transmitter
[…]
Microcode
UHF
Ultra High Frequency
UID
Unique Identifier
USRP
Universal Software Radio Peripheral
USRP2
Universal Software Radio Peripheral (version 2)
USB
Universal Serial Bus
VHF
Very High Frequency
VLF
Very Low Frequency
VHDL
VHSIC (Very High Speed Integrated Circuit) Hardware Description Language
VR
Voltage Regulator
WLAN
Wireless Local Area Network
XEX
Xor-Encrypt-Xor
XTS
XEX-based Tweaked-Codebook Mode with Ciphertext Stealing
XOR
Exclusive OR
IR
Intermediate Representation
OCD
On-Chip Debug
OS
Operating System
TSX
Transactional Synchronization Extensions
TZASC
TrustZone Address Space Controller
VM
Virtual Machine
vGPU
Virtual Graphics Processing Unit
CC
Confidential Computing
CCA
Confidential Compute Architecture
TDX
Trust Domain Extensions
Confidential Computing on Heterogeneous CPU-GPU Systems: Survey and Future Directions
Qifan Wang University of BirminghamBirminghamUK q.wang.3@bham.ac.uk  and  David Oswald University of BirminghamBirminghamUK d.f.oswald@bham.ac.uk
Abstract.
In recent years, the widespread informatization and rapid data explosion have increased the demand for high-performance heterogeneous systems that integrate multiple computing cores such as Central Processing Units, Graphics Processing Units, Application Specific Integrated Circuits, and Field Programmable Gate Arrays. The combination of CPU and GPU is particularly popular due to its versatility. However, these heterogeneous systems face significant security and privacy risks. Advances in privacy-preserving techniques, especially hardware-based Trusted Execution Environments, offer effective protection for GPU applications. Nonetheless, the potential security risks involved in extending TEEs to GPUs in heterogeneous systems remain uncertain and need further investigation. To investigate these risks in depth, we study the existing popular GPU TEE designs and summarize and compare their key implications. Additionally, we review existing powerful attacks on GPUs and traditional TEEs deployed on CPUs, along with the efforts to mitigate these threats. We identify potential attack surfaces introduced by GPU TEEs and provide insights into key considerations for designing secure GPU TEEs. This survey is timely as new TEEs for heterogeneous systems, particularly GPUs, are being developed, highlighting the need to understand potential security threats and build both efficient and secure systems.
Confidential computing, Trusted Execution Environment on GPU, hardware security, GPU attacks
ccs: General and reference Surveys and overviewsccs: Computer systems organization Heterogeneous (hybrid) systemsccs: Security and privacy Hardware attacks and countermeasuresccs: Security and privacy Trusted computing
1. Introduction
[…]
To ensure the security and confidentiality of heterogeneous computing systems, researchers have focused on adapting privacy-preserving techniques traditionally used on CPUs to accelerators like GPUs. These techniques include cryptographic methods such as Homomorphic Encryption (HE(Mishra et al., 2020; Ng and Chow, 2021; Fan et al., 2023) and Secure Multi-Party Computation (MPC(Mishra et al., 2020; Ng and Chow, 2021; Tan et al., 2021; Watson et al., 2022; Jawalkar et al., 2023), as well as TEEs tailored for accelerators like GPUs (Volos et al., 2018; Jang et al., 2019; Zhu et al., 2020; NVIDIA, 2022a; Jiang et al., 2022; Lee et al., 2022; Yudha et al., 2022; Deng et al., 2022; Ivanov et al., 2023; Mai et al., 2023; Wang et al., 2024b). TEEs offer certain advantages over cryptographic methods, particularly in performance and ease of deployment, providing a robust hardware-based security mechanism that is relatively straightforward to implement and manage. However, in contrast to cryptographic techniques, the security of TEEs relies on trust in the manufacturer and is difficult to analyse and prove rigorously.
Recent research has explored extending TEEs to accelerators like GPUs, both for x86 (Volos et al., 2018; Jang et al., 2019; Zhu et al., 2020; NVIDIA, 2022a; Lee et al., 2022; Yudha et al., 2022; Ivanov et al., 2023; Mai et al., 2023) and Arm (Jiang et al., 2022; Deng et al., 2022; Wang et al., 2024b) architectures. Most of these approaches involve modifying existing CPU TEEs (especially process-based TEEs like Intel Software Guard Extensions (SGX) and Arm TrustZone (TrustZone)) to secure the GPU software stack, such as GPU drivers, and proposing strategies to ensure the isolation and confidentiality of GPU kernel executions. Apart from this, NVIDIA has introduced a pioneering commercial platform in its Hopper architecture (NVIDIA, 2022a), which supports confidential computing through features like memory encryption and isolated execution contexts. However, the security threats associated with extending TEEs across both CPUs and accelerators in heterogeneous computing systems remain uncertain and require further investigation.
Figure 1. Current state of security research on heterogeneous computing systems with GPU. GPU Crypto refers to papers that protect the privacy of GPU-based applications using cryptographic techniques such as HE and MPC. AA, MSCA, MCCA, PSCA, SFIA refer to architectural, microarchitectural side-channel/covert-channel, physical side-channel, and software-based fault injection attacks.
1999200420092014201920242029 First GPU, NVIDIA GeForce256 ARM TrustZone Launched in ARM11 First Virtualized GPU (NVIDIA VGX) First Attack Exploiting GPU Vuln. (Lee et al., 2014; Maurice et al., 2014) First MSCA on GPU (Luo et al., 2015) Intel SGX Launched in Skylake First MCCA on GPU (Naghibijouybari et al., 2017) AMD SEV Launched in EPYC First FIA on GPU (Frigo et al., 2018) First x86-based GPU TEE (Volos et al., 2018) First MIG-enabled GPU, NVIDIA A100 Intel TDX Announced AMD SEV-SNP Available in Milan First Arm-based GPU TEE (Deng et al., 2022) First Commercial GPU TEE (NVIDIA, 2022a) First MCCA on MIG-enabled GPU (Zhang et al., 2023) First Arm CCA-based GPU TEE (Wang et al., 2024b)
Figure 2. Timeline of significant events in secure heterogeneous computing systems.
In this survey, we delve into the potential vulnerabilities in current TEE designs on accelerators, with a specific emphasis on GPUs (referred to as GPU TEEs). To this end, we begin by reviewing and classifying research efforts dedicated to creating GPU TEE prototypes. These solutions target various aspects: some aim to secure the GPU software stack, others focus on safeguarding the confidentiality of data processed on the GPU, and some provide robust isolation mechanisms. Consequently, each design presents different attack surfaces. To identify these potential attack surfaces, we categorize and analyze existing attacks on GPUs, including architectural/logical attacks, microarchitectural side-channel and covert-channel attacks, physical side-channel attacks, and fault injection. The unique architecture and accessibility of GPUs within heterogeneous computing systems pose significant differences compared to traditional CPUs, which have historically been the primary targets of these attacks. Additionally, we also review existing attacks against CPU TEEs such as Intel SGX, Arm TrustZone, and AMD Secure Encrypted Virtualization (SEV), and investigate their implications for code and data running on GPU TEEs. Finally, by synthesizing these insights, we discuss the potential attack surfaces in GPU TEEs exposed by these powerful attack vectors.
As shown in Figure 1 and Figure 2, the evolution of GPUs has led to the development of increasingly high-performance architectures. Consequently, research has increasingly focused on investigating potential attack surfaces in GPUs by exploring new architectural vulnerabilities. Likewise, we expect the potential security threats of recently emerging GPU TEEs to receive growing attention. We believe that this survey provides critical insights into key considerations for designing GPU TEEs, the primary workflows of existing attacks targeting heterogeneous computing systems, and the potential attack surface introduced for both CPU TEEs and GPUs. This understanding can guide future research into developing defenses for secure high-performance computing on heterogeneous systems and enhance confidential computing on GPU TEEs by refining the Trusted Computing Base (TCB). Moreover, we believe our work can inspire the creation of more GPU TEE prototypes that are well-suited for various architectures and real-world applications.
2. Background
In this section, we introduce the basics of GPUs, common attack types, and the most existing popular TEE designs.
2.1. General Purpose Graphics Processing Units
Table 1. Summary of existing popular GPUs
Accelerator Year Features Virtualization Memory Sharing Memory Size
None L1/L2 Cache DRAM L1 L2 DRAM
Intel HD/Iris/UHD/Xe 2010-2020 Integrated graphics GVT $\checkmark$ $\checkmark$ Shared Shared Shared
NVIDIA P100 2016 HBM2 vGPU/MPS $\checkmark$ 64 KB 4MB 16 GB
V100 2017 Optimized SM for ML vGPU/MPS $\checkmark$ 128 KB 6MB 16 GB
A100 2020 3G Tensor Cores and NVLink vGPU/MPS/MIG $\checkmark$ 192 KB 40 MB 40 GB
H100 2022 Confidential computing vGPU/MPS/MIG $\checkmark$ 256 KB 50 MB 80 GB
AMD Radeon RX 580 2017 FreeSync, CrossFire N/A $\checkmark$ 16 KB 2 MB 4/8 GB
Radeon RX 6900 2020 Infinity Cache, SAM N/A $\checkmark$ 80 KB 4 MB 16 GB
Arm Mali-T624 GPU 2013 Unified Shader Architecture N/A $\checkmark$ 16 KB 256 KB Shared
Qualcomm Adreno GPU 800 2020 AI Acceleration N/A $\checkmark$ 32-64 KB 256 KB-1 MB Shared
In Virtualization, vGPU, GVT, MPS, and MIG denote Virtual Graphics Processing Unit, Graphics Virtualization Technology, Multi-Process Service, and Multi-Instance GPU. Memory Sharing denotes whether a level of the memory hierarchy is shared between the CPU and the accelerator.
General Purpose GPUs are specialized accelerators designed to handle diverse compute-intensive workloads beyond traditional graphics rendering. With their robust parallel computing capabilities, these GPUs are ideal for applications such as data analysis, ML, cryptography, etc. Typically, a GPGPU consists of multiple processing units (i.e., CUDA cores in NVIDIA GPUs and shader cores and execution units in Arm GPUs) capable of executing thousands of threads simultaneously. Aside from GPUs, there are several prominent accelerators designed to boost performance for specific computational tasks. These include FPGAs, ASICs, and NPUs. In this survey, our primary focus is on heterogeneous systems utilizing GPUs.
With the evolution of GPUs, vendors like Intel, NVIDIA, AMD, Arm, Qualcomm, and others have introduced GPUs with distinct architectural features and capabilities. Different architectural features, such as the memory hierarchy, may introduce different attack surfaces on GPUs. Existing widely used GPUs, detailed in Table 1, vary in their architectural designs, particularly concerning their memory hierarchy, which can affect their vulnerability to certain types of attacks (e.g., microarchitectural attacks on integrated and discrete GPUs in Section 5). To provide a comprehensive overview of these GPUs, we consider the aspects in the following sections.
Integrated/Discrete GPU. Integrated GPUs, such as Intel HD/Iris/UHD/Xe Graphics, Arm, and Qualcomm GPUs listed in Table 1, are integrated on the same die or package as the CPU. These GPUs share the system’s memory with the CPU. Discrete GPUs are standalone and dedicated graphics processing units separate from the CPU. They have their dedicated memory, which is distinct from the system memory (e.g., NVIDIA and AMD GPUs listed in Table 1).
GPU virtualization. Graphics Virtualization Technology (GVT), a.k.a. Intel GVT, enables a single physical GPU to be shared among multiple Virtual Machines with dedicated GPU resources. Multi-Process Service (MPS) from NVIDIA manages multiple CUDA applications concurrently on a single GPU but lacks the address space isolation of true GPU virtualization, making it unsuitable for multi-tenant systems (Zhang et al., 2023). NVIDIA’s Virtual Graphics Processing Unit (vGPU), unlike MPS, ensures separate GPU address spaces and execution contexts. Multi-Instance GPU (MIG), introduced in NVIDIA’s A100 GPU architecture, partitions a single physical GPU into multiple instances. Each MIG instance operates as a separate GPU with dedicated resources, including compute cores, memory, and cache.
Open-source status of GPU source code. The availability of GPU source code, including drivers and low-level architectural details, varies depending on the vendor’s policies and the specific GPU architecture. For instance, vendors like AMD (e.g., AMD’s ROCm (Hunt et al., 2020)) and Intel (Mesa 3D Graphics (Intel, 2024)) offer open-source GPU drivers for specific platforms. However, NVIDIA, ARM, and Qualcomm typically provide proprietary drivers, limiting developers’ access to design details. As a result, many research efforts on GPU attacks often rely on reverse engineering to acquire details like Translation Lookaside Buffer (TLB) levels, warp schedulers, and more (see Section 5).
2.2. Types of GPU Attacks
In recent decades, a range of attacks has emerged as significant threats to modern CPU-based computing systems. As the use of GPUs becomes more prevalent, there is growing interest in exploring potential security vulnerabilities that extend from the CPU to the GPU. In this survey, we detail various attack vectors, including architectural/logical attacks, microarchitectural side-channel and covert-channel attacks, physical side-channel attacks, and fault injection.
2.2.1. Architectural/Logical Attacks
These attacks encompass a class of attacks that exploit vulnerabilities or weaknesses in the underlying hardware architecture. These attacks aim to compromise system by targeting the physical or logical design of key hardware elements, including processing units (e.g., processors, hardware accelerators, and co-processors), the memory subsystem, and communication components responsible for interfacing with other system elements (Ghasempouri et al., 2023).
2.2.2. Microarchitectural Side-Channel and Covert-Channel Attacks
Microarchitectural side-channel attacks exploit unintended behaviours of a processor’s microarchitecture to extract sensitive information. They target components like cache behaviour, branch prediction, or speculative execution. Initially demonstrated on CPUs, these attacks exploit various optimization structures such as caches (Gruss et al., 2016; Liu et al., 2015; Gruss et al., 2015; Osvik et al., 2006) and branch predictors (Evtyushkin et al., 2016a, b). Recently, a new category of microarchitectural attacks, known as transient-execution attacks (Xiong and Szefer, 2021), has emerged on CPUs. Examples include Meltdown (Lipp et al., 2020), Spectre (Kocher et al., 2020), and Microarchitectural Data Sampling (MDS(Van Schaik et al., 2019; Ragab et al., 2021), which rely on out-of-order and transient (speculative) execution in modern CPUs. Unlike physical side-channel attacks, microarchitectural attacks necessitate an understanding of the specific microarchitecture and its inherent vulnerabilities.
A microarchitectural covert-channel attack is a subtype of a side-channel attack. Covert channels involve two processes establishing an unauthorized communication pathway through an unintended channel to leak information, making them relevant in cloud computing environments (Muchene et al., 2013; Zhenyu et al., 2012). Due to imperfections in isolation and shared resource usage, malicious processes can exploit such resources to establish a communication channel with the victim process. The covert channel typically involves a sender, known as the “trojan”, and a receiver, often referred to as the “spy”. Through covert channels, the trojan can transmit sensitive information or leak data to the spy (Ahn et al., 2021b).
2.2.3. Physical Side-Channel Attacks
These passive attacks exploit the physical characteristics of a computing device, such as power consumption, Electro-Magnetic (EM), or timing information, to gather insights into the device’s operations. Examples include attacks such as Differential Power Analysis (DPA), which exploits fluctuations in power consumption to extract cryptographic keys or other sensitive data. Unlike microarchitectural attacks, physical side-channel attacks do not always demand knowledge of the device’s internal architecture or vulnerabilities. Instead, they use the unintentional leakage of information through physical signals and behaviour.
2.2.4. Fault Injection Attack
Fault injection attacks involve actively introducing faults or errors into a system to alter its behavior or extract sensitive information. These attacks can be generally categorized into non-invasive, semi-invasive, and invasive (Shuvo et al., 2023). Non-invasive attacks allow adversaries to use methods such as clock and voltage glitching, overheating, EM faults, software faults, and remote hardware faults, with either physical or non-physical access. Semi-invasive attacks require the (partial) decapsulation of the target chip for techniques like laser and optical fault injection, while invasive attacks involve more direct methods, such as micro-probing and Integrated Circuit (IC) modification, both of which necessitate deep physical access. In this work, we primarily focus on non-invasive fault injection attacks.
2.3. TEE
Most existing GPU TEEs rely on a CPU TEE to protect the GPU drivers on the CPU host, which can be compromised by the adversary. This section introduces CPU TEEs, including process-based TEEs such as Intel SGX and Arm TrustZone, which provide isolation at the process level, and virtualization-based TEEs like AMD SEV, Arm Confidential Compute Architecture (CCA), and Intel Trust Domain Extensions (TDX), which offer isolation at the VM level.
2.3.1. Intel SGX Overview
SGX offers a secure and isolated environment, termed enclave, for sensitive workloads without the need to rely on the security of the untrusted host Operating System (OS). SGX ensures the authenticity of executed code, maintains the integrity of runtime states (e.g., CPU registers, memory, and sensitive I/O), and preserves the confidentiality of enclave code, data, and runtime states stored in persistent memory through a sealing process (Costan and Devadas, 2016).
Threat model of Intel SGX. An attacker is assumed to control the entire software stack of the underlying computing platform, including the hypervisor, BIOS, and OS. Specifically, within this context, the attacker can (1) execute arbitrary instructions at privileged levels; (2) launch, pause, and destroy enclave instances as desired; (3) control the software stack within the non-enclave environment, including tasks such as memory mapping, I/O device, and CPU scheduling (Li et al., 2024).
TCB of Intel SGX. The TCB of SGX refers to all of a system’s hardware, firmware, and software components considered trusted and involved in providing a secure environment. A vulnerability in any TCB component would jeopardize the security of SGX. TCB size serves as a common metric to assess the security of a TEE design, where a smaller TCB size typically indicates fewer lines of code or hardware elements, implying a reduced vulnerability risk. However, comparing the security of two TEE designs based solely on TCB size is difficult and unreasonable (Li et al., 2024). Ongoing research efforts are focused on building smaller TEE OS with memory-safe programming languages (e.g., Rust-SGX (Wang et al., 2019)).
2.3.2. Arm TrustZone Overview
TrustZone has been integrated into Arm processors (Cortex-A) since 2004 (Arm, 2004), primarily employed in mobile devices. Unlike Intel SGX, TrustZone is not inherently tailored for cloud platforms and lacks support for remote attestation, which would enable remote parties to verify the initial states of the TEE instances (Li et al., 2024). TrustZone operates around the notion of protection domains referred to as the Secure World (SW) and the Normal World (NW). It also extends the concept of ”privilege rings” by introducing Exception Levels in the ARMv8 Instruction Set Architecture (ISA). The ”Secure Monitor” operates in the highest privilege level EL3, overseeing both the SW and NW. The untrusted OS runs within the NW, referred to as the Rich Execution Environment (REE). Within the SW, the trusted OS kernel operates in EL1, while the secure users-pace operates in EL0. The separation between SW and NW ensures that specific RAM regions and bus peripherals are designated as secure, limiting access solely to the SW.
Threat model of Arm TrustZone. Currently, there is a lack of a formal threat model for TrustZone. Commonly assumed attacker capabilities are as follows: The attacker (1) has full access to the REE of the device, encompassing the OS. This could be an authenticated (root) user, malicious third-party software installed on the device, or a compromised OS; (2) exploits unauthorized access to resources within the SW and extract sensitive information (Pinto and Santos, 2019).
TCB of Arm TrustZone. Within the SW, the trusted OS operates in EL1 and provides runtime support to sustain the lifecycle of Trusted Applications, ensuring their secure execution in user mode (EL0). The core of the trusted OS is the trusted kernel, which furnishes key OS primitives for the scheduling and management of TAs. The secure monitor implements mechanisms for secure context switching between worlds and operates with the highest privilege in EL3. Meanwhile, the TEE bootloader initiates the TEE system into a secure state (Cerdeira et al., 2020). In a nutshell, the trusted OS, secure monitor, and TEE bootloader constitute the TCB of a TrustZone-assisted TEE.
2.3.3. Virtualization-based TEEs
Process-based TEEs such as Intel SGX require partitioning an application into a trusted part, the enclave, and an untrusted part, which can complicate development efforts as developers must adhere to this security model and segment their applications accordingly. In response, vendors now iterated on their previous trusted computing CPU extensions to facilitate virtualization-based confidential computing across various server platforms, eliminating the need for extensive modifications to application code. Here, we outline three such virtualization-based confidential computing solutions: Intel TDX (Intel, 2023b), AMD SEV (AMD, 2023) and Arm CCA (Arm, 2023a).
AMD SEV. AMD introduced SEV with the Zen architecture to enable confidential computing in cloud environments. Recent AMD SEV- Secure Nested Paging (SNP(Sev-Snp, 2020) extends the capabilities of AMD SEV by introducing features such as secure nested paging to enhance the integrity of encrypted memory.
Arm CCA. Arm CCA is an advanced isolation technology introduced in Armv9. It preserves TrustZone’s NW and SW while introducing a new realm world designed for running multiple confidential “realms”. These realms, though isolated from other worlds, are managed by untrusted software components (e.g., hypervisor). A lightweight Realm Management Monitor (RMM(Arm, 2023b), operating within the hypervisor layer of the realm world, ensures memory isolation between realms. CCA incorporates a root world to accommodate the highest privilege monitor code stored in firmware (Wang et al., 2024b).
Intel TDX. Intel TDX leverages Intel’s Multi-Key Total Memory Encryption (MKTME(Intel, 2023a) and a CPU-attested software module to facilitate the isolated execution of secure VMs. The TDX module, digitally signed and executed in a novel processor mode called SEAM, functions as a separation kernel. It provides an interface to the hypervisor for scheduling, creating, and managing secure VMs, which are called trusted domains. For an in-depth comparison of AMD SEV, Arm CCA, and Intel TDX, we refer to the survey in (Guanciale et al., 2022).
3. Overview and Taxonomy
{forest}
forked edges, for tree= grow=east, reversed=true,anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, rounded corners,align=left, minimum width=2.5em, s sep=3pt, inner xsep=2pt, inner ysep=1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.8em,font=,, where level=2text width=5.6em,font=, where level=3text width=9.0em,font=,[Confidential Computing on Heterogeneous CPU-GPU Systems, ver [GPU TEEs
(Section 4) [x86-based
GPU TEEs [Graviton (Volos et al., 2018), HIX (Jang et al., 2019), HETEE (Zhu et al., 2020), NVIDIA H100 (NVIDIA, 2022a), TNPU (Lee et al., 2022), LITE (Yudha et al., 2022), SAGE (Ivanov et al., 2023), Honeycomb (Mai et al., 2023),leaf,text width=38em] ] [Arm-based
GPU TEEs [StrongBox (Deng et al., 2022), CRONUS (Jiang et al., 2022), CAGE (Wang et al., 2024b),leaf,text width=15em] ] ] [Security of GPUs
(Section 5) [Architectural/
Logical Attacks [Lee et al. (Lee et al., 2014), Maurice et al. (Maurice et al., 2014), Pietro et al. (Pietro et al., 2016), Zhou et al. (Zhou et al., 2016), HE et al. (HE et al., 2020),leaf,text width=30em] ] [Microarchitectural
Side/Covert-
Channel Attacks [Attacks on Integrated GPU [Frigo et al. (Frigo et al., 2018), HE et al. (HE et al., 2020), Dutta et al. (Dutta et al., 2021), Taneja et al. (Taneja et al., 2023), Wang et al. (Wang et al., 2024a),leaf,text width=31em] ] [Attacks on Discrete GPU [Jiang et al. (Jiang et al., 2016), Jiang et al. (Jiang et al., 2017), Ahn et al. (Ahn et al., 2021a), Giner et al. (Giner et al., 2024), Luo et al. (Luo et al., 2019), Zhu et al. (Zhu et al., 2021),
Hu et al. (Hu et al., 2020), Naghibijouybari et al. (Naghibijouybari et al., 2017), Wei et al. (Wei et al., 2020), Ahn et al. (Ahn et al., 2021b), Nayak et al. (Nayak et al., 2021),
Cronin et al. (Cronin et al., 2021), Yang et al. (Yang et al., 2022), Dutta et al. (Dutta et al., 2023), Zhang et al. (Zhang et al., 2023),leaf,text width=32em] ] [Countermeasures [Wang and Lee (Wang and Lee, 2007), Kong et al. (Kong et al., 2009), Liu et al. (Liu et al., 2016), Zhang et al. (Zhang et al., 2023), Wang et al. (Wang et al., 2014),
Shafiee et al. (Shafiee et al., 2015), Di et al. (Di et al., 2021), Kadam et al. (Kadam et al., 2018), Kadam et al. (Kadam et al., 2020), Lin et al. (Lin et al., 2019),leaf,text width=30em] ] ] [Physical Side-
Channel Attacks [Physical Access-based [Luo et al. (Luo et al., 2015), Gao et al. (Gao et al., 2018), Hu et al. (Hu et al., 2020), Chmielewski and Weissbart (Chmielewski and Weissbart, 2021), Maia et al. (Maia et al., 2022),
Horvath et al. (Horvath et al., 2023), Taneja et al. (Taneja et al., 2023),leaf,text width=32em] ] [Remote Access-based [Zhan et al. (Zhan et al., 2022), Liang et al. (Liang et al., 2022),leaf,text width=12em] ] [Countermeasures [Mangard et al. (Mangard et al., 2008), Zhan et al. (Zhan et al., 2022),leaf,text width=13em] ] ] [Fault Injection
Attacks [Software-based [Frigo et al. (Frigo et al., 2018), Sabbagh et al. (Sabbagh et al., 2020), Sabbagh et al. (Sabbagh et al., 2021), Sun et al. (Sun et al., 2023), Santos et al. (Santos et al., 2021),leaf,text width=32em] ] ] ] [Security of CPU
TEEs (Section 6) [Security of Intel
SGX [Controlled Channel/Address
Translation-based Attacks [Xu et al. (Xu et al., 2015), Shinde et al. (Shinde et al., 2016), Van Bulck et al. (Van Bulck et al., 2017a), Van Bulck et al. (Van Bulck et al., 2017b), Wang et al. (Wang et al., 2017),
Gyselinck et al. (Gyselinck et al., 2018), Kim et al. (Kim et al., 2019), Fei et al. (Fei et al., 2021),leaf,text width=33em] ] [Cache-based Attacks [Götzfried et al. (Götzfried et al., 2017), Moghimi et al. (Moghimi et al., 2017), Schwarz et al. (Schwarz et al., 2017), Brasser et al. (Brasser et al., 2017b), Dall et al. (Dall et al., 2018),
Moghimi et al. (Moghimi et al., 2019),leaf,text width=32em] ] [Branch Prediction Attacks [Fei et al. (Fei et al., 2021), Lee et al. (Lee et al., 2017), Chen et al. (Chen et al., 2019), Evtyushkin et al. (Evtyushkin et al., 2018), Huo et al. (Huo et al., 2020), Koruyeh et al. (Koruyeh et al., 2018),leaf,text width=34em] ] [Transient Execution Attacks [Chen et al. (Chen et al., 2019), Van Bulck et al. (Van Bulck et al., 2018), Van Schaik et al. (Van Schaik et al., 2020), Ragab et al. (Ragab et al., 2021),leaf,text width=27em] ] [Software-based Fault
Injection Attacks [Murdock et al. (Murdock et al., 2020), Kenjar et al. (Kenjar et al., 2020), Qiu et al. (Qiu et al., 2020), Jang et al. (Jang et al., 2017),leaf,text width=24em] ] [Hardware-based Attacks [Chen et al. (Chen et al., 2021), Chen and Oswald (Chen and Oswald, 2023), Lee et al. (Lee et al., 2020),leaf,text width=18em] ] [Interface-based Attacks [Van Bulck et al. (Van Bulck et al., 2019), Alder et al. (Alder et al., 2020), Cloosters et al. (Cloosters et al., 2020), Khandaker et al. (Khandaker et al., 2020), Alder et al. (Alder et al., 2024),leaf,text width=32em] ] ] [Security of
Arm TrustZone [Implementation Bugs [Beniamini (Beniamini, 2016), Shen (Shen, 2015), Komaromy (Komaromy, 2018),leaf,text width=15em] ] [Cache-based Attacks [Zhang et al. (Zhang et al., 2016c), Guanciale et al. (Guanciale et al., 2016), Lipp et al. (Lipp et al., 2016), Zhang et al. (Zhang et al., 2016b), Cho et al. (Cho et al., 2018),leaf,text width=30em] ] [Branch Prediction Attacks [Ryan (Ryan, 2019),leaf,text width=4em] ] [Software-based Fault
Injection Attacks [Tang et al. (Tang et al., 2017), Qiu et al. (Qiu et al., 2020), Carru (Carru, 2017),leaf,text width=16em] ] ] [Security of
AMD SEV [Memory Encryption Issues [Hetzelt and Buhren (Hetzelt and Buhren, 2017), Buhren et al. (Buhren et al., 2017), Du et al. (Du et al., 2017), Morbitzer et al. (Morbitzer et al., 2018), Li et al. (Li et al., 2021b),
Li (Li, 2022), Li et al. (Li et al., 2021a), Li et al. (Li et al., 2021c),leaf,text width=32em] ] [Software-based Fault
Injection Attacks [Morbitzer et al. (Morbitzer et al., 2021), Buhren et al. (Buhren et al., 2021), Zhang et al. (Zhang et al., 2024), Schlüter et al. (Schlüter et al., 2024a), Schlüter et al. (Schlüter et al., 2024b),leaf,text width=33em] ] [Interrupt-driven Attacks [Wilke et al. (Wilke et al., 2024),leaf,text width=6em] ] ] ] [Potential Attacks
against GPU TEEs
(Section 7) [x86-based
GPU TEE [Existing Attacks against Intel SGX,leaf,text width=12em] [ Dynamic Voltage and Frequency Scaling (DVFS)-based Attacks,leaf,text width=21em] [Physical Attacks on PCIe bus and GPU Memory,leaf,text width=17em] [Architectural/Logical attacks on GPU,leaf,text width=13em] ] [NVIDIA Hopper
H100-based
GPU TEE [Existing attacks against VM-based TEEs,leaf,text width=14em] [DVFS-based Attacks,leaf,text width=8em] [Microarchitectural Side/Covert-Channel Attacks on GPU,leaf,text width=19em] ] [Arm-based
GPU TEE [Existing attacks against Arm TrustZone,leaf,text width=13em] [Microarchitectural Side/Covert-Channel Attacks on Integrated GPU,leaf,text width=22em] [Physical Attacks on GPU Memory,leaf,text width=12em] [DVFS-based Attacks,leaf,text width=8em] ] ] ]
Figure 3. Taxonomy of confidential computing on heterogeneous CPU-GPU systems
Confidential computing on GPUs is an emerging field that ensures sensitive data processed on GPUs remains protected from unauthorized access and tampering even for attackers with privileged access. Other privacy-preserving techniques, such as secure multi-party computation, homomorphic encryption, and TEEs, offer different security guarantees. However, in this work, our primary focus is on the extension of TEEs to GPUs (i.e., GPU TEEs). Moving from homogeneous to such heterogeneous systems inevitably creates new threats and security vulnerabilities. It is crucial to understand these new attack vectors to inform the design of systems that can effectively mitigate them. In comparison to existing CPU TEEs, GPU TEEs present unique opportunities for attackers from several perspectives:
GPU software stack on CPU: GPU drivers handle critical aspects such as resource allocation, task scheduling, and dynamic power management. Specifically, in resource allocation, GPU drivers manage and allocate resources, including memory, processing units, and other hardware components. For task scheduling, they efficiently schedule and manage tasks to ensure optimal performance, balancing workloads across GPU cores. In dynamic power management, drivers adjust parameters such as voltage based on workload, optimizing performance and efficiency to extend battery life in mobile devices and reduce energy costs in data centers. In current GPU TEEs, the protection of GPU drivers, particularly their positioning in the CPU host, may introduce new attack surface. Furthermore, some state-of-the-art attacks against CPU TEEs may also affect GPU drivers.
Hardware components on GPU: GPU TEEs are still in the early stages of development. Most designs are experimental research and often lack extensive hardware modifications, resulting in insufficient hardware protection for components such as GPU memory, the Peripheral Component Interconnect Express (PCIe) bus, and power management components. Even products designed by GPU manufacturers (e.g., NVIDIA H100 GPU) might lack protection for certain hardware components, such as power management components. Thus, existing attacks against these components may also translate to weaknesses in GPU TEEs.
Architectural complexity: Although the security of GPUs is just starting to be explored, several vulnerabilities and attacks have already been documented. Due to the architectural complexity of GPUs, such as memory sharing among different concurrent kernels and the variety of GPU types (e.g., integrated and discrete), some existing GPU TEEs may not provide adequate protection. Additionally, many designs overlook cache-level access protection, making them susceptible to cache-based attacks on GPUs.
In Figure 3, we summarise our taxonomy underlying this survey based on relevant related work. In the remainder of this paper, we first review and discuss existing GPU TEEs in Section 4, covering their threat models, hardware modifications, TCB size, and performance overhead. In Section 5, we give an overview of existing attacks targeting GPUs based on the attack types listed in Section 2. We also review state-of-the-art attacks against existing CPU TEEs, such as Intel SGX, Arm TrustZone, and AMD SEV in Section 6. Finally, based on these GPU TEE designs and the summarized attacks, we discuss potential attacks against existing (and future) GPU TEEs in Section 7.
4. Confidential Computing on GPUs
We summarize the state-of-the-art GPU TEE designs based on Intel, AMD, and ARM in Table 2. Next, we first review all existing schemes and then discuss them based on aspects such as threat models, hardware modifications, TCB size and performance overhead resulting from implementing trusted execution on accelerators.
4.1. Trusted Execution Environments on GPUs
Traditional CPUs and cloud systems have slowly adopted hardware-based TEEs to securely isolate computations from a malicious OS or certain hardware attacks, such as snooping on the memory bus. However, GPUs and their cloud deployments have largely yet to incorporate such support for hardware-based trusted computing. With large amounts of sensitive data being offloaded to GPUs for acceleration in cloud environments, especially for sensitive ML workloads, there is an urgent need to address the trusted computing requirements for ubiquitous GPUs. Researchers and commercial vendors have recently proposed various designs to support confidential computing on GPUs.
Table 2. Summary of proposed heterogeneous TEE for accelerators
Scheme ISA Accelerator Memory GPU Driver Location Methods HW Changes Key Components of GPU TEE
Attestation Secure RM Secure I/O Integrity GPU ME
Graviton (Volos et al., 2018) Intel x86 NVIDIA GPU Dedicated Untrusted OS Modify GPU command processor $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
HIX (Jang et al., 2019) Intel x86 NVIDIA GPU Dedicated Part inside CPU enclave SGX Enclave $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
HETEE (Zhu et al., 2020) Intel x86 NVIDIA GPU Dedicated Inside CPU enclave Dynamic allocation of accelerators within CPU enclave and extra FPGA $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
NVIDIA H100 (NVIDIA, 2022a) Intel x86 NVIDIA GPU Dedicated Inside confidential VM Hardware-level protection $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
[…]
LITE (Yudha et al., 2022) Intel/AMD x86 NVIDIA GPU Dedicated N/A Software-based unified encryption domain $\checkmark$ $\checkmark$ $\checkmark$
[…]
SAGE (Ivanov et al., 2023) Intel x86 NVIDIA GPU Dedicated N/A Software-based verification function $\checkmark$ $\checkmark$
[…]
x86-based GPU TEEs. Graviton (Volos et al., 2018) initiated the work of trusted execution on hardware accelerators like GPUs. Graviton directly establishes a TEE within the GPU, necessitating hardware modifications but avoiding the need for CPU-side TEE like SGX. Graviton establishes a secure channel between the untrusted host and GPU, monitoring command submission and data transfer by directing all resource allocation requests through the GPU’s command processor instead of GPU driver. Nevertheless, for data protection, Graviton assumes the GPU’s on-package memory can be trusted. The assumption is that even with physical access, it is extremely difficult for an attacker to breach the integrity of the GPU package and snoop on the silicon interconnect between GPU and stacked memory.
HIX (Jang et al., 2019) removes the key functionalities for controlling the GPU from the OS-resident driver and places those within a trusted GPU enclave, extending an SGX enclave and incorporating new SGX-style instructions like $\mathtt{ECREATE}$ and $\mathtt{EADD}$. Furthermore, HIX establishes a secure hardware I/O path, preventing the OS from altering the virtual-to-physical address mapping for the GPU Memory-Mapped I/O (MMIO) region and from deducing commands and data transmitted via MMIO and Direct Memory Access (DMA). Like Graviton, HIX lacks a trusted GPU memory region, leaving data in GPU memory accessible in plaintext.
To eliminate the need for modifications to CPU or GPU chips, HETEE (Zhu et al., 2020) introduces a data-center level GPU TEE designed to support extensive confidential computing, without any chip-level change. It facilitates dynamic allocation of computing resources for both secure and non-sensitive computing tasks across all servers within a rack. They incorporate a small TCB, such as a single security controller with customized code supporting task isolation, reducing design complexity and minimizing the attack surface associated with resource sharing with the untrusted OS.
Similar to HIX, Honeycomb (Mai et al., 2023) ensures secure communication between the user and the GPU by using a CPU-side TEE (e.g., AMD SEV-SNP) to transfer sensitive data to the GPU and manage access to GPU MMIO. Unlike previous designs, Honeycomb employs static analysis to confirm that mutually distrusting GPU applications are contained within their enclaves. This approach enables more efficient implementations by moving runtime checks to load time. Moreover, Honeycomb reduces the TCB by excluding both the user-space and kernel-space GPU drivers from it, instead using two security monitors to intercept and control all traffic between applications and the GPU. Honeycomb makes the same assumption as Graviton in trusting the GPU’s device memory based on the fact that modern GPUs typically integrate device memory using 2.5D/3D silicon interposers within the same package.
SAGE (Ivanov et al., 2023) offers a software-based alternative for trusted GPU execution. It employs an SGX enclave running on the host as a local verifier and uses this enclave to bootstrap a software primitive that establishes a dynamic root of trust on the GPU. It ensures that the user kernel on the untrusted device remains unmodified, is invoked for execution on the untrusted GPU, and runs without tampering. It complements existing hardware-based GPU TEEs.
Given that previous designs ignore memory security (i.e., they do not encrypt GPU memory), LITE (Yudha et al., 2022) introduces a co-designed GPU TEE with CPU to establish a unified encryption domain, where data is stored in ciphertext form in GPU caches and memory. Specifically, LITE ensures that data enters the GPU chip encrypted. The software then decrypts the data using the Arithmetic Logic Unit (ALU), operating on the data in registers. Additionally, data may be temporarily stored in plaintext in the on-chip shared memory.
TNPU (Lee et al., 2022) proposes tree-less off-chip memory protection for Deep Neural Networks accelerators. The core idea is to group contiguous memory blocks into tiles and provide freshness guarantees at the tile level. TNPU employs a dedicated software module running on the host CPU to manage version numbers, which are stored in a “Tensor Table” within the host CPU’s secure memory. The tensor table is protected by an integrity tree.
NVIDIA’s H100 Tensor Core GPU (NVIDIA, 2022a), based on the Hopper architecture, introduces advanced features for confidential computing. To enable confidential computing with the H100, it requires specific CPU TEEs, such as Intel TDX, AMD SEV-SNP, and Arm CCA. In addition to ensuring the confidentiality and integrity of data and code, H100 can defend against basic physical attacks targeting the PCIe bus and Double Data Rate RAM (DDR). H100 supports three different operation modes, namely CC-Off, CC-On, and CC-DevTools. In CC-On mode, the H100, along with the drivers on the CPU, fully activates all available confidential computing features. Some hardware resources, such as performance counters, are disabled in CC-On mode to prevent potential side-channel attacks, as performance counters could be used to infer the behavior of device usage (Naghibijouybari et al., 2018). However, details about specific hardware changes, such as memory encryption in the H100, remain unclear at present.
Arm-based GPU TEEs. Previous designs for Intel/AMD-based platforms cannot be readily extended to Arm GPUs due to numerous architectural differences. StrongBox (Deng et al., 2022) thus introduces the first GPU TEE for secure general computation on Arm endpoints without necessitating hardware modifications or architectural changes. It deploys the StrongBox runtime within the secure monitor to ensure that the GPU executes in isolation. Through a combination of Stage-2 translation and the TrustZone Address Space Controller (TZASC), StrongBox specifies six distinct access permission types to effectively manage secure access from DMA, GPU, and other peripherals.
CRONUS (Jiang et al., 2022) introduces a design that supports various heterogeneous accelerators, enables spatial sharing within a single accelerator, and ensures robust isolation across multiple accelerators. Leveraging Arm secure virtualization technology, CRONUS partitions heterogeneous computation into isolated TEE enclaves. Each enclave encapsulates a specific type of computation like GPU computation and allows multiple enclaves to spatially share a single accelerator.
To support confidential GPU computing on next-generation Arm devices, CAGE (Wang et al., 2024b) presents a new scheme for Arm CCA’s realm-style architecture. It offloads resource-intensive but data-independent tasks to NW, while protecting sensitive data within realms through a shadow task mechanism. By configuring Granule Protection Check (GPC), CAGE restricts access from untrusted components. Additionally, it isolates realms in GPU computing by providing each realm with a distinct GPU memory view in the GPU GPC. CAGE retains high hardware compatibility, as it requires no additional hardware or modifications to existing CPU and GPU configurations.
4.2. Discussion
In this section, we discuss several key metrics in designing an effective and secure GPU TEEs.
Threat models. Essentially all GPU TEE designs adopt a threat model similar to that of CPU TEEs, aiming to create secure execution environments for GPUs. Specifically, they consider an adversary capable of controlling the entire software stack, including device drivers, the guest OS, and the hypervisor. Additionally, the adversary is assumed to have limited physical access to the hardware, allowing for passive physical attacks, such as snooping on PCIe traffic. This means that such an adversary could potentially leak or tamper with sensitive data and execution results of GPU applications. They could also access or tamper with user data in DMA buffers or with commands submitted by the victim application to the GPU.
Regarding GPU memory protection, many designs assume that the adversary cannot access secrets in the GPU memory or explicitly state that physical attacks targeting memory are out of their scope. For instance, Graviton (Volos et al., 2018) and Honeycomb (Mai et al., 2023) trust the device memory of the GPU, as modern GPUs typically integrate device memory using 2.5D/3D silicon interposers within the same package. These designs mainly focus on preventing adversaries from accessing sensitive data stored in GPU memory with finer-grained access control. To fill this gap, LITE (Yudha et al., 2022) designs a software-based unified encryption domain to protect sensitive data stored in GPU memory, including L1/L2 cache and device memory. Additionally, NVIDIA H100 (NVIDIA, 2022a) and TNPU (Lee et al., 2022) support hardware-based memory encryption, while CAGE (Wang et al., 2024b) requires hardware-assisted memory encryption support built into future CCA devices.
Hardware changes. Earlier works like Graviton (Volos et al., 2018) and HIX (Jang et al., 2019) require non-trivial modifications to peripheral components, such as the GPU’s command processor, the introduction of new SGX instructions, specific Memory Management Unit (MMU) page table walkers, and other hardware changes. The extent to which a GPU TEE design modifies hardware components determines its practical compatibility, i.e., its suitability for various GPUs. Many designs necessitate hardware changes that have long lead times before they can be deployed in production environments. Consequently, works like LITE (Yudha et al., 2022), SAGE (Ivanov et al., 2023), and Honeycomb (Mai et al., 2023), seek purely software-based approaches for trusted GPU execution. CAGE (Wang et al., 2024b) primarily leverages generic hardware features in Arm CCA and GPUs without requiring additional hardware changes or customization of existing CPUs and GPUs.
TCB size. One of the key challenges in GPU TEE design is reducing the TCB size as much as possible. A large software stack increases the TCB, which can introduce more vulnerabilities and hence threaten security. An effective and efficient way to reduce the TCB is to place most of the GPU software stack in the untrusted OS, while keeping only the key functions for controlling the GPU in a trusted area of the CPU, such as an SGX enclave. For example, CAGE (Wang et al., 2024b) achieves a ‘thin’ TCB for realms by using its shadow task mechanism instead of loading the heavyweight GPU software into the TEE. It introduces 2–-26K lines of code for a CPU-side isolation and 1,301 lines of code for monitor, resulting in the lowest TCB size among current Arm-based GPU TEEs.
Performance overhead. The ultimate goal of designing a GPU TEE is to ensure strong isolation of GPU execution while minimizing the performance overhead associated with providing security guarantees. Comparing the performance overhead of different schemes is challenging due to the costly re-implementation efforts required, as each scheme targets different platforms and is implemented under varying conditions. For example, in HIX, the software components are implemented on top of an emulated system (like KVM and QEMU), which supports hardware modifications. Several factors, such as the number of cryptographic operations, the frequency of switching between secure and normal worlds, and the integrity checks, play significant roles in optimizing performance compared to insecure baselines on GPUs. For example, CAGE (Wang et al., 2024b) outperforms StrongBox (Deng et al., 2022) by 1.96%–13.16% on six Rodinia benchmarks. This performance gain is attributed to CAGE’s ability to securely transfer plaintext data to buffers and store the results without introducing additional cryptographic operations. It is reasonable to speculate that the NVIDIA H100 GPU introduces a smaller performance burden due to its hardware-based encryption support. It would be insightful to conduct a comprehensive performance comparison across different platforms once the entire workflow for confidential computing on the NVIDIA H100 GPU is fully available in the future.
5. Security of GPUs
Due to their widespread adoption as accelerators, GPUs have come under scrutiny from both offensive and defensive perspectives in recent years. Researchers have explored GPU vulnerabilities at hardware architecture and software levels, showcasing practical attacks and suggesting mitigations. We believe that summarizing and analyzing attacks on GPUs will facilitate the exploration of potential attacks on GPU TEEs. While Naghibijouybari et al. (Naghibijouybari et al., 2022) offer a comprehensive survey categorizing GPU security vulnerabilities and potential countermeasures, their analysis does not cover significant advancements that have appeared since the latter half of 2022. Moreover, their emphasis primarily is on microarchitectural attacks across diverse accelerators without considering attacks that leverage physical access or measurements. Nonetheless, the consideration of physical channels is essential for potential attacks on GPU TEE. In this section, we review existing GPU attacks, including microarchitectural side-channel/covert-channel attacks, physical side-channel attacks, and fault injection attacks. LABEL:tab:GPU_attacks provides a summary of attacks, target GPUs, threat model, attack goals and methods.
{ThreePartTable}{TableNotes}
CNN/DNN denotes Convolutional and Deep Neural Network. In Attacks, SCA, CovA, and FIA denote Side-Channel Analysis, Covert-Channel, and Fault Injection Attacks, respectively. AA/LA represents an attack that exploits an architectural or logical vulnerability on GPU. Physical denotes whether this attack leverages physical properties (e.g., power consumption, EM, acoustics, etc.). In Model, sGPU, mGPU, and iGPU denote the attack is deployed on a single GPU, multiple GPUs or integrated GPU, respectively. Note that multiple GPUs represent multiple discrete GPU cards or multiple GPU instances (e.g., NVIDIA’s MIG feature). In Main results, we report the central outcomes, e.g., bandwidth and classification accuracy for CovA and SCA, respectively.
Table 3. Comparison of different attacks on heterogeneous systems (oldest results first)
Work&Year Accelerator Attack Threat Model Phy- sical Attack Scenarios & Goals Method Main results
Model Permissions
Lee et al.’14 (Lee et al., 2014) NVIDIA/AMD GPU AA/LA sGPU User-level permissions Infer which web pages a victim user has visited Exploit that GPUs do not initialize newly allocated memory pages $95.4\%$ accuracy
Maurice et al.’14 (Maurice et al., 2014) NVIDIA GPU AA/LA sGPU Require root privileges based on accessing GPU memory via GPU runtime or PCI configuration Recover data of victim’s previously execution on GPU in virtualized environment Exploit the vulnerability of memory clearing on GPU Access data stored in GPU global memory after soft reboot
Luo et al.’15 (Luo et al., 2015) NVIDIA GPU SCA sGPU Unprivileged attacker has physical access to the victim GPU $\checkmark$ Recover all of the last round Advanced Encryption Standard (AES) key byte Apply correlation power analysis to analyze the power consumption of the GPU Recover the key with 160,000 power traces
Pietro et al.’16 (Pietro et al., 2016) NVIDIA GPU AA/LA sGPU User-level permissions Read information stored in memory and infer the AES keys Leverage the lack of adequate data allocation and data cleaning $70\%$, $28\%$ leakages
Zhou et al.’16 (Zhou et al., 2016) NVIDIA/AMD GPU AA/LA sGPU User-level permissions Recover users’ profile images and email contents Reconstruct images from identified image-like tiles in GPU memory $79.3\%$ successful recovery
Jiang et al.’16 (Jiang et al., 2016) NVIDIA GPU SCA sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Leverage the relation between encryption kernel’s runtime and the key 30 min with 1M timing samples
Jiang et al.’17 (Jiang et al., 2017) NVIDIA GPU SCA sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Exploit timing variability due to shared memory bank conflicts. Use 10M timing samples
Naghibijouybari et al.’17 (Naghibijouybari et al., 2017) NVIDIA GPU CovA sGPU Trojan and spy co-exist on a GPU, both with user-level permission Infer transmitted data Reverse engineer warp schedulers; learn contention on caches and memory $4$ MB/s (error-free)
Frigo et al.’18 (Frigo et al., 2018) Intel HD, Qualcomm Adreno SCA, FIA iGPU Access iGPU through a malicious application or when victim visits malicious JavaScript website Escape the Firefox sandbox on Android platforms and break ASLR Detect contiguous memory allocation via SCA and then trigger Rowhammer bit flips Hijack the browser and break ASLR in 116 s and 27 s
Gao et al.’18 (Gao et al., 2018) NVIDIA GPU SCA sGPU Unprivileged attacker has physical access to the victim GPU $\checkmark$ Recover AES private key Sample MMIO traces via trigger; exploit parallel MMIO leakages Recover the key with 11k MMIO traces
Naghibijouybari et al.’18 (Naghibijouybari et al., 2018) NVIDIA GPU SCA sGPU Spy co-locates with the victim program with user-level permission Infer victim’s accessed website data and the number of neurons Learn contention via memory allocation, performance counter, and timing $90\%$, $94.8\%$, $88.6\%$ accuracy
Luo et al.’19 (Luo et al., 2019) NVIDIA GPU SCA sGPU User-level permissions; available to timing information Recover RSA private key Capture the parallel characteristics of an RSA on GPU via timing model Recover 512/2048-bit key in 15 min/8 h
Sabbagh et al.’20 (Sabbagh et al., 2020) AMD GPU FIA sGPU DVFS interfaces access Recover AES key Exploit DVFS to introduce random faults during kernel execution Recover key in 15 min
Wei et al.’20 (Wei et al., 2020) NVIDIA GPU SCA sGPU Spy co-locates with the victim program with user-level permission Recover the layer composition and hyper-parameters of DNN Leverage context-switching penalties when switching different kernels $76.9\%$-$100\%$ accuracy
Hu et al.’20 (Hu et al., 2020) NVIDIA GPU SCA sGPU Physically access the victim GPU with user-level permissions $\checkmark$ Obtain the complete model architecture information Learn the inter-layer temporal association information from kernel execution $75.9\%$ accuracy
HE et al.’20 (HE et al., 2020) Intel HD 530 AA/LA, CovA iGPU Unprivileged adversary and the victim co-locate on the same GPU, but are isolated (1) Recover AES key; (2) monitor the browser activity; (3) conduct a covert channel to exchange data Exploit context management with uncleared registers and shared memory during context switches Recover AES-128/192 in 0.15 s/2 min; $90.7\%$ accuracy; 1.3–8 Gbps
Santos et al.’21 (Santos et al., 2021) NVIDIA GPU FIA sGPU Inject faults into pipeline registers, ALU, and memory Improve the probability of incorrect output by injecting faults Combine register transfer level FIA with software FIA Underestimate the error rate by up to $48\%$
Ahn et al.’21 (Ahn et al., 2021b) NVIDIA GPU CovA sGPU Trojan and spy co-exist on a GPU, both with user-level permissions Infer transmitted data Exploit interconnect channel contention $24$ MB/s
Cronin et al.’21 (Cronin et al., 2021) Apple GPU SCA sGPU Exploit JavaScript website and trick user into installing a malicious app Track the user’s web activity Exploit the shared cache between the CPU-cores and peripherals $90\%$, $95\%$ accuracy
Zhu et al.’21 (Zhu et al., 2021) NVIDIA GPU Snooping sGPU Install a PCIe bus snooping device to monitor and log the PCIe traffic Recover the DNN model (semantically identical architecture) Reverse engineer GPU command headers; collect PCIe packets via snooping Similar accuracy as original ones
Ahn et al.’21 (Ahn et al., 2021a) NVIDIA GPU SCA sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Recover early/later key bytes via timing correlation and cache-collision Use 100k timing samples
Dutta et al.’21 (Dutta et al., 2021) Intel HD Graphics CovA iGPU Spy and trojan run on CPU and GPU with user-level permissions Infer transmitted data and track user’s cache activity CovA1: prime+probe attack on LLC; CovA2: observe delays on the ring bus CovA1: 120 KB/s, CovA2: 400 KB/s
Nayak et al.’21 (Nayak et al., 2021) NVIDIA GPU CovA sGPU User-level permissions Infer the accessed data from Virginian database Leverage the shared last-level TLB (reverse engineering) and prime+probe 81 KB/s (MPS-enabled)
Yang et al.’22 (Yang et al., 2022) Qualcomm Adreno SCA sGPU Install and launch an eavesdropping application on the victim device Eavesdrop the user’s input credentials (login usernames and passwords) through onscreen keyboard Read raw values from the GPU device file and identify GPU events from key presses via ML models Correctly infer more than $80\%$ of user’s credential inputs
Zhan et al.’22 (Zhan et al., 2022) NVIDIA/AMD GPU SCA sGPU Unprivileged attacker eavesdrops on a victim $\checkmark$ Website fingerprinting and keystroke timing inference Exploit EM caused by DVFS $95.6\%$ accuracy and $90\%$ keystrokes
Liang et al.’22 (Liang et al., 2022) NVIDIA GPU SCA sGPU Attacker and victim GPUs may be separated by meters $\checkmark$ Infer the number of layers/kernels, sizes of layers and strides Exploit far-field EM signals emanated from a GPU Recover VGG’s and ResNet’s parameters
Maia et al.’22 (Maia et al., 2022) NVIDIA GPU SCA sGPU Unprivileged attacker with physical proximity $\checkmark$ Infer neural network architectures and layer parameters Exploit the magnetic flux emanating from a GPU’s power cable Up to $96.8\%$ accuracy
Horvath et al.’23 (Horvath et al., 2023) NVIDIA GPU SCA sGPU User-level permission with physical access and know the CNN architecture $\checkmark$ Recover the weights and biases of CNN Correlation electromagnetic analysis N/A
Dutta et al.’23 (Dutta et al., 2023) NVIDIA GPU SCA, CovA mGPU Spy and victim connect via NVLink with user-level permissions CovA: infer message; SCA: infer the number of neurons of NN Mount Prime+Probe attack on the L2 cache of victim GPU CovA: 4 MB/s; SCA: $99.91\%$ accuracy
Zhang et al.’23 (Zhang et al., 2023) NVIDIA GPU CovA mGPU Unprivileged spy and trojan on an MIG-enabled GPU Infer transmitted data and classify the running ML frameworks Mount Prime+Probe attack on the shared L3-TLB $31.47$ KB/s
Taneja et al.’23 (Taneja et al., 2023) Various GPUs SCA iGPU, sGPU Read frequency, power, and temperature from internal sensors (may require privileged access) $\checkmark$ (1) Observe CPU/GPU behaviour; (2) recover images and user’s browsing history; (3) fingerprint websites (1) Use data from internal sensors; (2) observe throttling via timing; (3) leverage frequency and power bursts on iGPUs Pixel stealing with accuracy of 60%–90%
Wang et al.’24 (Wang et al., 2024a) Various GPUs SCA iGPU, sGPU Observe coarse-grained pattern redundancy using timer Perform cross-origin pixel stealing attacks in the browser Infer data-dependent DRAM traffic and cache usage via compression channel Pixel stealing with accuracy 68.6%–99.6%
Giner et al.’24 (Giner et al., 2024) NVIDIA/AMD GPU SCA, CovA sGPU SCA: Victim passively visits attacker’s webpage; CovA: Spy accesses the website visited by victim SCA: Detect typed characters and recover the victim’s AES key; CovA: Infer transmitted data Exploit cache contention on the L2 cache via Prime+Probe attack, parallelize eviction set construction SCA: $98\%$ accuracy, recover key in 6  min; CovA: $10.9$ KB/s
[…]
5.1. Architectural/Logical Attacks
There has been an increase in research examining attacks that target GPUs or their software stack on the CPU. Schemes (Lee et al., 2014; Maurice et al., 2014; Pietro et al., 2016; Zhou et al., 2016) focus on information leakage due to leftover memory from processes that recently terminated. Lee et al. (Lee et al., 2014) exploit the vulnerability arising from the GPUs not initializing newly allocated memory pages, thereby revealing the victim’s data stored in GPU memory both during its execution and immediately after termination. In practice, they infer the webpages visited by a victim user when analyzing remaining textures from Chromium and Firefox. Maurice et al. (Maurice et al., 2014) recover data from previously executed GPU applications by observing that GPU global memory is not always zeroed. Pietro et al. (Pietro et al., 2016) leverage interleaved access and inadequate data handling to read information stored in shared or global memory by another process. Zhou et al. (Zhou et al., 2016) demonstrate how malicious programs can exploit GPU memory management strategies to breach the memory isolation boundary and recovers images from other processes. HE et al. (HE et al., 2020) observed a vulnerability in Intel integrated GPU caused by flawed context management, where residual register values and shared memory are not cleared during a context switch. This class of vulnerability can usually be fixed by clearing memory (i.e., memory zeroing) before reallocating it to another application.
5.2. Microarchitectural Side-Channel/Covert-Channel Attacks
In this subsection, we review existing attacks on GPUs w.r.t. microarchitectural side-channel and covert-channel attacks.
5.2.1. Attacks on Integrated GPUs.
Attacks on integrated GPUs are gaining attention due to their security risks. The integration opens up new attack opportunities by exploiting shared resources, leading to cross-component microarchitectural attacks. A key requirement for these attacks is co-location, where the attacker runs code close enough to interact with the victim (Naghibijouybari et al., 2022). This interaction could involve inducing faults or contention to measure side-channel leakage. Therefore, integrated GPUs, which share the same chip as the CPU, offer more attack opportunities.
Dutta et al. (Dutta et al., 2021) develop two covert channels, enabling two malicious applications located on separate components (CPU and iGPU) to exchange confidential data through shared hardware resources. One channel exploits the shared Last-Level Cache (LLC) within Intel’s integrated GPU, while the other relies on the contention from the ring bus. Wang et al. (Wang et al., 2024a) develop cross-origin pixel stealing attacks in web browsers by exploiting the iGPU compression channel to induce data-dependent DRAM traffic and cache utilization. Given that the LLC is shared between the CPU and GPU, they demonstrate how to use LLC walk time to infer the GPU-induced CPU cache state changes. Wang et al. (Wang et al., 2021) extend the previous timing side-channel attack on AES implementation (Jiang et al., 2016) to the iGPU.
Some recent studies have also delved into GPUs such as the Qualcomm Adreno GPU and Apple GPU. Building upon Arm’s system-level cache, Cronin et al. (Cronin et al., 2021) exploit the GPU and shared cache architecture of Arm DynamIQ to create a website fingerprinting side channel. Yang et al. (Yang et al., 2022) target the Qualcomm Adreno GPUs, investigating the GPU overdraw induced by rendering popups of user key presses. Each key press leads to unique variations in selected GPU performance counters, allowing accurate inference of these key presses. To access GPU performance counters on Android, they read the raw values of GPU PCs from the GPU device file (e.g., Qualcomm Adreno GPU drivers’ open-source header file, msm_kgsl.h).
5.2.2. Attacks on Discrete GPUs.
Attacks on discrete GPUs can generally be categorized in two ways: (i) where the spy and victim (or trojan) applications are co-located on a single GPU, denoted as sGPU; and (ii) where the spy and victim (or trojan) applications are co-located on different GPUs, denoted as mGPU. For instance, this scenario could involve a system equipped with multiple GPUs, such as NVIDIA’s Pascal-based DGX-1 system comprising eight Tesla P100 GPUs. Alternatively, it could involve a system equipped with an MIG-enabled GPU, where a single physical GPU is partitioned into multiple instances, each acting as an independent GPU with dedicated resources, such as NVIDIA A100 and H100.
Attacks on a single GPU. This category of attacks typically assumes that an attacker with only user-level permissions either launches a GPU kernel on a single GPU to gain insights from measurements such as execution time, performance counters, and other leakage vectors; or co-locates with the victim process on a single GPU to establish a covert channel. In the following, we review the existing research and categorize them into the following aspects:
Recovering private encryption keys in GPU-based cryptography implementations: Recent works demonstrate the capability of recovering encryption keys (e.g., AES and Rivest Shamir and Adleman (RSA) private keys) from encryption applications running on GPUs. Jiang et al. (Jiang et al., 2016) exploit timing differences between addresses generated by different threads accessing memory, where execution time is affected by the number of unique memory requests after coalescing. Consequently, the key affects the address pattern accessed by threads, thereby impacting the observed runtime of the AES encryption. In their another study (Jiang et al., 2017), they establish a correlation between the execution time of a warp’s table lookup and the number of shared memory bank conflicts generated by threads within the warp. These key-dependent timing differences are then used to correlate the measured execution time with the key at the last round of the AES encryption as it executes on the GPU. Ahn et al. (Ahn et al., 2021a) exploits negative timing correlation to recover earlier key bytes of AES, while employing cache-collision attacks for the latter AES key bytes. More recently, Giner et al. (Giner et al., 2024) recover keys from an AES T-table GPU implementation. They allocate a sizable buffer to occupy a significant portion of the cache, execute an AES encryption, and then use Prime+Probe to identify evicted buffer offsets that correlate with the encryption’s inputs and outputs. They demonstrate superior performance by recovering the entire AES key within 6 min, whereas previous studies took between 15 and 30 min. Furthermore, aside from GPU-based AES implementation, Luo et al. (Luo et al., 2019) present a timing side-channel attack on an RSA implementation. They use the correlation between decryption time and the number of reductions per decryption window to extract the RSA private key.
(a) The attacker can snoop the PCIe traffic or GDDR5
using a bus snooping device
(b) Spy and victim are on two different applications (or processes and VMs) sharing the same GPU
Figure 4. Attacker models for inferring NN models on a single GPU.
Inferring sensitive information from GPU-based Neural Network (NN): Other attacks exploit side-channel leakage to infer information from ML models like NNs. Zhu et al. (Zhu et al., 2021) and Hu et al. (Hu et al., 2020) focus on the attack surface of the bus, particularly PCIe (as depicted in Figure 4(a)). Zhu et al. (Zhu et al., 2021) reverse-engineer the critical data structures (e.g., GPU command headers) offline and collect victim PCIe packets via a snooping device to reconstruct models in the online phase. The stolen DNN models have the same parameters and semantically identical architecture as the original ones. One potential mitigation for this attack is encrypting PCIe traffic. In contrast, Hu et al. (Hu et al., 2020) assume that the adversary cannot access data passing through buses, only the addresses, allowing attacks even with encrypted data. They conduct bus snooping attacks on discrete GPU’s memory and PCIe bus, obtaining kernel read/write access volume and memory address traces. Correlating extracted architectural events with model internal architectures. Another class of attacks occurs between two concurrent GPU applications, as shown in Figure 4(b). In (Naghibijouybari et al., 2017), they (i) implement website fingerprinting attack via the GPU memory utilization Application Programming Interface (API) or GPU performance counters; (ii) monitor user’s web activities during interactions or while typing characters on a keyboard through keystroke timing analysis; (iii) utilize a CUDA spy within the computational stack of GPU to deduce the NN structure by leveraging shared resources. In contrast to (Naghibijouybari et al., 2017), Wei et al. (Wei et al., 2020) consider fine-grained time-sliced sharing of applications on different VMs and exploit context-switching penalties on performance counters as leakage vector to recover the layer composition and hyperparameters of a DNN.
Most of these attacks require extensive reverse-engineering of NVIDIA GPU details. This is becauseGPU driver and internal GPU workings are closed-source, and crucial data structures are left undocumented (cf. Section 2.1). In other words, these factors make the attack more difficult to some extent, but do not fundamentally prevent them.
Covert-channel attacks where both spy and victim concurrently run on a single GPU: The key point of a successful covert-channel attack is identifying the shared resources utilized by both the trojan and spy, often requiring extensive reverse engineering of GPU architectures. This process is necessary to uncover details about GPU components, such as the TLB and other shared resources, which can be exploited for covert communication. Naghibijouybari et al. (Naghibijouybari et al., 2017) first reverse-engineer the hardware block and warp-to-warp schedulers to establish the co-location, then learn contention on caches, functional units and memory, and finally construct covert channels based on these resources with an error-free bandwidth exceeding 4 Mbps. With this, they recover the number of neurons in an NN. Ahn et al. (Ahn et al., 2021b) initially reverse-engineer the structure of on-chip networks within modern GPUs and note that the hierarchical organization of the GPU leads to the sharing of interconnect bandwidth among adjacent cores. They then exploit the contention of the interconnect channel, leverage clock registers for sender-receiver synchronization and ultimately establish microarchitectural covert channels with a surprisingly high bandwidth of 24 Mbps. Nayak et al. (Nayak et al., 2021) reverse-engineer the details of the TLB hierarchy and the details of all TLB levels, such as the number of entries and associativity, of the NVIDIA 1080Ti. They create a covert timing channel via the shared L3 TLB using Prime+Probe. With MPS, they increase the bandwidth of the GPU’s L3 TLB-based covert channel to 81 Kbps. Unlike previous work relies on access to the native GPU APIs through CUDA or OpenGL for high-precision performance counters monitoring, Giner et al. (Giner et al., 2024) do not rely on WebGPU providing hardware timers. Instead, they use a shared memory buffer and a dedicated thread to increment a shared variable constantly, functioning as a timer. With this timer and their GPU-accelerated cache eviction set construction, they establish a Prime+Probe cache covert channel for the L2 cache with a bandwidth of 10.9 KB/s.
Attacks on multiple GPUs. In NVIDIA’s DGX machines, where discrete GPUs are connected via NVLink and PCIe, Dutta et al. (Dutta et al., 2023) reverse-engineer the cache hierarchy and show that an attacker on one GPU can create contention in the L2 cache of another GPU through NVLink. They establish a Prime+Probe covert-channel across two GPUs, achieving a bandwidth of 4 MB/s. Additionally, their side-channel attack allows an attacker to monitor another application running on different GPUs, enabling the recovery of the number of neurons of the NN. By reverse-engineering the TLB structure in recent NVIDIA GPUs such as A100 and H100, Zhang et al. (Zhang et al., 2023) discover a design flaw in the NVIDIA MIG feature: MIG does not partition the last-level TLB, which is shared by all GPU instances. They dump GPU memory to locate corresponding table entries and access the target entry through the 1 MB GPU memory window (MMIO) for runtime modifications. They finally establish a cross-GPU-instance covert-channel based on the shared TLB. In summary, both studies begin by identifying shared resources present in multi-GPU systems, such as the shared remote L2 cache (Dutta et al., 2023) and the shared L3 TLB (Zhang et al., 2023), and then create visible contention on these shared resources.
5.2.3. Countermeasures
The countermeasures can be summarized as follows:
(i) Partitioning resources: Co-location between the attacker and the victim is essential for those attacks that exploit shared microarchitectural resources. One solution to mitigate this threat is to partition resources (e.g., LLC) either statically or dynamically (Wang and Lee, 2007; Kong et al., 2009; Liu et al., 2016). By assigning the spy and trojan to different cache partitions, they cannot replace each other’s cache lines. The MIG introduced in Section 2.1 potentially mitigates these attacks by isolating resources between independent GPU instances. However, Zhang et al. (Zhang et al., 2023) point out that the last-level TLB between MIG-enabled GPU instances is not properly partitioned, allowing for the creation of a covert channel using this vulnerability.
(ii) Eliminating contention: Contentions among processes can be eliminated by controlling traffic in memory controllers. This involves grouping memory requests from each processor into the same queue, potentially accessing the same memory bank/port (Wang et al., 2014; Shafiee et al., 2015). For example, TLB-pilot (Di et al., 2021) binds thread blocks from different kernels to different groups of streaming multiprocessors by considering hardware isolation of last-level TLB and the application’s resource needs. It gathers kernel information before kernel launches and reduces load imbalance by coordinating software and hardware-based scheduling and employing a kernel splitting strategy. There are also defences against coalescing-based correlation timing attacks. For example, Kadam et al. (Kadam et al., 2018, 2020) add extra memory accesses to make execution time less predictable. To protect against timing and cache-based attacks in GPU-based AES, Lin et al.(Lin et al., 2019) use scatter and gather techniques to reorganize AES tables, avoiding key-dependent lookups from timing or address leakage.
5.3. Physical Side-Channel Attacks
We classify attacks allowing an attacker to have direct proximity or contact with the GPU or its surrounding hardware components as physical access-based attacks. Such an attacker is physically close to the device and can directly interact with it. This may involve placing sensors or monitoring equipment directly on or near the device to capture physical signals or emissions. Note that we term attacks that perform monitoring or manipulation of the physical properties of the target device from a distance, without direct physical access to the device itself, as remote access-based attacks (called hybrid attacks in (Taneja et al., 2023)). Such attacks use e.g., wireless sensors or network-connected devices to capture signals or emissions emitted by the target device from a distance.
5.3.1. Physical Access-based.
Luo et al. (Luo et al., 2015) and Gao et al. (Gao et al., 2018) focus on recovering the AES private key by exploiting physical properties. Luo et al. utilize correlation power analysis to analyze the GPU’s power consumption (i.e., power traces), necessitating the attacker to instrument the power supply. Gao et al. sample EM traces employing a precise trigger mechanism and establish a heuristic leakage model to exploit the concurrent MMIO leakages in parallel scenarios. This method requires removing the GPU’s heat sink and positioning a probe near the GPU chip surface.
Hu et al. (Hu et al., 2020) employ MMIO side-channel attacks to capture memory access volume and kernel execution time by targeting CPU-GPU interconnections or co-locating a CUDA spy. This allows them to deduce the model architecture without any prior knowledge of the NN model. Similarly, Chmielewski and Weissbart (Chmielewski and Weissbart, 2021) use simple MMIO analysis to recover partial information of an NN model. Maia et al. (Maia et al., 2022) show how to recover the full model architectures of 64 NNs by exploiting the magnetic flux emanating from a GPU’s power cable via a physical sensor. Horvath et al. (Horvath et al., 2023) use correlation EM analysis to recover the weights and biases in the convolutional layers of a CNN model.
Taneja et al. (Taneja et al., 2023) capture side-channel leakage of frequency, power, and temperature (monitored via internal sensors) in response to the workload executed in both integrated and discrete GPUs. Their attacks, which observe instruction and data-dependent behaviour, require different privilege levels depending on vendors. For instance, Apple permits unprivileged access to both, whereas Google restricts temperature readings but allows unprivileged access to power.
5.3.2. Remote Access-based.
Zhan et al.(Zhan et al., 2022) identify the DVFS feature used in GPU as the underlying cause of an MMIO side-channel vulnerability. In scenarios where a victim employs an NVIDIA or AMD GPU without sharing, an attacker can monitor the victim’s activities and discern the visited webpages (i.e., a website fingerprinting attack). The key idea is capturing the fluctuating workload of a GPU during webpage rendering, which influences the MMIO signal it emits. These workload variations, coupled with the need to manage heat generation, fan noise, and minimize power consumption, lead to changes in the GPU’s performance levels (e.g., P-states in NVIDIA GPUs). Consequently, when a performance level switches on or off, distinct MMIO signals corresponding to the GPU’s memory clock frequencies appear or vanish in the spectrum. Remarkably, this attack can intercept MMIO signal leakage at distances of up to 6 meters, even through a wall. Liang et al. (Liang et al., 2022) conduct a similar far-field MMIO side-channel attack to infer the number of layers and their types, the number of kernels, sizes of layers and strides in an NN model.
5.3.3. Countermeasures
MMIO side channels from GPUs can be exploited when these EM signals are dependent on computations, revealing information about ongoing activities. Such side-channel signals are easily measurable when they are strong and can propagate over several meters. Two countermeasures can mitigate such vulnerabilities.
(i) Hiding and masking (Mangard et al., 2008): Hiding aims to reduce the Signal to Noise Ratio (SNR) available to attackers. Generating MMIO noise is possible, but shielding the computer to decrease the intensity of emitted MMIO signals is more effective. Hardware manufacturers could also focus on minimizing MMIO emissions in future product designs. Masking involves combining data with random mask values to ensure the observable leakage is independent of the secret. For example, creating random GPU workloads during sensitive tasks can disrupt usage patterns and reduce predictability (Mahmoud et al., 2022; Zhan et al., 2022). However, masking can be complicated due to the need for random masks and the added complexity of the algorithms.
(ii) Coarsening the granularity: Another approach is to make GPU DVFS less sensitive to workload changes. This would keep performance more stable, leading to less detailed leaked information and reducing the chance of sensitive data being inferred (Zhan et al., 2022). However, such changes might negatively affect the efficiency benefits of DVFS.
5.4. Fault Injection Attacks
5.4.1. Software-based.
Frigo et al. (Frigo et al., 2018) use GPU-based timers to execute a JavaScript-based side-channel attack, enabling attackers to identify contiguous regions of physical memory. This method allows for remote triggering of Rowhammer bit flips in the identified memory areas. Consequently, an attacker can breach Address Space Layout Randomization (ASLR) and evade the Firefox sandbox on Android platforms featuring integrated Qualcomm Adreno GPUs. Sabbagh et al. (Sabbagh et al., 2020) introduced the first non-invasive fault attack named “overdrive” on a GPU. By exploiting the DVFS interface, an adversary can configure out-of-specification voltage and frequency combinations on the host CPU and send malicious overdrive commands to the GPU control registers, inducing random faults during GPU kernel execution. They also perform a misclassification attack on victim CNN inference (Sabbagh et al., 2021), improving the overdrive attack by addressing its limitations in timing precision and the lack of analysis on how fault injection affects silent data corruption.
However, unlike encryption algorithms, where hardware fault injection attacks are highly effective, ML models, especially well-trained commercial DNN models, are inherently more fault-tolerant. This is because computation errors from random hardware faults are usually absorbed by later layers and do not greatly impact accuracy. Therefore, inducing hardware faults at specific timing points, optimizing the glitch voltage-frequency pair, and determining the appropriate fault duration are critical for generating controllable and impactful faults in ML models. To this end, Sun et al. (Sun et al., 2023) focus on searching sensitive targets of a CNN model, identifying proper parameters for timing and position, refining fault injection parameters, and ultimately reducing inference accuracy by an average of $69.1\%$. Santos et al. (Santos et al., 2021) propose a two-level fault injection framework combining the accuracy of Register Transfer Level (RTL) fault injection with the efficiency of software fault injection.
Overall, researchers have primarily leveraged software interfaces to inject faults, often utilizing DVFS or Rowhammer attacks. Rowhammer attacks exploit the DRAM vulnerability where repeated accesses to a memory row (  i.e. hammering) can cause bit flips in adjacent rows, enabling dangerous exploits like privilege escalation from a normal user to a system administrator. The high parallelism and memory bandwidth inherent in GPUs potentially make them more vulnerable to Rowhammer attacks, as they can generate memory accesses faster than CPUs.
5.4.2. Countermeasures
Protections against Rowhammer exploits are most effective when implemented in the main memory or the memory controller. Frigo et al.(Frigo et al., 2018) investigate preventing attackers from targeting valuable data by enforcing stricter memory reuse policies. One solution involves enhancing the physical compartmentalization suggested by CATT (Brasser et al., 2017a) to user-space applications. However, this approach brings trade-offs in complexity, performance, and capacity, since dynamically allocating (isolated) pages increases complexity and can have performance implications.
Regarding DVFS-based attacks, it is feasible to make it more difficult for adversaries to control and exploit the interfaces. Voltage and clock signal monitors can be used to detect unconventional changes111cf. e.g., Intel’s approach: https://www.intel.com/content/www/us/en/newsroom/news/the-story-behind-new-intel-security-feature.html. Additionally, some vendors have disabled all software-accessible interfaces for DVFS. For example, Intel defends against DVFS-based attacks on Intel SGX (Murdock et al., 2020) with a BIOS setting to disable the overclocking mailbox interface and with a microcode update that includes the current state of this setting in the SGX TCB attestation. A similar strategy could also be applied to GPUs.
6. Security of CPU TEEs
In this section, we introduce typical attacks against CPU TEEs. As there already exist survey papers focusing on reviewing existing attacks on CPU TEEs, and the aim of this survey is to explore potential attacks on GPU TEEs, we simply categorize the most common attacks based on their technical details and existing surveys (Fei et al., 2021; Nilsson et al., 2020; Cerdeira et al., 2020; Li, 2022) here.
6.1. Security of Intel SGX
6.1.1. Controlled Channel and Address Translation-based Attacks
Controlled channels attack, a concept introduced by Xu et al. (Xu et al., 2015) in 2015, exploits the privileges of the OS to monitor and manipulate page tables. This allows the attacker to uncover the memory access patterns of the enclave code, cf. also (Shinde et al., 2016). This attack is especially effective because, while Intel SGX enclaves protect sensitive computations from a compromised OS, the OS still manages memory and can observe page faults, thereby inferring details about the enclave’s operations. The SGX-Step (Van Bulck et al., 2017a) serves as a practical tool for, among others, executing such controlled channel attacks by configuring Advanced Programmable Interrupt Controller (APIC) timers, issuing interrupts, and tracking page-table entries. SGX-Step enables deterministic single-stepping of enclave execution, allowing for precise monitoring of memory access patterns over time.
In fault-less page table attacks (Kim et al., 2019; Van Bulck et al., 2017b; Wang et al., 2017), attackers exploit the page table itself rather than relying on page faults. This allows them to infer the target’s memory access patterns by examining page table attributes and/or observing the caching behaviour of unprotected page table memory. However, these attacks are limited to page-level accuracy. Segmentation attacks (Gyselinck et al., 2018) allow attackers to deduce finer memory access patterns by manipulating the segmentation unit, but this only works on 32-bit enclaves, as segmentation is disabled in 64-bit enclaves (Fei et al., 2021).
6.1.2. Cache-based Attacks
Typical cache timing-channel attacks (in the non-SGX scenario) comprise four primary variants, including Evict+Time (Tromer et al., 2010), Evict+Reload (Gruss et al., 2015), Prime+Probe (Osvik et al., 2006), and Flush+Reload (Yarom and Falkner, 2014), all relying on the time difference between cache hits and misses. Recent studies (Götzfried et al., 2017; Moghimi et al., 2017; Schwarz et al., 2017; Brasser et al., 2017b; Dall et al., 2018; Moghimi et al., 2019) show that SGX enclaves are susceptible to the same cache attacks as any other software application. In fact, they may be even more vulnerable due to the enhanced attackers within SGX’s threat model. Several papers (Götzfried et al., 2017; Moghimi et al., 2017; Schwarz et al., 2017; Brasser et al., 2017b) exploit L1 cache and the shared LLC as the attack surface by employing Prime+Probe method. Dall et al. (Dall et al., 2018) take a slightly different approach by using Prime+Probe on Intel’s provisioning enclave, allowing Intel themselves to compromise the Enhanced Privacy ID (EPID) unlinkability property. Memjam (Moghimi et al., 2019) utilizes read-after-write false dependencies caused by the 4K aliasing of the L1 cache to conduct an Evict+Time-style attack.
6.1.3. Branch Prediction Attacks
Branch prediction attacks offer a more detailed level of utilization compared to address translation-based and cache-based attacks. They target the branch prediction unit, including the Branch Target Buffer (BTB(Lee et al., 2017; Chen et al., 2019), the Page History Table (PHT(Evtyushkin et al., 2018; Huo et al., 2020), and the Return Stack Buffer (RSB(Koruyeh et al., 2018). Attackers exploit the BTB to ascertain whether a particular branch instruction is taken by the victim. The PHT attack seeks to provoke collisions between branches of victim and attacker processes and subsequently infer the direction of the victim process branch. RSB attacks rely on contaminating the RSB to incite mis-speculation, consequently resulting in the leakage.
6.1.4. Transient Execution Attacks
Transient execution attacks, which exploit out-of-order and speculative execution of instructions (summarized in section 2.2), have also been studied in the context of SGX. The aforementioned BTB and RSB attacks (Chen et al., 2019; Koruyeh et al., 2018) can also fall into this class of attacks, stemming from Spectre (Kocher et al., 2020). Additionally, the Foreshadow attack (Van Bulck et al., 2018), which targets SGX by employing Meltdown-like techniques, breaches enclave execution confidentiality. While MDS attacks (Canella et al., 2019; Van Schaik et al., 2019; Schwarz et al., 2019) also leverage speculative and out-of-order execution, they primarily exploit information leakage from various implementation-specific and undocumented intermediary buffers of the targeted microarchitecture. SGAxe (Van Schaik et al., 2020) and Crosstalk (Ragab et al., 2021) extend MDS to extract private keys from SGX enclaves.
6.1.5. Software-based Fault Injection and Side-Channel Attacks
In fault injection attacks, the attacker disrupts the normal operation of a system by introducing malicious faults and then derives secret information from the faulty execution. Software-based fault attacks expand the threat model from a local attacker, who requires physical access to the target device, to a potentially remote attacker with only local code execution capabilities. Several studies (Murdock et al., 2020; Kenjar et al., 2020; Qiu et al., 2020) exploit privileged DVFS on x86 CPUs to reliably corrupt enclave computations, thereby compromising the confidentiality and integrity of enclaves. Chen and Oswald (Chen and Oswald, 2023) show that software fault attacks on SGX can also be mounted by pivoting to management chips on Supermicro server systems, and can even permanently damage the host CPUs. Apart from voltage scaling-based attacks, SGX-Bomb (Jang et al., 2017) leverages the Rowhammer bug to induce bit flips in enclave memory, resulting in failed data integrity checks and subsequent system lock-down.
In addition to these fault injection attacks, Lipp et al. (Lipp et al., 2021) exploits a side-channel attack by leveraging unprivileged access to the Intel Running Average Power Limits interface. This interface reveals values linked to power consumption, creating a power side channel exploitable across Intel server, desktop, and laptop CPUs. Wang et al. (Wang et al., 2022b) show that power leakage can be measured through indirect channels such as frequency scaling, and Liu et al. (Liu et al., 2022) use the related feature of throttling to among others leak secrets from SGX enclaves.
6.1.6. Hardware-based Attacks
Apart from software-based fault injection attacks, Chen et al. (Chen et al., 2021; Chen and Oswald, 2023) explore hardware-level fault injection attacks on SGX enclaves. “VoltPillager” is a low-cost tool that injects messages on the Serial Voltage Identification (SVID) bus between the CPU and the voltage regulator, breaching the confidentiality and integrity of Intel SGX enclaves (Chen et al., 2021). Chen and Oswald later demonstrates that undervolting via the Power Management Bus (PMBus) can similarly break SGX enclave integrity, bypassing Intel’s countermeasures against previous software undervolting attacks (Murdock et al., 2020; Kenjar et al., 2020; Qiu et al., 2020). “MEMBUSTER” (Lee et al., 2020) introduces an off-chip attack that breaks SGX enclave confidentiality by snooping the memory bus (with high-spec, expensive lab equipment), extracting memory access patterns and increasing cache misses using techniques like critical page whitelisting and cache squeezing.
6.1.7. Interface-based Attacks
SGX enclave security can be compromised if they fail to strictly adhere to secure interfaces between trusted and untrusted code. Despite well-designed open-source Software Development Kits, Van Bulck et al. (Van Bulck et al., 2019) identify several sanitization vulnerabilities in Application Binary Interface (ABI) and API, leading to memory safety and side-channel vulnerabilities. Alder et al. (Alder et al., 2020) explores the attack surface affecting floating-point computations in SGX enclaves via ABI, showing that control and state registers are not always properly sanitized. Recently, there has been a focus on symbolic execution tools for SGX (Cloosters et al., 2020; Khandaker et al., 2020; Alder et al., 2024). Pandora (Alder et al., 2024) enables runtime-agnostic symbolic execution of the exact attested enclave binary and validates the enclave shielding runtime. Pandora can autonomously discover 200 new and 69 known vulnerable code locations across 11 different SGX shielding runtimes.
6.1.8. Countermeasures
In this section, we review and categorize the most relevant countermeasures based on both prior surveys (Nilsson et al., 2020; Fei et al., 2021) and potential solutions proposed in our summarized attack papers.
(i) System/Microcode-level countermeasures: A CPU’s complex instructions are often managed by low-level software called microcode, rather than being entirely implemented in hardware. Thus, Intel can issue microcode updates, which are typically the most effective method. For instance, in response to DVFS-based attacks like Plundervolt (Murdock et al., 2020), Intel released microcode updates that fully disable the software voltage scaling interface via MSR $\mathtt{0x150}$.
(ii) Compiler-based countermeasures: Beyond Intel’s solutions, compiler-based approaches are also effective. Deterministic multiplexing (Shinde et al., 2016) ensures that page-fault access patterns remain consistent regardless of input values by proactively accessing all data and code in a predetermined sequence. “T-SGX” (Shih et al., 2017) use Intel’s Transactional Synchronization Extensions (TSX) to redirect exceptions and interrupts to a specific page, terminating the program if a page fault is detected. Chen et al. (Chen et al., 2017) propose shielded execution with reliable time measurement to check the execution time of target programs. “Cloak” (Gruss et al., 2017) uses Hardware Transactional Memory (HTM) to protect against malicious cache observations. Similarly, glitch detection (Spensky et al., 2021) can help mitigate DVFS-based fault injection attacks.
(iii) Randomization: “SGX-Shield” (Seo et al., 2017) leverages ASLR to create a secure in-enclave loader that secretly randomizes memory space layout, hiding sensitive operations from attackers. Hosseinzadeh et al. (Hosseinzadeh et al., 2018) propose runtime control flow randomization, implemented as compiler extensions on llvm. “DR.SGX” (Brasser et al., 2019a) disrupts memory observations by permuting data locations.
(iv) Application/Source code design: Finally, developers can design enclave applications in a side-channel-resistant manner. Early works like “OBFUSCURO” (Ahmad et al., 2019) enforce operations with Oblivious RAM (ORAM) protection to prevent attackers from inferring access patterns. Several schemes (Ohrimenko et al., 2016; Poddar et al., 2020; Wang et al., 2022a) utilize oblivious primitives, implemented in x86_64 assembly, to design data-oblivious applications. These primitives operate exclusively on registers, which are private to the processor, ensuring that the contents hidden from code outside the enclave and any register-to-register data manipulation is data-oblivious.
6.2. Security of Arm TrustZone
6.2.1. Implementation Bugs
Cerdeira et al. (Cerdeira et al., 2020) examine bug reports obtained from public Common Vulnerabilities and Exposures (CVE) databases and vendor bulletin reports, categorizing them into validation bugs, functional bugs, and extrinsic bugs. These bugs are pervasive and often exploited to escalate privileges, enabling attackers to fully hijack kernels in devices from Qualcomm (Beniamini, 2016) and Huawei (Shen, 2015) or compromise client applications like Samsung Pay (Komaromy, 2018).
6.2.2. Cache-based Attacks
Given that the cache is shared between NW and SW in TrustZone-enabled processors, several studies capitalize on the contention arising from this cache coherency to extract sensitive information (Zhang et al., 2016c; Guanciale et al., 2016; Lipp et al., 2016; Zhang et al., 2016b; Cho et al., 2018). Several authors (Lipp et al., 2016; Zhang et al., 2016b; Cho et al., 2018) use the Prime+Probe method to monitor the cache activities within TrustZone.
6.2.3. Branch Prediction Attacks
Similar to the attacks on Intel SGX mentioned earlier, the branch predictor can also be exploited to target TrustZone. Since the BTB is shared between NW and SW, techniques like Prime+Probe can be used to disclose secure information to NW. For example, Ryan (Ryan, 2019) successfully recovered a 256-bit private key from Qualcomm’s hardware-backed keystore.
6.2.4. Software-based Fault Injection Attacks
Several studies have also expanded software-based fault injection attacks to TrustZone or TrustZone-based systems. “CLKSCREW” (Tang et al., 2017) leverages manipulation on DVFS to induce faulty computations to breach TrustZone hardware-enforced boundaries, extract secret keys, and bypass code signing operations. Rowhammer-based attacks have also been adapted to undermine TrustZone security. Given that overclocking typically involves running the processor at high frequencies, which can be relatively easy to detect and prevent (e.g., through hardware frequency locking), “VoltJockey” (Qiu et al., 2020) adopts a different strategy by manipulating voltages rather than frequencies to generate hardware faults on the victim cores. Carru (Carru, 2017) use bitflips to bypass security measures and access secure memory, enabling the extraction of RSA keys and undermining the secure storage provided by TrustZone.
6.2.5. Countermeasures
In this section, we review and categorize countermeasures based on both previous survey (Cerdeira et al., 2020) and potential solutions from the summarized attack papers.
In addressing the architectural defences of Arm TrustZone, some research efforts (Sun et al., 2015; Brasser et al., 2019b) enhance isolation granularity between TEE components and reduce the amount of code running in SW, thereby lowering the risk of severe privilege escalation attacks within SW. Countermeasures (Jang et al., 2015; Jang and Kang, 2018) tackle the weaknesses of inadequate/weak authentication when accessing TEE resources from NW, and potentially insecure shared memory used for data exchange across the NW-SW boundary. Unlike Intel SGX, TrustZone lacks built-in on-chip memory encryption, prompting some works (Zhang et al., 2016a; Yun and Zhong, 2019) to design encryption strategies to bridge this gap. To further strengthen the verification of both TEE and TA binaries’ integrity and identity, several studies (Ferraiuolo et al., 2017; Lentz et al., 2018) introduced additional trusted computing primitives, such as enhanced remote attestation and sealed storage. To address implementation bugs, “RustZone” (Evenchick, 2018) extends TrustZone by implementing TAs using Rust, which inherently provides memory safety and thread safety.
Cache side-channel attacks can be mitigated by carefully implementing cryptographic algorithms in software (Guanciale et al., 2016) or using dedicated hardware (Lipp et al., 2016). Cache maintenance techniques can also be effective in mitigating leakage. For example, in terms of BTB attacks, flushing the shared microarchitectural structures when transferring between NW and SW can prevent attacks from targeting keys through the BTB. However, as observed in the Nexus 5X firmware by (Ryan, 2019), the L1 cache is not flushed, making Prime+Probe-based attacks still feasible.
Countermeasures against DVFS-based fault injection attacks are similar to those used for SGX. In comparison to other attacks on TrustZone, Rowhammer attacks are relatively straightforward to mitigate. They rely on inducing bit flips in adjacent memory rows, making them feasible only if non-secure memory is situated next to a memory row within SW. If secure memory resides on a separate memory storage device, Rowhammer attacks become impractical.
6.3. Security of AMD SEV
6.3.1. Memory Encryption Issues
Since the release of AMD SEV (AMD, 2023), numerous attacks have targeted its early memory security mechanisms (Hetzelt and Buhren, 2017; Buhren et al., 2017; Du et al., 2017; Morbitzer et al., 2018). Hetzelt and Buhren (Hetzelt and Buhren, 2017) utilize a return-oriented programming-like attack to exploit an unencrypted VM control block, enabling arbitrary reading and writing of encrypted memory in the guest VM. Such vulnerabilities have been addressed in SEV-ES (Kaplan, 2017). Du et al. (Du et al., 2017) leverage Electronic Code Book (ECB) mode in memory encryption to conduct a chosen-plaintext attack via an HTTP server installed on the guest VM. Sevurity (Wilke et al., 2020) exploits vulnerabilities in the Xor-Encrypt-Xor (XEX) mode to insert arbitrary 2-byte instructions into encrypted memory, while Morbitzer et al. (Morbitzer et al., 2018) breach confidentiality by manipulating nested page tables, altering the virtual memory of the guest VMs. Most of these attacks arise from the lack of integrity protection, but SEV-SNP (Sev-Snp, 2020) addressed them through the Reverse Map Table. Conversely, “CIPHERLEAKS” (Li et al., 2021b) allows a privileged adversary to infer guest VM execution states or recover certain plaintext by monitoring changes in the ciphertext of the victim VM. The authors highlight that AMD has confirmed that SEV-SNP is also vulnerable to the CIPHERLEAKS attack (Li, 2022).
Additionally, “CROSSLINE” and its variants by Li et al. (Li et al., 2021a; Li, 2022) extract the memory content of victim VMs by exploiting improper use of the Address Space Identifier (ASID) for controlling VM accesses to encrypted memory pages, cache lines, and TLB entries. They also compromise the integrity and confidentiality of SEV VMs by poisoning TLB entries (Li et al., 2021c). It is noted that SEV-SNP is expected to address these issues, including TLB misuse.
6.3.2. Software-based Fault Injection Attacks
Morbitzer et al. (Morbitzer et al., 2021) exploit the system’s lack of memory integrity protection to inject faults and execute arbitrary code within SEV VMs through I/O channels, using the hypervisor to locate and trigger the execution of encrypted payloads. However, SEV-SNP (Sev-Snp, 2020) can mitigate this issue through Reverse Map Table (RMP). Buhren et al. (Buhren et al., 2021) present a voltage fault injection attack enabling the execution of custom payloads on AMD secure processors, subsequently decrypting a VM’s memory by deploying custom SEV firmware. “CacheWarp” (Zhang et al., 2024) presents a new software-based fault attack targeting AMD SEV-ES and SEV-SNP, exploiting the possibility to revert modified cache lines of guest VMs to their previous (stale) state at the architectural level. This attack only requires interrupting the VM at a point chosen by the attacker to invalidate modified cache lines without writing them back to memory. More recently, Schlüter et al. (Schlüter et al., 2024a) introduce a novel attack where the hypervisor injects well-crafted, malicious #$\mathtt{\#VC}$ (a new exception by AMD facilitating communication between the VM and untrusted hypervisor) into a victim VM’s CPU, enabling the attacker to induce arbitrary behaviour in the VM. In another related attack, “HECKLER” (Schlüter et al., 2024b) shows that the hypervisor can inject malicious non-timer interrupts to compromise VMs. They leverage interrupt handlers with global effects, allowing manipulation of a VM’s register states to alter data and control flow. In general, SEV-SNP remains vulnerable to such software-based fault injection attacks.
6.3.3. Interrupt-driven Attacks
Similar to SGX-Step framework (Van Bulck et al., 2017a), SEV-Step (Wilke et al., 2024) introduces reliable single-stepping against SEV VMs. In addition to this functionality, SEV-Step facilitates access to common attack techniques, such as page fault tracking and cache attacks targeting SEV.
6.3.4. Countermeasures
Most memory-related attacks have been addressed in recent AMD product updates. Here, we focus on countermeasures against fault injection attacks. For malicious voltage drops or glitches, similar to countermeasures used for SGX and TrustZone, hardware-level voltage monitoring circuits and software-based detection mechanisms can be implemented. To counter fault attacks that enable fine-grained memory write suppression through specific hypervisor instructions, restricting the use of those instructions at the hardware level or implementing compiler-level solutions can ensure correct memory operations. For attacks that inject malicious interrupts, one can block or detect the injections by monitoring external interrupts or disabling the interrupt handler.
7. Conclusions on future directions in GPU TEE
In this section, based on the relevant work discussed before, we sketch potential attacks that might be relevant to GPU TEEs in the future. We note that as GPU TEEs are not widely commercially available, it is not yet possible to practically evaluate these attack vectors, but believe that according vulnerabilities might be discovered in the future.
7.1. Potential Attacks against x86-based GPU TEE Prototype
Figure 5. Potential attacks against x86-based GPU TEE prototype.
As summarized in Section 4, the state-of-the-art designs targeting x86 architecture each have unique characteristics aimed at reducing the TCB size and minimizing hardware modifications. The primary goal of these designs is to protect GPU resources from a compromised OS, including the GPU driver. We outline the key design ideas and provide a general overview in Figure 5. Specifically, in this threat model, the adversary controls the entire software stack, encompassing OS, hypervisor, and device drivers. The adversary also aims to compromise the hardware and software I/O data path between a user application and the GPU. The functionalities of GPU drivers are often partitioned into secure and non-secure parts instead of porting the entire driver into the TEE enclave, which would increase the TCB size. Given that the adversary can compromise the OS, the GPU driver located in the OS is also within this scope. The device memory of the GPU is trusted and the adversary cannot observe or corrupt the data stored in it. Moreover, we recognize that the power and frequency management modules residing in both the CPU and GPU sides are outside the scope of protection. Next, based on this security model, we will analyze potential attacks that could pose threats and leak information when such designs are deployed in real-life applications.
7.1.1. Existing Attacks against Intel SGX
Given that the essential GPU resource management functions within the GPU driver reside within the TEE enclave, the majority of attacks outlined in Section 6.1 could potentially target this enclave to glean secrets (cf. attack surface ➀ in Figure 5). It is worth noting that while some research may assert that most side-channel attacks against SGX are out of scope, it is critical to provide the corresponding protection when considering the deployment of GPU TEE designs in real-world cloud platforms (where even some forms of physical access by malicious insiders might be relevant). These potential leakages can be categorized as follows: (i) Observing memory access patterns allows attackers to deduce the control flow of programs within the SGX enclave, including decision-making processes. Armed with this insight, attackers could monitor GPU commands and data transfers, thereby inferring the tasks executed on the GPU. For instance, it might become feasible to infer the quantity (and size) of detected objects in each frame of a CNN-based video analysis task (Poddar et al., 2020). (ii) Monitoring page fault patterns enables attackers to extract cryptographic keys from implementations of certain cryptographic open-source software libraries (e.g., OpenSSL and Libgcrypt) (Shinde et al., 2016). With this knowledge, attackers could subsequently access all sensitive information transmitted to the GPU. (iii) Exploiting timing channels allows attackers to infer sensitive information about kernels running on the GPU. For instance, Telekine (Hunt et al., 2020) demonstrates how attackers can accurately classify images from ImageNet by solely observing the timing of GPU kernel execution, without needing access to the images themselves.
7.1.2. DVFS-based Attacks
To enhance energy efficiency, many current CPU processors enable the DVFS extension, which adjusts the frequency and voltage of processor cores based on real-time computational loads. This capability now extends to accelerators like NVIDIA GPUs. As demonstrated by Lightning (Sun et al., 2023), NVIDIA GPU working voltages can be altered through software instructions executed on the CPU via NVIDIA drivers, potentially enabling attackers to induce transient voltage glitches for fault injection attacks. Furthermore, current NVIDIA GPUs offer unrestricted control for Phase Locked Loop (PLL) and clock divider, allowing easy overclocking by adjusting the frequency offset using NVIDIA, thus potentially facilitating frequency-based fault injection attacks. Consequently, for a x86-based GPU TEE prototype, we identify two attack surfaces (cf. ➁ and ➅) regarding DVFS-based attacks in Figure 5.
On the CPU side. In many existing GPU TEEs prototypes (Volos et al., 2018; Jang et al., 2019; Zhu et al., 2020; Lee et al., 2022; Yudha et al., 2022; Ivanov et al., 2023), the part of the GPU driver responsible for DVFS parameter control lies outside their TCB. Consequently, an attacker can flexibly adjust GPU voltage and frequency from the CPU side to induce faults on the GPU side. Additionally, it warrants investigation whether previous attacks against SGX could gain access to the relevant memory region to alter DVFS parameters, even if the GPU driver managing these parameters is protected by running in an SGX enclave.
On the GPU side. Due to the absence of protection on GPU hardware, attackers with physical access to GPU hardware could directly manipulate GPU voltage, injecting faults into running GPU kernels. For instance, “VoltPillager” (Chen et al., 2021) can inject messages on the SVID bus between the CPU and the voltage regulator on the motherboard, enabling precise control over CPU core voltage. Ultimately, this jeopardizes the confidentiality and integrity of Intel SGX enclaves. Exploring whether attackers can inject malicious SVID packets (or similar) into the respective bus on the GPU card presents an intriguing avenue for further investigation.
7.1.3. Physical Attacks on PCIe Bus and GPU Memory
In many existing GPU TEEs prototypes lacking vendor hardware support (cf. attack surfaces ➂ and ➄ in Figure 5), GPUs lack a trusted memory region, leaving data in GPU memory exposed in plaintext. Consequently, direct physical access to GPU memory can compromise sensitive data. Further, cold boot attacks (Halderman et al., [n. d.]) could be relevant, if an adversary can ensure (e.g., by cooling) that TEE-protected memory contents are retained and can be readout after a reset of the GPU.
Additionally, as PCIe interconnects are exposed, injecting malicious PCIe packets via specialized hardware is feasible. Simply securing the routing path to the GPU is not sufficient to protect GPU control via MMIO or DMA (Jang et al., 2019). While most designs employ authenticated encryption to ensure data integrity during transfer, this alone is insufficient to protect data confidentiality and integrity. For instance, certain designs lack a hardware root of trust necessary for secure authentication and establishment of trusted communication channels, involving secure storage for cryptographic keys and secure execution environments for cryptographic operations. Furthermore, ensuring secure boot and firmware integrity is crucial to guarantee the correctness of the secure channel establishment process.
7.1.4. Architectural/Logical Attacks on GPU
Previous attacks on memory vulnerabilities (see Section 5.1) are also possible, as current GPU TEE designs often overlook memory clearing between different GPU contexts (which is also relevant across resets in the case of cold boot attacks mentioned above). Therefore, a secure memory management strategy, including memory initialization and clearance, is critical to the integrity and security of the GPU TEE workflow.
7.2. Potential Attacks against NVIDIA Hopper H100-based GPU TEE
Figure 6. Potential attacks against NVIDIA Hopper H100-based GPU TEE.
In addition to the state-of-the-art x86-based GPU TEE prototypes, NVIDIA’s Hopper H100 Tensor Core GPU, based on the Hopper architecture, recently introduces advanced features for confidential computing (only available to select users at present). According to NVIDIA’s whitepaper (NVIDIA, 2022a), the H100 relies on specific CPU TEEs, such as Intel TDX, AMD SEV-SNP, or Arm CCA, to enable confidential computing. Here, we provide a general system model of an NVIDIA Hopper H100-based GPU TEE based on the whitepaper in Figure 6. Specifically, on the CPU side, this GPU TEE operates within the standard threat model of VM-based TEEs (Schlüter et al., 2024a, b; Zhang et al., 2024). The untrusted hypervisor loads the VM image into memory and manages the initial configurations. Remote attestation measures the VM’s initial memory state before initiating the boot process. The software executing within the VM (including the guest OS, user applications, and trusted modules for TEEs) is part of the TCB. The specifications require the hypervisor to set up certain initial states (e.g., the number of vCPUs, supported hardware features, and memory size). The hardware verifies these configurations and only enters the VM if the setup is correct. Additionally, the hardware zeroes out certain values (e.g., specific general-purpose registers) before exiting the VM and manages context saving and restoring during VM context switches. On the GPU side, the H100 GPU is expected to feature a memory encryption engine (most likely an Memory Encryption Engine (MEE(Corporation, 2015) similar to SGX) and other capabilities to ensure the confidentiality and integrity of data and code, even defending against basic physical attacks targeting the PCIe bus and GPU memory. However, the whitepaper does not provide detailed technical information about these features.
7.2.1. Existing Attacks against VM-based TEEs
Perhaps because VM-based TEEs have only been proposed in recent years, there is currently limited research on attacks against them. However, recent developments have shown that software-based fault injection attacks can be highly effective against AMD SEV-SNP. These attacks would allow an adversary to compromise the GPU driver within the VM, enabling precise control over sensitive tasks running on the GPU (cf. attack surface ➀ in Figure 6). For example, the “WeSee” attack (Schlüter et al., 2024a) targets AMD SEV-SNP used in NVIDIA H100-based GPU TEEs, allowing an attacker to perform arbitrary memory writes by faking a #$\mathtt{\#VC}$ for MMIO reads. This capability would enabls the attacker to modify commands and data sent to the GPU. More critically, attacks like CacheWarp (Zhang et al., 2024), which remain unmitigated in the latest SEV-SNP implementation, do not depend on the specifics of the guest VM. Therefore, it is crucial to equip NVIDIA H100-based GPU TEEs with protection against these attacks to preserve the confidentiality and integrity of GPU data. It is also important to explore potential attacks against other VM-based TEEs, such as Intel TDX, as these are integral to NVIDIA H100-based GPU TEEs.
7.2.2. DVFS-based Attacks
Similar to GPU TEE prototypes, we identify two attack surfaces (cf. ➀ and ➃) regarding DVFS-based attacks in Figure 6.
On the CPU side. Recall that existing attacks targeting VM-based TEEs (e.g., AMD SEV-SNP) allow attackers to escalate privileges to root and write any value of their choice to any location in the victim VM. Thus, attackers with these abilities might be able to manipulate the GPU’s voltage and frequency using software instructions performed on the CPU via the NVIDIA driver.
On the GPU side. The H100 whitepaper (NVIDIA, 2022b) lacks detail about voltage and frequency management. Assuming an attacker has physical access, they could potentially execute attacks like “VoltPillager” (Chen et al., 2021) to inject malicious SVID packets or equivalent into the relevant GPU card’s bus, thereby altering the GPU voltage.
In summary, an attacker with the described capabilities could modify the voltage and frequency from both the CPU and GPU sides. However, it remains uncertain whether and how these changes affect tasks running on the NVIDIA H100 GPU. Specifically, it is unclear if the induced faults significantly impact tasks, such as degrading ML model accuracy or fully extracting private encryption keys. At the time of writing this survey, the complete confidential workflow of the NVIDIA H100-based GPU TEE is unavailable, as the corresponding NVIDIA libraries are still under development.
7.2.3. Microarchitectural Side-Channel and Covert-Channel Attacks on GPU
The H100 whitepaper (NVIDIA, 2022b) emphasizes the enhanced security features of the Hopper architecture for confidential computing, particularly in collaborative multi-party computing scenarios, namely single tenant with multiple GPUs and multiple tenants with a single GPU. However, vulnerabilities discussed in Section 5.2 may potentially lead to security breaches on the NVIDIA H100 GPU.
[…]
Multiple tenants with a single GPU. In the case of multiple tenants with a single GPU (cf. attack surface ➂ in Figure 6), facilitated by the MIG feature, an NVIDIA H100 GPU can be partitioned into multiple GPU instances, each isolated in terms of memory system paths, including the on-chip crossbar ports, L2 cache banks, memory controllers, and DRAM address busses. However, as discussed in Section 5.2, Zhang et al. (Zhang et al., 2023) demonstrate a design flaw of the MIG feature where the last-level TLB is not securely partitioned by all GPU instances. Consequently, this flaw enables covert-channel attacks between different GPU instances. Given that this mode supports multiple tenants, it is pertinent to investigate whether each GPU instance created with MIG truly has separate and isolated paths through other microarchitectural components such as L1 cache, warp scheduler, and command processor.
7.3. Potential Attacks against Arm-based GPU TEE Prototype
Figure 7. Potential attacks against Arm-based GPU TEE.
Due to various architectural disparities in TEEs for Arm GPUs, several studies have explored Arm-based GPU TEEs. In Figure 7, we base it on the design proposed by StrongBox (Deng et al., 2022). StrongBox serves as a suitable foundation as it was recently introduced and does not rely on specific Arm endpoint features or necessitate hardware modifications to GPU or CPU chips. Notably, while secure GPU computing on Arm or Intel platforms may not align with Arm’s CCA realm-style architecture (Wang et al., 2024b), we omit insights regarding this, as it closely resembles that of NVIDIA Hopper H100-based GPU TEE. In Figure 7, the attacker gains control over the kernel and the entire GPU software stack, including the GPU driver and runtime. This control enables them to manipulate sensitive data and code within GPU applications. They can directly access unified memory used for GPU tasks or manipulate peripherals to evade detection. Furthermore, by compromising the GPU driver, the attacker can undermine the memory management of GPU applications, potentially mapping sensitive data to unprotected regions. Similar to existing x86-based GPU TEEs, trust is placed in the GPU and the secure world. Based on this threat model, we discuss potential attacks in the following.
7.3.1. Existing Attacks against Arm TrustZone
The attacks summarized in Section 6.2 that target TrustZone are also applicable to this Arm-based GPU TEE prototype. The impact of these side-channel attacks on TrustZone indirectly affects Arm-based GPU TEE (cf. attack surfaces ➀ and ➁ in Figure 7). The potential consequences can be summarized as follows: (i) An attacker could escalate privileges to root to submit numerous malicious tasks (Beniamini, 2016; Shen, 2015). However, establishing covert channels between malicious and victim tasks is unlikely, as GPU tasks are executed sequentially on Arm endpoint GPUs (Deng et al., 2022). (ii) An attacker may monitor cache activity to extract AES keys from TrustZone (Guanciale et al., 2016; Zhang et al., 2016b; Lipp et al., 2016), thus compromising the confidentiality of data transferred between the CPU and GPU.
7.3.2. Microarchitectural Side-Channel and Covert-Channel Attacks on Integrated GPU
As summarized in LABEL:tab:GPU_attacks, several studies have investigated side-channel and covert-channel attacks on integrated GPUs, with some of these potentially applicable to Arm-based GPU TEEs prototype (cf. attack surface ➂ in Figure 7).
(i) Covert channels between CPU and iGPU: Dutta et al. (Dutta et al., 2021) establish bidirectional covert channels between a trojan process on the CPU and a spy process on the Intel integrated GPU (vice versa) via the LLC. Considering this attack in an Arm-based GPU TEE, a possible scenario is that of using a trojan process launched on the normal world of the CPU, communicating the bits to the GPU. However, challenges arise due to undocumented LLC/system-level cache behavior and the asymmetric view of cache utilization between the CPU and GPU in the Arm architecture.
(ii) The existing threat model of Arm-based GPU TEEs may not effectively mitigate attacks exploiting data-dependent DRAM traffic and cache utilization induced by compression to reveal user pixels in the normal world (Wang et al., 2024a). Further testing is needed to assess the success rate of such attacks on Arm-based GPU TEEs.
7.3.3. Physical Attacks on GPU Memory
Most existing Arm-based GPU TEE prototypes primarily focus on delivering hardware-enforced isolation to protect sensitive code and data, rather than employing memory encryption mechanisms like Intel SGX’s MEE. Consequently, direct physical access to GPU memory (again also possibly via cold boot attacks (Halderman et al., [n. d.])) could potentially compromise sensitive data (cf. attack surface ➃ in Figure 7). For instance, in StrongBox (Deng et al., 2022), a region in the unified memory defined as Secure Task RAM serves as a fixed and non-secure memory region reserved for confidential GPU applications. Its main purpose is to dynamically allocate memory and establish GPU page table mappings and thus prevent unauthorized access control to this region.
7.3.4. DVFS-based Attacks
In contrast to the DVFS in NVIDIA GPUs, the voltage and frequency of an Arm GPU (e.g., Mali series) are typically controlled by the Power Management Integrated Circuit (PMIC). This component is responsible for managing power distribution to various system elements, including both CPU and GPU. Thus, as depicted in Figure 7 (cf. attack surface ➄), attackers with privileges can potentially induce hardware faults by manipulating processor voltage and frequency settings through the PMIC. Additionally, assuming an attacker gains physical access to the voltage regulator or clock control module within the System on Chip (SoC), this prompts further investigation into whether hardware-based attacks targeting these components (Chen et al., 2021) are also applicable to Arm GPUs.
8. Conclusion
With the proliferation of heterogeneous systems integrating GPUs, GPU TEEs are emerging as effective solutions for ensuring confidentiality and integrity. In this survey, we first examined existing GPU TEEs prototypes and NVIDIA’s first commercial GPU TEE, providing key insights into their design implications and characteristics. We reviewed recent attacks extended across various types of GPUs, including integrated GPUs, single GPUs, and multiple GPUs, along with the efforts to mitigate these threats. Additionally, we summarized popular attacks against common CPU TEEs such as Intel SGX, Arm TrustZone, and AMD SEV. Based on these insights, we explored potential attacks at both physical and software levels on x86-based and Arm-based GPU TEEs prototypes, as well as the NVIDIA H100 GPU, highlighting the need for careful analysis of the already-complex attack surface. With innovations in GPU architecture and isolation mechanisms in GPU TEEs, this attack surface will further expand, necessitating the evolution of secure designs and defence strategies as these systems are widely deployed across computing domains.
Acknowledgements
This research is partially funded by the Engineering and Physical Sciences Research Council (EPSRC) under grants EP/X03738X/1, EP/V000454/1, and EP/R012598/1. The results feed into DsbDtech.

### v3

@@
 Virtual Graphics Processing Unit
-CC
+$C$
 Confidential Computing
@@
 Trust Domain Extensions
+LLM
+Large Language Model
 Confidential Computing on Heterogeneous CPU-GPU Systems: Survey and Future Directions
-Qifan Wang University of BirminghamBirminghamUK q.wang.3@bham.ac.uk  and  David Oswald University of BirminghamBirminghamUK d.f.oswald@bham.ac.uk
+Qifan Wang Durham UniversityDurhamUnited Kingdom University of BirminghamBirminghamUnited Kingdom qifan.wang2@durham.ac.uk and David Oswald Durham UniversityDurhamUnited Kingdom University of BirminghamBirminghamUnited Kingdom david.f.oswald@durham.ac.uk
 Abstract.
@@
 Confidential computing, Trusted Execution Environment on GPU, hardware security, GPU attacks
-ccs: General and reference Surveys and overviewsccs: Computer systems organization Heterogeneous (hybrid) systemsccs: Security and privacy Hardware attacks and countermeasuresccs: Security and privacy Trusted computing
+journal: CSURjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1journalyear: 2026copyright: acmlicensedccs: General and reference Surveys and overviewsccs: Computer systems organization Heterogeneous (hybrid) systemsccs: Security and privacy Hardware attacks and countermeasuresccs: Security and privacy Trusted computing
 1. Introduction
@@
 To ensure the security and confidentiality of heterogeneous computing systems, researchers have focused on adapting privacy-preserving techniques traditionally used on CPUs to accelerators like GPUs.
-These techniques include cryptographic methods such as Homomorphic Encryption (HE(Mishra et al., 2020; Ng and Chow, 2021; Fan et al., 2023) and Secure Multi-Party Computation (MPC(Mishra et al., 2020; Ng and Chow, 2021; Tan et al., 2021; Watson et al., 2022; Jawalkar et al., 2023), as well as TEEs tailored for accelerators like GPUs (Volos et al., 2018; Jang et al., 2019; Zhu et al., 2020; NVIDIA, 2022a; Jiang et al., 2022; Lee et al., 2022; Yudha et al., 2022; Deng et al., 2022; Ivanov et al., 2023; Mai et al., 2023; Wang et al., 2024b).
+These techniques include cryptographic methods such as Homomorphic Encryption (HE(Mishra et al., 2020; Ng and Chow, 2021; Fan et al., 2023) and Secure Multi-Party Computation (MPC(Mishra et al., 2020; Ng and Chow, 2021; Tan et al., 2021; Watson et al., 2022; Jawalkar et al., 2024), as well as TEEs tailored for accelerators like GPUs (Volos et al., 2018; Jang et al., 2019; Zhu et al., 2020; NVIDIA, 2022a; Jiang et al., 2022; Lee et al., 2022; Yudha et al., 2022; Deng et al., 2022; Ivanov et al., 2023; Mai et al., 2023; Wang et al., 2024a).
 TEEs offer certain advantages over cryptographic methods, particularly in performance and ease of deployment, providing a robust hardware-based security mechanism that is relatively straightforward to implement and manage.
 However, in contrast to cryptographic techniques, the security of TEEs relies on trust in the manufacturer and is difficult to analyse and prove rigorously.
-Recent research has explored extending TEEs to accelerators like GPUs, both for x86 (Volos et al., 2018; Jang et al., 2019; Zhu et al., 2020; NVIDIA, 2022a; Lee et al., 2022; Yudha et al., 2022; Ivanov et al., 2023; Mai et al., 2023) and Arm (Jiang et al., 2022; Deng et al., 2022; Wang et al., 2024b) architectures.
+Recent research has explored extending TEEs to accelerators like GPUs, both for x86 (Volos et al., 2018; Jang et al., 2019; Zhu et al., 2020; NVIDIA, 2022a; Lee et al., 2022; Yudha et al., 2022; Ivanov et al., 2023; Mai et al., 2023) and Arm (Jiang et al., 2022; Deng et al., 2022; Wang et al., 2024a) architectures.
 Most of these approaches involve modifying existing CPU TEEs (especially process-based TEEs like Intel Software Guard Extensions (SGX) and Arm TrustZone (TrustZone)) to secure the GPU software stack, such as GPU drivers, and proposing strategies to ensure the isolation and confidentiality of GPU kernel executions.
@@
 Figure 1.
-Current state of security research on heterogeneous computing systems with GPU.
+Current state of security research on heterogeneous computing systems with GPU (from the past 15 years).
 GPU Crypto refers to papers that protect the privacy of GPU-based applications using cryptographic techniques such as HE and MPC.
-AA, MSCA, MCCA, PSCA, SFIA refer to architectural, microarchitectural side-channel/covert-channel, physical side-channel, and software-based fault injection attacks.
+GPU attacks refer to architectural (AA), microarchitectural side-channel/covert-channel (MSCA/MCCA), physical side-channel (PSCA), and software-based fault injection attacks (SFIA).
 1999200420092014201920242029 First GPU, NVIDIA GeForce256 ARM TrustZone Launched in ARM11 First Virtualized GPU (NVIDIA VGX) First Attack Exploiting GPU Vuln.
-(Lee et al., 2014; Maurice et al., 2014) First MSCA on GPU (Luo et al., 2015) Intel SGX Launched in Skylake First MCCA on GPU (Naghibijouybari et al., 2017) AMD SEV Launched in EPYC First FIA on GPU (Frigo et al., 2018) First x86-based GPU TEE (Volos et al., 2018) First MIG-enabled GPU, NVIDIA A100 Intel TDX Announced AMD SEV-SNP Available in Milan First Arm-based GPU TEE (Deng et al., 2022) First Commercial GPU TEE (NVIDIA, 2022a) First MCCA on MIG-enabled GPU (Zhang et al., 2023) First Arm CCA-based GPU TEE (Wang et al., 2024b)
+(Lee et al., 2014; Maurice et al., 2014) First MSCA on GPU (Luo et al., 2015) Intel SGX Launched in Skylake First MCCA on GPU (Naghibijouybari et al., 2017) AMD SEV Launched in EPYC First FIA on GPU (Frigo et al., 2018) First x86-based GPU TEE (Volos et al., 2018) First MIG-enabled GPU, NVIDIA A100 Intel TDX Announced AMD SEV-SNP Available in Milan First Arm-based GPU TEE (Deng et al., 2022) First Commercial GPU TEE (NVIDIA, 2022a) First MCCA on MIG-enabled GPU (Zhang et al., 2023) First Arm CCA-based GPU TEE (Wang et al., 2024a)
 Figure 2. Timeline of significant events in secure heterogeneous computing systems.
@@
 Table 1. Summary of existing popular GPUs
-Accelerator Year Features Virtualization Memory Sharing Memory Size
-None L1/L2 Cache DRAM L1 L2 DRAM
+Accelerator Year Features Virtualization Shared Memory Hierarchy Memory Size
+None L2/L3 Cache DRAM L1 L2 DRAM
 Intel HD/Iris/UHD/Xe 2010-2020 Integrated graphics GVT $\checkmark$ $\checkmark$ Shared Shared Shared
@@
 Integrated/Discrete GPU.
-Integrated GPUs, such as Intel HD/Iris/UHD/Xe Graphics, Arm, and Qualcomm GPUs listed in Table 1, are integrated on the same die or package as the CPU.
+Integrated GPUs, e.g., Intel HD/Iris/UHD/Xe Graphics, Arm, and Qualcomm GPUs in Table 1, are integrated on the same die or package as the CPU.
 These GPUs share the system’s memory with the CPU.
 Discrete GPUs are standalone and dedicated graphics processing units separate from the CPU.
-They have their dedicated memory, which is distinct from the system memory (e.g., NVIDIA and AMD GPUs listed in Table 1).
+They have their dedicated memory, which is distinct from the system memory (e.g., NVIDIA and AMD GPUs in Table 1).
+NVIDIA GPUs have long supported host memory access via PCIe and Unified Memory, but recent architectures like the H100—especially in Grace Hopper configurations—enable tighter CPU-GPU integration with coherent memory sharing over NVLink-C2C (NVIDIA, 2022b).
 GPU virtualization.
@@
 In this survey, we detail various attack vectors, including architectural/logical attacks, microarchitectural side-channel and covert-channel attacks, physical side-channel attacks, and fault injection.
-2.2.1. Architectural/Logical Attacks
-These attacks encompass a class of attacks that exploit vulnerabilities or weaknesses in the underlying hardware architecture.
-These attacks aim to compromise system by targeting the physical or logical design of key hardware elements, including processing units (e.g., processors, hardware accelerators, and co-processors), the memory subsystem, and communication components responsible for interfacing with other system elements (Ghasempouri et al., 2023).
+2.2.1. Architectural Attacks
+These attacks exploit vulnerabilities in the design and configuration of hardware components.
+This includes misuse of GPU or processor features, privilege escalation, and weak memory isolation.
+Attackers may target processing units (e.g., processors, hardware accelerators, and co-processors), the memory subsystem, and communication components responsible for interfacing with other system elements to bypass protections or leak sensitive data (Ghasempouri et al., 2023).
+Logical attacks, a subset of architectural attacks, exploit flaws in the intended logic or functional behavior of the hardware.
+These may involve undocumented instructions, incorrect permission checks, or logic bugs that allow attackers to alter control flow or data access patterns.
 2.2.2. Microarchitectural Side-Channel and Covert-Channel Attacks
@@
 In this work, we primarily focus on non-invasive fault injection attacks.
+2.2.5. Comparison of Attacks on CPUs and GPUs
+While attacks on CPU and GPU share some fundamental principles—such as exploiting timing variations, memory leakage, or shared resources—their feasibility, targeted attack surfaces, and required attacker capabilities diverge significantly due to underlying architectural and operational differences.
+CPUs generally enforce stricter privilege separation and memory isolation, making attacks like residual memory leakage less effective.
+In contrast, GPUs have historically lacked such safeguards, making them more vulnerable to memory remanence and weak access control.
+For microarchitectural attacks, GPUs introduce new leakage vectors due to their massive parallelism and SIMT execution model, allowing attackers to exploit fine-grained resource contention (e.g.,, warps, shared caches) at scale.
+GPUs also expose performance counters and APIs more openly than CPUs, increasing their attack surface.
+In physical side-channel attacks, GPUs tend to emit more observable signals due to their higher power usage and bandwidth, and their close coupling with CPUs introduces additional leakage through shared interconnects or MMIO.
+For fault injection, while CPUs are often targeted through physical glitching, GPU attacks frequently exploit software-exposed features like DVFS or power-state manipulation to induce faults during kernel execution.
+ML workloads on GPUs are resilient to random faults but remain vulnerable to precise, targeted injections.
+Overall, while CPUs benefit from more mature defense mechanisms, GPU security is still evolving.
+Adapting proven CPU defenses to GPU-specific designs and threat models remains an important direction for future research.
 2.3. TEE
@@
 A lightweight Realm Management Monitor (RMM(Arm, 2023b), operating within the hypervisor layer of the realm world, ensures memory isolation between realms.
-CCA incorporates a root world to accommodate the highest privilege monitor code stored in firmware (Wang et al., 2024b).
+CCA incorporates a root world to accommodate the highest privilege monitor code stored in firmware (Wang et al., 2024a).
 Intel TDX.
@@
 {forest}
-forked edges, for tree= grow=east, reversed=true,anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, rounded corners,align=left, minimum width=2.5em, s sep=3pt, inner xsep=2pt, inner ysep=1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.8em,font=,, where level=2text width=5.6em,font=, where level=3text width=9.0em,font=,[Confidential Computing on Heterogeneous CPU-GPU Systems, ver [GPU TEEs
-(Section 4) [x86-based
-GPU TEEs [Graviton (Volos et al., 2018), HIX (Jang et al., 2019), HETEE (Zhu et al., 2020), NVIDIA H100 (NVIDIA, 2022a), TNPU (Lee et al., 2022), LITE (Yudha et al., 2022), SAGE (Ivanov et al., 2023), Honeycomb (Mai et al., 2023),leaf,text width=38em] ] [Arm-based
-GPU TEEs [StrongBox (Deng et al., 2022), CRONUS (Jiang et al., 2022), CAGE (Wang et al., 2024b),leaf,text width=15em] ] ] [Security of GPUs
-(Section 5) [Architectural/
-Logical Attacks [Lee et al.
-(Lee et al., 2014), Maurice et al.
-(Maurice et al., 2014), Pietro et al.
-(Pietro et al., 2016), Zhou et al.
-(Zhou et al., 2016), HE et al.
-(HE et al., 2020),leaf,text width=30em] ] [Microarchitectural
-Side/Covert-
-Channel Attacks [Attacks on Integrated GPU [Frigo et al.
-(Frigo et al., 2018), HE et al.
-(HE et al., 2020), Dutta et al.
-(Dutta et al., 2021), Taneja et al.
-(Taneja et al., 2023), Wang et al.
-(Wang et al., 2024a),leaf,text width=31em] ] [Attacks on Discrete GPU [Jiang et al.
-(Jiang et al., 2016), Jiang et al.
-(Jiang et al., 2017), Ahn et al.
-(Ahn et al., 2021a), Giner et al.
-(Giner et al., 2024), Luo et al.
-(Luo et al., 2019), Zhu et al.
-(Zhu et al., 2021),
-Hu et al. (Hu et al., 2020), Naghibijouybari et al. (Naghibijouybari et al., 2017), Wei et al. (Wei et al., 2020), Ahn et al. (Ahn et al., 2021b), Nayak et al. (Nayak et al., 2021),
-Cronin et al.
-(Cronin et al., 2021), Yang et al.
-(Yang et al., 2022), Dutta et al.
-(Dutta et al., 2023), Zhang et al.
-(Zhang et al., 2023),leaf,text width=32em] ] [Countermeasures [Wang and Lee (Wang and Lee, 2007), Kong et al.
-(Kong et al., 2009), Liu et al.
-(Liu et al., 2016), Zhang et al.
-(Zhang et al., 2023), Wang et al.
-(Wang et al., 2014),
-Shafiee et al.
-(Shafiee et al., 2015), Di et al.
-(Di et al., 2021), Kadam et al.
-(Kadam et al., 2018), Kadam et al.
-(Kadam et al., 2020), Lin et al.
-(Lin et al., 2019),leaf,text width=30em] ] ] [Physical Side-
-Channel Attacks [Physical Access-based [Luo et al.
-(Luo et al., 2015), Gao et al.
-(Gao et al., 2018), Hu et al.
-(Hu et al., 2020), Chmielewski and Weissbart (Chmielewski and Weissbart, 2021), Maia et al.
-(Maia et al., 2022),
-Horvath et al.
-(Horvath et al., 2023), Taneja et al.
-(Taneja et al., 2023),leaf,text width=32em] ] [Remote Access-based [Zhan et al.
-(Zhan et al., 2022), Liang et al.
-(Liang et al., 2022),leaf,text width=12em] ] [Countermeasures [Mangard et al.
-(Mangard et al., 2008), Zhan et al.
-(Zhan et al., 2022),leaf,text width=13em] ] ] [Fault Injection
-Attacks [Software-based [Frigo et al.
-(Frigo et al., 2018), Sabbagh et al.
-(Sabbagh et al., 2020), Sabbagh et al.
-(Sabbagh et al., 2021), Sun et al.
-(Sun et al., 2023), Santos et al.
-(Santos et al., 2021),leaf,text width=32em] ] ] ] [Security of CPU
-TEEs (Section 6) [Security of Intel
-SGX [Controlled Channel/Address
-Translation-based Attacks [Xu et al.
-(Xu et al., 2015), Shinde et al.
-(Shinde et al., 2016), Van Bulck et al.
-(Van Bulck et al., 2017a), Van Bulck et al.
-(Van Bulck et al., 2017b), Wang et al.
-(Wang et al., 2017),
-Gyselinck et al.
-(Gyselinck et al., 2018), Kim et al.
-(Kim et al., 2019), Fei et al.
-(Fei et al., 2021),leaf,text width=33em] ] [Cache-based Attacks [Götzfried et al.
-(Götzfried et al., 2017), Moghimi et al.
-(Moghimi et al., 2017), Schwarz et al.
-(Schwarz et al., 2017), Brasser et al.
-(Brasser et al., 2017b), Dall et al.
-(Dall et al., 2018),
-Moghimi et al.
-(Moghimi et al., 2019),leaf,text width=32em] ] [Branch Prediction Attacks [Fei et al.
-(Fei et al., 2021), Lee et al.
-(Lee et al., 2017), Chen et al.
-(Chen et al., 2019), Evtyushkin et al.
-(Evtyushkin et al., 2018), Huo et al.
-(Huo et al., 2020), Koruyeh et al.
-(Koruyeh et al., 2018),leaf,text width=34em] ] [Transient Execution Attacks [Chen et al.
-(Chen et al., 2019), Van Bulck et al.
-(Van Bulck et al., 2018), Van Schaik et al.
-(Van Schaik et al., 2020), Ragab et al.
-(Ragab et al., 2021),leaf,text width=27em] ] [Software-based Fault
-Injection Attacks [Murdock et al.
-(Murdock et al., 2020), Kenjar et al.
-(Kenjar et al., 2020), Qiu et al.
-(Qiu et al., 2020), Jang et al.
-(Jang et al., 2017),leaf,text width=24em] ] [Hardware-based Attacks [Chen et al.
-(Chen et al., 2021), Chen and Oswald (Chen and Oswald, 2023), Lee et al.
-(Lee et al., 2020),leaf,text width=18em] ] [Interface-based Attacks [Van Bulck et al.
-(Van Bulck et al., 2019), Alder et al.
-(Alder et al., 2020), Cloosters et al.
-(Cloosters et al., 2020), Khandaker et al.
-(Khandaker et al., 2020), Alder et al.
-(Alder et al., 2024),leaf,text width=32em] ] ] [Security of
-Arm TrustZone [Implementation Bugs [Beniamini (Beniamini, 2016), Shen (Shen, 2015), Komaromy (Komaromy, 2018),leaf,text width=15em] ] [Cache-based Attacks [Zhang et al.
-(Zhang et al., 2016c), Guanciale et al.
-(Guanciale et al., 2016), Lipp et al.
-(Lipp et al., 2016), Zhang et al.
-(Zhang et al., 2016b), Cho et al.
-(Cho et al., 2018),leaf,text width=30em] ] [Branch Prediction Attacks [Ryan (Ryan, 2019),leaf,text width=4em] ] [Software-based Fault
-Injection Attacks [Tang et al. (Tang et al., 2017), Qiu et al. (Qiu et al., 2020), Carru (Carru, 2017),leaf,text width=16em] ] ] [Security of
-AMD SEV [Memory Encryption Issues [Hetzelt and Buhren (Hetzelt and Buhren, 2017), Buhren et al.
-(Buhren et al., 2017), Du et al.
-(Du et al., 2017), Morbitzer et al.
-(Morbitzer et al., 2018), Li et al.
-(Li et al., 2021b),
-Li (Li, 2022), Li et al. (Li et al., 2021a), Li et al. (Li et al., 2021c),leaf,text width=32em] ] [Software-based Fault
-Injection Attacks [Morbitzer et al.
-(Morbitzer et al., 2021), Buhren et al.
-(Buhren et al., 2021), Zhang et al.
-(Zhang et al., 2024), Schlüter et al.
-(Schlüter et al., 2024a), Schlüter et al.
-(Schlüter et al., 2024b),leaf,text width=33em] ] [Interrupt-driven Attacks [Wilke et al.
-(Wilke et al., 2024),leaf,text width=6em] ] ] ] [Potential Attacks
-against GPU TEEs
-(Section 7) [x86-based
-GPU TEE [Existing Attacks against Intel SGX,leaf,text width=12em] [ Dynamic Voltage and Frequency Scaling (DVFS)-based Attacks,leaf,text width=21em] [Physical Attacks on PCIe bus and GPU Memory,leaf,text width=17em] [Architectural/Logical attacks on GPU,leaf,text width=13em] ] [NVIDIA Hopper
-H100-based
-GPU TEE [Existing attacks against VM-based TEEs,leaf,text width=14em] [DVFS-based Attacks,leaf,text width=8em] [Microarchitectural Side/Covert-Channel Attacks on GPU,leaf,text width=19em] ] [Arm-based
-GPU TEE [Existing attacks against Arm TrustZone,leaf,text width=13em] [Microarchitectural Side/Covert-Channel Attacks on Integrated GPU,leaf,text width=22em] [Physical Attacks on GPU Memory,leaf,text width=12em] [DVFS-based Attacks,leaf,text width=8em] ] ] ]
-Figure 3. Taxonomy of confidential computing on heterogeneous CPU-GPU systems
+Figure 3. Overview of confidential computing on heterogeneous CPU-GPU systems
 Confidential computing on GPUs is an emerging field that ensures sensitive data processed on GPUs remains protected from unauthorized access and tampering even for attackers with privileged access.
@@
 In Section 5, we give an overview of existing attacks targeting GPUs based on the attack types listed in Section 2.
-We also review state-of-the-art attacks against existing CPU TEEs, such as Intel SGX, Arm TrustZone, and AMD SEV in Section 6.
+We also review state-of-the-art attacks against existing CPU TEEs, such as Intel SGX, Arm TrustZone, and AMD SEV in the supplemental material due to the limited space of this article.
 Finally, based on these GPU TEE designs and the summarized attacks, we discuss potential attacks against existing (and future) GPU TEEs in Section 6.
@@
 NVIDIA H100 (NVIDIA, 2022a) Intel x86 NVIDIA GPU Dedicated Inside confidential VM Hardware-level protection $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-[…]
+TNPU (Lee et al., 2022) Intel x86 NPU Unified Inside CPU enclave SGX Enclave $\checkmark$} $\checkmark$} $\checkmark$ $\checkmark$ $\checkmark$
 LITE (Yudha et al., 2022) Intel/AMD x86 NVIDIA GPU Dedicated N/A Software-based unified encryption domain $\checkmark$ $\checkmark$ $\checkmark$
-[…]
-[…]
+StrongBox (Deng et al., 2022) Arm Arm Mali GPU Unified Inside normal world TrustZone $\checkmark$} $\checkmark$} $\checkmark$ $\checkmark$
+CRONUS (Jiang et al., 2022) Arm General accelerators Dedicated Inside secure world TrustZone virtualization $\checkmark$} $\checkmark$} $\checkmark$ $\checkmark$
 SAGE (Ivanov et al., 2023) Intel x86 NVIDIA GPU Dedicated N/A Software-based verification function $\checkmark$ $\checkmark$
-[…]
-[…]
-[…]
+Honeycomb (Mai et al., 2023) AMD x86 AMD GPU Dedicated Inside confidential VM AMD SEV-SNP $\checkmark$} $\checkmark$} $\checkmark$ $\checkmark$
+CAGE (Wang et al., 2024a) Arm Arm Mali GPU Unified Inside normal world RME, GPC $\checkmark$} $\checkmark$ $\checkmark$
+Memory indicates whether the accelerators share the memory with the CPU or not.
+HW Modification explains if need modification on hardware (● = yes, ○ = no).
+Secure RM and GPU ME are Secure Resource Management and GPU Memory Encryption.
+$\dagger$}} indicates that scheme needs additional support (e.g., hardware security features provided in NVIDIA H100 GPU or next-generation Arm devices), as they are not currently considering them.
 x86-based GPU TEEs.
@@
 Each enclave encapsulates a specific type of computation like GPU computation and allows multiple enclaves to spatially share a single accelerator.
-To support confidential GPU computing on next-generation Arm devices, CAGE (Wang et al., 2024b) presents a new scheme for Arm CCA’s realm-style architecture.
+To support confidential GPU computing on next-generation Arm devices, CAGE (Wang et al., 2024a) presents a new scheme for Arm CCA’s realm-style architecture.
 It offloads resource-intensive but data-independent tasks to NW, while protecting sensitive data within realms through a shadow task mechanism.
@@
 In this section, we discuss several key metrics in designing an effective and secure GPU TEEs.
+Comparison to CPU TEEs.
+While both CPU and GPU TEEs aim to provide secure and isolated execution environments, their designs differ significantly due to fundamental differences in microarchitecture and usage patterns.
+CPU TEEs, such as Intel SGX and AMD SEV-SNP, operate in control-centric environments with centralized execution and predictable memory behavior, making them well-suited for fine-grained access control, secure paging, and privilege enforcement.
+In contrast, GPUs are optimized for massive data-parallel workloads, with thousands of lightweight threads, wide memory bandwidth, and deep memory hierarchies.
+These characteristics make it difficult to directly apply traditional TEE techniques like enclave management or synchronized memory encryption.
+Moreover, GPUs often lack built-in privilege separation and rely on the CPU for task scheduling and memory control, expanding the attack surface.
+As a result, GPU TEE designs typically require either hardware modifications to decouple from the host (e.g., Graviton, HIX) or leverage CPU-side TEEs for protection and communication (e.g., Honeycomb, SAGE).
+This architectural gap explains the greater diversity and complexity in GPU TEE designs compared to their CPU counterparts.
 Threat models.
@@
 To fill this gap, LITE (Yudha et al., 2022) designs a software-based unified encryption domain to protect sensitive data stored in GPU memory, including L1/L2 cache and device memory.
-Additionally, NVIDIA H100 (NVIDIA, 2022a) and TNPU (Lee et al., 2022) support hardware-based memory encryption, while CAGE (Wang et al., 2024b) requires hardware-assisted memory encryption support built into future CCA devices.
+Additionally, NVIDIA H100 (NVIDIA, 2022a) and TNPU (Lee et al., 2022) support hardware-based memory encryption, while CAGE (Wang et al., 2024a) requires hardware-assisted memory encryption support built into future CCA devices.
 Hardware changes.
@@
 Consequently, works like LITE (Yudha et al., 2022), SAGE (Ivanov et al., 2023), and Honeycomb (Mai et al., 2023), seek purely software-based approaches for trusted GPU execution.
-CAGE (Wang et al., 2024b) primarily leverages generic hardware features in Arm CCA and GPUs without requiring additional hardware changes or customization of existing CPUs and GPUs.
+CAGE (Wang et al., 2024a) primarily leverages generic hardware features in Arm CCA and GPUs without requiring additional hardware changes or customization of existing CPUs and GPUs.
 TCB size.
@@
 An effective and efficient way to reduce the TCB is to place most of the GPU software stack in the untrusted OS, while keeping only the key functions for controlling the GPU in a trusted area of the CPU, such as an SGX enclave.
-For example, CAGE (Wang et al., 2024b) achieves a ‘thin’ TCB for realms by using its shadow task mechanism instead of loading the heavyweight GPU software into the TEE.
+For example, CAGE (Wang et al., 2024a) achieves a ‘thin’ TCB for realms by using its shadow task mechanism instead of loading the heavyweight GPU software into the TEE.
 It introduces 2–-26K lines of code for a CPU-side isolation and 1,301 lines of code for monitor, resulting in the lowest TCB size among current Arm-based GPU TEEs.
@@
 Several factors, such as the number of cryptographic operations, the frequency of switching between secure and normal worlds, and the integrity checks, play significant roles in optimizing performance compared to insecure baselines on GPUs.
-For example, CAGE (Wang et al., 2024b) outperforms StrongBox (Deng et al., 2022) by 1.96%–13.16% on six Rodinia benchmarks.
+For example, CAGE (Wang et al., 2024a) outperforms StrongBox (Deng et al., 2022) by 1.96%–13.16% on six Rodinia benchmarks.
 This performance gain is attributed to CAGE’s ability to securely transfer plaintext data to buffers and store the results without introducing additional cryptographic operations.
@@
 It would be insightful to conduct a comprehensive performance comparison across different platforms once the entire workflow for confidential computing on the NVIDIA H100 GPU is fully available in the future.
+Trade-off between cryptographic methods and GPU TEEs.
+In general, while cryptographic approaches like HE and MPC offer strong confidentiality guarantees, they remain impractical for many GPU workloads due to significant computation/communication overhead, memory usage, and developer complexity.
+Recent GPU-based MPC systems such as CryptGPU (Tan et al., 2021), Piranha (Watson et al., 2022), and Orca (Jawalkar et al., 2024) show that performance is highly sensitive to the number of computing participants and choice of operators, making them difficult to develop.
+Piranha further highlights that communication overhead becomes a major bottleneck in wide-area network settings—an important consideration since real-world deployments often involve geographically distributed GPU providers.
+For HE, even with dedicated HE accelerator support, as in TensorFHE (Fan et al., 2023), significant modifications are needed to underlying schemes (e.g., CKKS) and operations (e.g., NTT).
+Applying HE to ML workloads also requires careful tuning of HE parameters and may necessitate changes to model architecture, complicating deployment in applications.
+Therefore, GPU TEEs offer a more practical path forward, enabling protection of existing workloads with less developer effort and minimal modification to ML models.
+While GPU TEEs are still in their early stages and only applicable on certain machines, we believe they hold strong potential for future GPU-based applications, especially in ML.
 5. Security of GPUs
@@
 LABEL:tab:GPU_attacks provides a summary of attacks, target GPUs, threat model, attack goals and methods.
-{ThreePartTable}{TableNotes}
+Based on LABEL:tab:GPU_attacks, we highlight several interesting take-away points: ❶ Most attacks target single discrete GPU systems, with memory and timing behavior serving as the primary leakage vectors.
+However, recent research has begun to explore attacks on multi-GPU and multi-instance GPU environments, revealing new vulnerabilities in GPU-to-GPU communication and resource partitioning.
+❷ A majority of attacks can be carried out with only user-level permissions, indicating that isolation mechanisms in modern GPU systems remain inadequate, particularly in multi-tenant or cloud-based deployments.
+❸ Attacks span across multiple GPU vendors (e.g., NVIDIA, AMD, Qualcomm) and consistently aim to recover sensitive information, such as AES keys, Neural Network (NN) model parameters, and user input keystrokes.
+Notably, attacks targeting more complex GPU workloads, such as Large Language Models, remain largely unexplored.
+❹ While impactful, fault injection and covert channels appear less frequently in the literature.
+These often require stronger attacker capabilities or intricate coordination between software and hardware layers, which may explain their limited exploration.
+❺ Emerging attacks have begun to explore cross-GPU timing channels and remote leakage via browsers, suggesting that cloud-hosted GPU services—especially those accessible by multiple users—may become an increasingly important and vulnerable attack surface.
+Table 3. Comparison of different attacks on heterogeneous systems (oldest results first)
+Work&Year Accelerator Threat Model Phy- sical Attack Scenarios & Goals Method Main results
+Model Permissions
+Architectural Attacks
+Lee et al.’14 (Lee et al., 2014) NVIDIA/AMD GPU sGPU User-level permissions Infer which web pages a victim user has visited Exploit that GPUs do not initialize newly allocated memory pages $95.4\%$ accuracy
+Maurice et al.’14 (Maurice et al., 2014) NVIDIA GPU sGPU Require root privileges based on accessing GPU memory via GPU runtime or PCI configuration Recover data of victim’s previously execution on GPU in virtualized environment Exploit the vulnerability of memory clearing on GPU Access data stored in GPU global memory after soft reboot
+Pietro et al.’16 (Pietro et al., 2016) NVIDIA GPU sGPU User-level permissions Read information stored in memory and infer the Advanced Encryption Standard (AES) keys Leverage the lack of adequate data allocation and data cleaning $70\%$, $28\%$ leakages
+Zhou et al.’16 (Zhou et al., 2016) NVIDIA/AMD GPU sGPU User-level permissions Recover users’ profile images and email contents Reconstruct images from identified image-like tiles in GPU memory $79.3\%$ successful recovery
+Microarchitectural Side-Channel/Covert-Channel Attacks
+Luo et al.’15 (Luo et al., 2015) NVIDIA GPU sGPU Unprivileged attacker has physical access to the victim GPU $\checkmark$ Recover all of the last round AES key byte Apply correlation power analysis to analyze the power consumption of the GPU Recover the key with 160,000 power traces
+Jiang et al.’16 (Jiang et al., 2016) NVIDIA GPU sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Leverage the relation between encryption kernel’s runtime and the key 30 min with 1M timing samples
+Jiang et al.’17 (Jiang et al., 2017) NVIDIA GPU sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Exploit timing variability due to shared memory bank conflicts.
+Use 10M timing samples
+Naghibijouybari et al.’17 (Naghibijouybari et al., 2017) NVIDIA GPU sGPU Trojan and spy co-exist on a GPU, both with user-level permission Infer transmitted data Reverse engineer warp schedulers; learn contention on caches and memory 44 MB/s (error-free)
+Gao et al.’18 (Gao et al., 2018) NVIDIA GPU sGPU Unprivileged attacker has physical access to the victim GPU $\checkmark$ Recover AES private key Sample MMIO traces via trigger; exploit parallel MMIO leakages Recover the key with 11k MMIO traces
+Naghibijouybari et al.’18 (Naghibijouybari et al., 2018) NVIDIA GPU sGPU Spy co-locates with the victim program with user-level permission Infer victim’s accessed website data and the number of neurons Learn contention via memory allocation, performance counter, and timing $90\%$, $94.8\%$, $88.6\%$ accuracy
+Luo et al.’19 (Luo et al., 2019) NVIDIA GPU sGPU User-level permissions; available to timing information Recover RSA private key Capture the parallel characteristics of an RSA on GPU via timing model Recover 512/2048-bit key in 15 min/8 h
+Wei et al.’20 (Wei et al., 2020) NVIDIA GPU sGPU Spy co-locates with the victim program with user-level permission Recover the layer composition and hyper-parameters of DNN Leverage context-switching penalties when switching different kernels $76.9\%$-$100\%$ accuracy
+Hu et al.’20 (Hu et al., 2020) NVIDIA GPU sGPU Physically access the victim GPU with user-level permissions $\checkmark$ Obtain the complete model architecture information Learn the inter-layer temporal association information from kernel execution $75.9\%$ accuracy
+HE et al.’20 (HE et al., 2020) Intel HD 530 iGPU Unprivileged adversary and the victim co-locate on the same GPU, but are isolated (1) Recover AES key; (2) monitor the browser activity; (3) conduct a covert channel to exchange data Exploit context management with uncleared registers and shared memory during context switches Recover AES-128/192 in 0.15 s/2 min; $90.7\%$ accuracy; 1.3–8 Gbps
+Ahn et al.’21 (Ahn et al., 2021b) NVIDIA GPU sGPU Trojan and spy co-exist on a GPU, both with user-level permissions Infer transmitted data Exploit interconnect channel contention $24$ MB/s
+Cronin et al.’21 (Cronin et al., 2021) Apple GPU sGPU Exploit JavaScript website and trick user into installing a malicious app Track the user’s web activity Exploit the shared cache between the CPU-cores and peripherals $90\%$, $95\%$ accuracy
+Zhu et al.’21 (Zhu et al., 2021) NVIDIA GPU sGPU Install a PCIe bus snooping device to monitor and log the PCIe traffic Recover the DNN model (semantically identical architecture) Reverse engineer GPU command headers; collect PCIe packets via snooping Similar accuracy as original ones
+Ahn et al.’21 (Ahn et al., 2021a) NVIDIA GPU sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Recover early/later key bytes via timing correlation and cache-collision Use 100k timing samples
+Dutta et al.’21 (Dutta et al., 2021) Intel HD Graphics iGPU Spy and trojan run on CPU and GPU with user-level permissions Infer transmitted data and track user’s cache activity CovA1: prime+probe attack on LLC; CovA2: observe delays on the ring bus CovA1: 120 KB/s, CovA2: 400 KB/s
+Nayak et al.’21 (Nayak et al., 2021) NVIDIA GPU sGPU User-level permissions Infer the accessed data from Virginian database Leverage the shared last-level TLB (reverse engineering) and prime+probe 81 KB/s (MPS-enabled)
+Yang et al.’22 (Yang et al., 2022) Qualcomm Adreno sGPU Install and launch an eavesdropping application on the victim device Eavesdrop the user’s input credentials (login usernames and passwords) through onscreen keyboard Read raw values from the GPU device file and identify GPU events from key presses via ML models Correctly infer more than $80\%$ of user’s credential inputs
+Zhan et al.’22 (Zhan et al., 2022) NVIDIA/AMD GPU sGPU Unprivileged attacker eavesdrops on a victim $\checkmark$ Website fingerprinting and keystroke timing inference Exploit EM caused by Dynamic Voltage and Frequency Scaling (DVFS) $95.6\%$ accuracy and $90\%$ keystrokes
+Liang et al.’22 (Liang et al., 2022) NVIDIA GPU sGPU Attacker and victim GPUs may be separated by meters $\checkmark$ Infer the number of layers/kernels, sizes of layers and strides Exploit far-field EM signals emanated from a GPU Recover VGG’s and ResNet’s parameters
+Maia et al.’22 (Maia et al., 2022) NVIDIA GPU sGPU Unprivileged attacker with physical proximity $\checkmark$ Infer NN architectures and layer parameters Exploit the magnetic flux emanating from a GPU’s power cable Up to $96.8\%$ accuracy
+Horvath et al.’23 (Horvath et al., 2023) NVIDIA GPU sGPU User-level permission with physical access and know the CNN architecture $\checkmark$ Recover the weights and biases of CNN Correlation electromagnetic analysis N/A
+Dutta et al.’23 (Dutta et al., 2023) NVIDIA GPU mGPU Spy and victim connect via NVLink with user-level permissions CovA: infer message; SCA: infer the number of neurons of NN Mount Prime+Probe attack on the L2 cache of victim GPU CovA: 4 MB/s; SCA: $99.91\%$ accuracy
+Zhang et al.’23 (Zhang et al., 2023) NVIDIA GPU mGPU Unprivileged spy and trojan on an MIG-enabled GPU Infer transmitted data and classify the running ML frameworks Mount Prime+Probe attack on the shared L3-TLB $31.47$ KB/s
+Taneja et al.’23 (Taneja et al., 2023) Various GPUs iGPU, sGPU Read frequency, power, and temperature from internal sensors (may require privileged access) $\checkmark$ (1) Observe CPU/GPU behaviour; (2) recover images and user’s browsing history; (3) fingerprint websites (1) Use data from internal sensors; (2) observe throttling via timing; (3) leverage frequency and power bursts on iGPUs Pixel stealing with accuracy of 60%–90%
+Wang et al.’24 (Wang et al., 2024b) Various GPUs iGPU, sGPU Observe coarse-grained pattern redundancy using timer Perform cross-origin pixel stealing attacks in the browser Infer data-dependent DRAM traffic and cache usage via compression channel Pixel stealing with accuracy 68.6%–99.6%
+Giner et al.’24 (Giner et al., 2024) NVIDIA/AMD GPU sGPU SCA: Victim passively visits attacker’s webpage; CovA: Spy accesses the website visited by victim SCA: Detect typed characters and recover the victim’s AES key; CovA: Infer transmitted data Exploit cache contention on the L2 cache via Prime+Probe attack, parallelize eviction set construction SCA: $98\%$ accuracy, recover key in 6  min; CovA: $10.9$ KB/s
+Fault Injection Attacks
+Frigo et al.’18 (Frigo et al., 2018) Intel HD, Qualcomm Adreno iGPU Access iGPU through a malicious application or when victim visits malicious JavaScript website Escape the Firefox sandbox on Android platforms and break ASLR Detect contiguous memory allocation via SCA and then trigger Rowhammer bit flips Hijack the browser and break ASLR in 116 s and 27 s
+Sabbagh et al.’20 (Sabbagh et al., 2020) AMD GPU sGPU DVFS interfaces access Recover AES key Exploit DVFS to introduce random faults during kernel execution Recover key in 15 min
+dos Santos et al.’21 (dos Santos et al., 2021) NVIDIA GPU sGPU Inject faults into pipeline registers, ALU, and memory Improve the probability of incorrect output by injecting faults Combine register transfer level FIA with software FIA Underestimate the error rate by up to $48\%$
 CNN/DNN denotes Convolutional and Deep Neural Network.
 In Attacks, SCA, CovA, and FIA denote Side-Channel Analysis, Covert-Channel, and Fault Injection Attacks, respectively.
-AA/LA represents an attack that exploits an architectural or logical vulnerability on GPU.
+AA represents an attack that exploits an architectural or logical vulnerability on GPU.
 Physical denotes whether this attack leverages physical properties (e.g., power consumption, EM, acoustics, etc.).
@@
 In Main results, we report the central outcomes, e.g., bandwidth and classification accuracy for CovA and SCA, respectively.
-Table 3. Comparison of different attacks on heterogeneous systems (oldest results first)
-Work&Year Accelerator Attack Threat Model Phy- sical Attack Scenarios & Goals Method Main results
-Model Permissions
-Lee et al.’14 (Lee et al., 2014) NVIDIA/AMD GPU AA/LA sGPU User-level permissions Infer which web pages a victim user has visited Exploit that GPUs do not initialize newly allocated memory pages $95.4\%$ accuracy
-Maurice et al.’14 (Maurice et al., 2014) NVIDIA GPU AA/LA sGPU Require root privileges based on accessing GPU memory via GPU runtime or PCI configuration Recover data of victim’s previously execution on GPU in virtualized environment Exploit the vulnerability of memory clearing on GPU Access data stored in GPU global memory after soft reboot
-Luo et al.’15 (Luo et al., 2015) NVIDIA GPU SCA sGPU Unprivileged attacker has physical access to the victim GPU $\checkmark$ Recover all of the last round Advanced Encryption Standard (AES) key byte Apply correlation power analysis to analyze the power consumption of the GPU Recover the key with 160,000 power traces
-Pietro et al.’16 (Pietro et al., 2016) NVIDIA GPU AA/LA sGPU User-level permissions Read information stored in memory and infer the AES keys Leverage the lack of adequate data allocation and data cleaning $70\%$, $28\%$ leakages
-Zhou et al.’16 (Zhou et al., 2016) NVIDIA/AMD GPU AA/LA sGPU User-level permissions Recover users’ profile images and email contents Reconstruct images from identified image-like tiles in GPU memory $79.3\%$ successful recovery
-Jiang et al.’16 (Jiang et al., 2016) NVIDIA GPU SCA sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Leverage the relation between encryption kernel’s runtime and the key 30 min with 1M timing samples
-Jiang et al.’17 (Jiang et al., 2017) NVIDIA GPU SCA sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Exploit timing variability due to shared memory bank conflicts.
-Use 10M timing samples
-Naghibijouybari et al.’17 (Naghibijouybari et al., 2017) NVIDIA GPU CovA sGPU Trojan and spy co-exist on a GPU, both with user-level permission Infer transmitted data Reverse engineer warp schedulers; learn contention on caches and memory $4$ MB/s (error-free)
-Frigo et al.’18 (Frigo et al., 2018) Intel HD, Qualcomm Adreno SCA, FIA iGPU Access iGPU through a malicious application or when victim visits malicious JavaScript website Escape the Firefox sandbox on Android platforms and break ASLR Detect contiguous memory allocation via SCA and then trigger Rowhammer bit flips Hijack the browser and break ASLR in 116 s and 27 s
-Gao et al.’18 (Gao et al., 2018) NVIDIA GPU SCA sGPU Unprivileged attacker has physical access to the victim GPU $\checkmark$ Recover AES private key Sample MMIO traces via trigger; exploit parallel MMIO leakages Recover the key with 11k MMIO traces
-Naghibijouybari et al.’18 (Naghibijouybari et al., 2018) NVIDIA GPU SCA sGPU Spy co-locates with the victim program with user-level permission Infer victim’s accessed website data and the number of neurons Learn contention via memory allocation, performance counter, and timing $90\%$, $94.8\%$, $88.6\%$ accuracy
-Luo et al.’19 (Luo et al., 2019) NVIDIA GPU SCA sGPU User-level permissions; available to timing information Recover RSA private key Capture the parallel characteristics of an RSA on GPU via timing model Recover 512/2048-bit key in 15 min/8 h
-Sabbagh et al.’20 (Sabbagh et al., 2020) AMD GPU FIA sGPU DVFS interfaces access Recover AES key Exploit DVFS to introduce random faults during kernel execution Recover key in 15 min
-Wei et al.’20 (Wei et al., 2020) NVIDIA GPU SCA sGPU Spy co-locates with the victim program with user-level permission Recover the layer composition and hyper-parameters of DNN Leverage context-switching penalties when switching different kernels $76.9\%$-$100\%$ accuracy
-Hu et al.’20 (Hu et al., 2020) NVIDIA GPU SCA sGPU Physically access the victim GPU with user-level permissions $\checkmark$ Obtain the complete model architecture information Learn the inter-layer temporal association information from kernel execution $75.9\%$ accuracy
-HE et al.’20 (HE et al., 2020) Intel HD 530 AA/LA, CovA iGPU Unprivileged adversary and the victim co-locate on the same GPU, but are isolated (1) Recover AES key; (2) monitor the browser activity; (3) conduct a covert channel to exchange data Exploit context management with uncleared registers and shared memory during context switches Recover AES-128/192 in 0.15 s/2 min; $90.7\%$ accuracy; 1.3–8 Gbps
-Santos et al.’21 (Santos et al., 2021) NVIDIA GPU FIA sGPU Inject faults into pipeline registers, ALU, and memory Improve the probability of incorrect output by injecting faults Combine register transfer level FIA with software FIA Underestimate the error rate by up to $48\%$
-Ahn et al.’21 (Ahn et al., 2021b) NVIDIA GPU CovA sGPU Trojan and spy co-exist on a GPU, both with user-level permissions Infer transmitted data Exploit interconnect channel contention $24$ MB/s
-Cronin et al.’21 (Cronin et al., 2021) Apple GPU SCA sGPU Exploit JavaScript website and trick user into installing a malicious app Track the user’s web activity Exploit the shared cache between the CPU-cores and peripherals $90\%$, $95\%$ accuracy
-Zhu et al.’21 (Zhu et al., 2021) NVIDIA GPU Snooping sGPU Install a PCIe bus snooping device to monitor and log the PCIe traffic Recover the DNN model (semantically identical architecture) Reverse engineer GPU command headers; collect PCIe packets via snooping Similar accuracy as original ones
-Ahn et al.’21 (Ahn et al., 2021a) NVIDIA GPU SCA sGPU Send plaintext to victim GPU for encryption with user-level permission Recover the 16-byte last-round key of the AES Recover early/later key bytes via timing correlation and cache-collision Use 100k timing samples
-Dutta et al.’21 (Dutta et al., 2021) Intel HD Graphics CovA iGPU Spy and trojan run on CPU and GPU with user-level permissions Infer transmitted data and track user’s cache activity CovA1: prime+probe attack on LLC; CovA2: observe delays on the ring bus CovA1: 120 KB/s, CovA2: 400 KB/s
-Nayak et al.’21 (Nayak et al., 2021) NVIDIA GPU CovA sGPU User-level permissions Infer the accessed data from Virginian database Leverage the shared last-level TLB (reverse engineering) and prime+probe 81 KB/s (MPS-enabled)
-Yang et al.’22 (Yang et al., 2022) Qualcomm Adreno SCA sGPU Install and launch an eavesdropping application on the victim device Eavesdrop the user’s input credentials (login usernames and passwords) through onscreen keyboard Read raw values from the GPU device file and identify GPU events from key presses via ML models Correctly infer more than $80\%$ of user’s credential inputs
-Zhan et al.’22 (Zhan et al., 2022) NVIDIA/AMD GPU SCA sGPU Unprivileged attacker eavesdrops on a victim $\checkmark$ Website fingerprinting and keystroke timing inference Exploit EM caused by DVFS $95.6\%$ accuracy and $90\%$ keystrokes
-Liang et al.’22 (Liang et al., 2022) NVIDIA GPU SCA sGPU Attacker and victim GPUs may be separated by meters $\checkmark$ Infer the number of layers/kernels, sizes of layers and strides Exploit far-field EM signals emanated from a GPU Recover VGG’s and ResNet’s parameters
-Maia et al.’22 (Maia et al., 2022) NVIDIA GPU SCA sGPU Unprivileged attacker with physical proximity $\checkmark$ Infer neural network architectures and layer parameters Exploit the magnetic flux emanating from a GPU’s power cable Up to $96.8\%$ accuracy
-Horvath et al.’23 (Horvath et al., 2023) NVIDIA GPU SCA sGPU User-level permission with physical access and know the CNN architecture $\checkmark$ Recover the weights and biases of CNN Correlation electromagnetic analysis N/A
-Dutta et al.’23 (Dutta et al., 2023) NVIDIA GPU SCA, CovA mGPU Spy and victim connect via NVLink with user-level permissions CovA: infer message; SCA: infer the number of neurons of NN Mount Prime+Probe attack on the L2 cache of victim GPU CovA: 4 MB/s; SCA: $99.91\%$ accuracy
-Zhang et al.’23 (Zhang et al., 2023) NVIDIA GPU CovA mGPU Unprivileged spy and trojan on an MIG-enabled GPU Infer transmitted data and classify the running ML frameworks Mount Prime+Probe attack on the shared L3-TLB $31.47$ KB/s
-Taneja et al.’23 (Taneja et al., 2023) Various GPUs SCA iGPU, sGPU Read frequency, power, and temperature from internal sensors (may require privileged access) $\checkmark$ (1) Observe CPU/GPU behaviour; (2) recover images and user’s browsing history; (3) fingerprint websites (1) Use data from internal sensors; (2) observe throttling via timing; (3) leverage frequency and power bursts on iGPUs Pixel stealing with accuracy of 60%–90%
-Wang et al.’24 (Wang et al., 2024a) Various GPUs SCA iGPU, sGPU Observe coarse-grained pattern redundancy using timer Perform cross-origin pixel stealing attacks in the browser Infer data-dependent DRAM traffic and cache usage via compression channel Pixel stealing with accuracy 68.6%–99.6%
-Giner et al.’24 (Giner et al., 2024) NVIDIA/AMD GPU SCA, CovA sGPU SCA: Victim passively visits attacker’s webpage; CovA: Spy accesses the website visited by victim SCA: Detect typed characters and recover the victim’s AES key; CovA: Infer transmitted data Exploit cache contention on the L2 cache via Prime+Probe attack, parallelize eviction set construction SCA: $98\%$ accuracy, recover key in 6  min; CovA: $10.9$ KB/s
-[…]
-5.1. Architectural/Logical Attacks
+5.1. Architectural Attacks
 There has been an increase in research examining attacks that target GPUs or their software stack on the CPU.
@@
 This class of vulnerability can usually be fixed by clearing memory (i.e., memory zeroing) before reallocating it to another application.
+Takeaway points.
+Architectural attacks often exploit leftover data in GPU memory due to missing or incomplete memory clearing.
+These attacks require only user-level access and affect both integrated and discrete GPUs, showing that such flaws are common across vendors.
+While many of these vulnerabilities have been patched in recent years, they highlight past weaknesses in GPU memory isolation.
 5.2. Microarchitectural Side-Channel/Covert-Channel Attacks
@@
 Wang et al.
-(Wang et al., 2024a) develop cross-origin pixel stealing attacks in web browsers by exploiting the iGPU compression channel to induce data-dependent DRAM traffic and cache utilization.
+(Wang et al., 2024b) develop cross-origin pixel stealing attacks in web browsers by exploiting the iGPU compression channel to induce data-dependent DRAM traffic and cache utilization.
 Given that the LLC is shared between the CPU and GPU, they demonstrate how to use LLC walk time to infer the GPU-induced CPU cache state changes.
@@
 To access GPU performance counters on Android, they read the raw values of GPU PCs from the GPU device file (e.g., Qualcomm Adreno GPU drivers’ open-source header file, msm_kgsl.h).
+Takeaway points.
+Integrated GPUs introduce new security challenges due to their tight coupling with CPUs and shared microarchitectural resources, such as LLC and interconnects.
+This shared environment enables cross-component attacks, exploiting shared caches, compression channels, and GPU performance counters.
+These attacks are particularly concerning in web and mobile environments, where integrated GPUs are widely deployed and users have minimal control over low-level hardware access.
 5.2.2. Attacks on Discrete GPUs.
@@
 Figure 4. Attacker models for inferring NN models on a single GPU.
-Inferring sensitive information from GPU-based Neural Network (NN): Other attacks exploit side-channel leakage to infer information from ML models like NNs.
+Inferring sensitive information from GPU-based NN: Other attacks exploit side-channel leakage to infer information from ML models like NNs.
 Zhu et al.
@@
 In other words, these factors make the attack more difficult to some extent, but do not fundamentally prevent them.
-Covert-channel attacks where both spy and victim concurrently run on a single GPU: The key point of a successful covert-channel attack is identifying the shared resources utilized by both the trojan and spy, often requiring extensive reverse engineering of GPU architectures.
+Figure 5. Key steps of covert-channel attacks on a single GPU.
+Covert-channel attacks where both spy and victim concurrently run on a single GPU: As summarized in Figure 5, the key point of a successful covert-channel attack is identifying the shared resources utilized by both the trojan and spy, often requiring extensive reverse engineering of GPU architectures.
 This process is necessary to uncover details about GPU components, such as the TLB and other shared resources, which can be exploited for covert communication.
@@
 Ahn et al.
-(Ahn et al., 2021b) initially reverse-engineer the structure of on-chip networks within modern GPUs and note that the hierarchical organization of the GPU leads to the sharing of interconnect bandwidth among adjacent cores.
-They then exploit the contention of the interconnect channel, leverage clock registers for sender-receiver synchronization and ultimately establish microarchitectural covert channels with a surprisingly high bandwidth of 24 Mbps.
+(Ahn et al., 2021b) initially reverse-engineer the structure of on-chip networks and find the hierarchical organization of the GPU leads to the sharing of interconnect bandwidth among adjacent cores.
+They then exploit the contention of the interconnect channel, leverage clock registers for sender-receiver synchronization and ultimately establish microarchitectural covert channels with a high bandwidth of 24 Mbps.
 Nayak et al.
@@
 In summary, both studies begin by identifying shared resources present in multi-GPU systems, such as the shared remote L2 cache (Dutta et al., 2023) and the shared L3 TLB (Zhang et al., 2023), and then create visible contention on these shared resources.
+Takeaway points.
+❶ Reverse engineering remains key to GPU attacks, highlighting both the limitations of security through obscurity and the lack of transparency from vendors that impedes systematic evaluation.
+❷ Attack bandwidths have significantly improved and are now comparable to early CPU-based covert channels, reflecting the growing sophistication of GPU exploitation.
+❸ Isolation mechanisms such as MIG may have flaws.
+Logical partitioning alone does not guarantee security and requires careful validation.
+❹ Attacks leveraging NVLink and PCIe show that GPU interconnects are an emerging and under-protected attack surface, particularly in AI datacenter environments.
+❺ ML workloads, especially DNNs, remain prime attack targets.
+Yet, the security of newer models such as LLMs and diffusion models across multi-GPU systems remains largely unexplored.
 5.2.3. Countermeasures
@@
 Zhan et al.(Zhan et al., 2022) identify the DVFS feature used in GPU as the underlying cause of an MMIO side-channel vulnerability.
-In scenarios where a victim employs an NVIDIA or AMD GPU without sharing, an attacker can monitor the victim’s activities and discern the visited webpages (i.e., a website fingerprinting attack).
+In scenarios where a victim employs an NVIDIA or AMD GPU, an attacker can monitor the victim’s activities and discern the visited webpages (i.e., a website fingerprinting attack).
 The key idea is capturing the fluctuating workload of a GPU during webpage rendering, which influences the MMIO signal it emits.
@@
 (Liang et al., 2022) conduct a similar far-field MMIO side-channel attack to infer the number of layers and their types, the number of kernels, sizes of layers and strides in an NN model.
+Takeaway points.MMIO channels are a consistent and critical leakage source, acting as a bridge between CPU and GPU for both local and remote side-channel attacks.
+❷ Access to internal sensors (e.g., temperature, power) varies by vendor.
+On platforms like Apple, where such access is unprivileged, this creates opportunities for exploitation.
+❸ Remote physical attacks are becoming increasingly feasible, with some able to succeed from several meters away, even through walls, without privileged access.
+This broadens the threat model to include fully remote, passive attackers.
 5.3.3. Countermeasures
@@
 (Sun et al., 2023) focus on searching sensitive targets of a CNN model, identifying proper parameters for timing and position, refining fault injection parameters, and ultimately reducing inference accuracy by an average of $69.1\%$.
-Santos et al.
-(Santos et al., 2021) propose a two-level fault injection framework combining the accuracy of Register Transfer Level (RTL) fault injection with the efficiency of software fault injection.
+dos Santos et al.
+(dos Santos et al., 2021) propose a two-level fault injection framework combining the accuracy of Register Transfer Level (RTL) fault injection with the efficiency of software fault injection.
 Overall, researchers have primarily leveraged software interfaces to inject faults, often utilizing DVFS or Rowhammer attacks.
@@
 The high parallelism and memory bandwidth inherent in GPUs potentially make them more vulnerable to Rowhammer attacks, as they can generate memory accesses faster than CPUs.
+Takeaway points.
+❶ Fault injection ‘can be triggered entirely through software by abusing out-of-spec voltage-frequency settings, exposing a lack of safeguards in GPU power management and its tight coupling with CPU controls.
+❷ While DNNs are generally resilient to random faults, targeted injection at specific layers or operations can significantly reduce accuracy.
+Achieving this, however, requires precise control over timing, location, and fault parameters.
+❸ The high memory bandwidth and parallelism of GPUs make them especially prone to Rowhammer attacks, as they can generate rapid memory access patterns that increase the risk of bit flips.
 5.4.2. Countermeasures
@@
 A similar strategy could also be applied to GPUs.
-6. Security of CPU TEEs
-In this section, we introduce typical attacks against CPU TEEs.
-As there already exist survey papers focusing on reviewing existing attacks on CPU TEEs, and the aim of this survey is to explore potential attacks on GPU TEEs, we simply categorize the most common attacks based on their technical details and existing surveys (Fei et al., 2021; Nilsson et al., 2020; Cerdeira et al., 2020; Li, 2022) here.
-6.1. Security of Intel SGX
-6.1.1. Controlled Channel and Address Translation-based Attacks
-Controlled channels attack, a concept introduced by Xu et al.
-(Xu et al., 2015) in 2015, exploits the privileges of the OS to monitor and manipulate page tables.
-This allows the attacker to uncover the memory access patterns of the enclave code, cf.
-also (Shinde et al., 2016).
-This attack is especially effective because, while Intel SGX enclaves protect sensitive computations from a compromised OS, the OS still manages memory and can observe page faults, thereby inferring details about the enclave’s operations.
-The SGX-Step (Van Bulck et al., 2017a) serves as a practical tool for, among others, executing such controlled channel attacks by configuring Advanced Programmable Interrupt Controller (APIC) timers, issuing interrupts, and tracking page-table entries.
-SGX-Step enables deterministic single-stepping of enclave execution, allowing for precise monitoring of memory access patterns over time.
-In fault-less page table attacks (Kim et al., 2019; Van Bulck et al., 2017b; Wang et al., 2017), attackers exploit the page table itself rather than relying on page faults.
-This allows them to infer the target’s memory access patterns by examining page table attributes and/or observing the caching behaviour of unprotected page table memory.
-However, these attacks are limited to page-level accuracy.
-Segmentation attacks (Gyselinck et al., 2018) allow attackers to deduce finer memory access patterns by manipulating the segmentation unit, but this only works on 32-bit enclaves, as segmentation is disabled in 64-bit enclaves (Fei et al., 2021).
-6.1.2. Cache-based Attacks
-Typical cache timing-channel attacks (in the non-SGX scenario) comprise four primary variants, including Evict+Time (Tromer et al., 2010), Evict+Reload (Gruss et al., 2015), Prime+Probe (Osvik et al., 2006), and Flush+Reload (Yarom and Falkner, 2014), all relying on the time difference between cache hits and misses.
-Recent studies (Götzfried et al., 2017; Moghimi et al., 2017; Schwarz et al., 2017; Brasser et al., 2017b; Dall et al., 2018; Moghimi et al., 2019) show that SGX enclaves are susceptible to the same cache attacks as any other software application.
-In fact, they may be even more vulnerable due to the enhanced attackers within SGX’s threat model.
-Several papers (Götzfried et al., 2017; Moghimi et al., 2017; Schwarz et al., 2017; Brasser et al., 2017b) exploit L1 cache and the shared LLC as the attack surface by employing Prime+Probe method.
-Dall et al.
-(Dall et al., 2018) take a slightly different approach by using Prime+Probe on Intel’s provisioning enclave, allowing Intel themselves to compromise the Enhanced Privacy ID (EPID) unlinkability property.
-Memjam (Moghimi et al., 2019) utilizes read-after-write false dependencies caused by the 4K aliasing of the L1 cache to conduct an Evict+Time-style attack.
-6.1.3. Branch Prediction Attacks
-Branch prediction attacks offer a more detailed level of utilization compared to address translation-based and cache-based attacks.
-They target the branch prediction unit, including the Branch Target Buffer (BTB(Lee et al., 2017; Chen et al., 2019), the Page History Table (PHT(Evtyushkin et al., 2018; Huo et al., 2020), and the Return Stack Buffer (RSB(Koruyeh et al., 2018).
-Attackers exploit the BTB to ascertain whether a particular branch instruction is taken by the victim.
-The PHT attack seeks to provoke collisions between branches of victim and attacker processes and subsequently infer the direction of the victim process branch.
-RSB attacks rely on contaminating the RSB to incite mis-speculation, consequently resulting in the leakage.
-6.1.4. Transient Execution Attacks
-Transient execution attacks, which exploit out-of-order and speculative execution of instructions (summarized in section 2.2), have also been studied in the context of SGX.
-The aforementioned BTB and RSB attacks (Chen et al., 2019; Koruyeh et al., 2018) can also fall into this class of attacks, stemming from Spectre (Kocher et al., 2020).
-Additionally, the Foreshadow attack (Van Bulck et al., 2018), which targets SGX by employing Meltdown-like techniques, breaches enclave execution confidentiality.
-While MDS attacks (Canella et al., 2019; Van Schaik et al., 2019; Schwarz et al., 2019) also leverage speculative and out-of-order execution, they primarily exploit information leakage from various implementation-specific and undocumented intermediary buffers of the targeted microarchitecture.
-SGAxe (Van Schaik et al., 2020) and Crosstalk (Ragab et al., 2021) extend MDS to extract private keys from SGX enclaves.
-6.1.5. Software-based Fault Injection and Side-Channel Attacks
-In fault injection attacks, the attacker disrupts the normal operation of a system by introducing malicious faults and then derives secret information from the faulty execution.
-Software-based fault attacks expand the threat model from a local attacker, who requires physical access to the target device, to a potentially remote attacker with only local code execution capabilities.
-Several studies (Murdock et al., 2020; Kenjar et al., 2020; Qiu et al., 2020) exploit privileged DVFS on x86 CPUs to reliably corrupt enclave computations, thereby compromising the confidentiality and integrity of enclaves.
-Chen and Oswald (Chen and Oswald, 2023) show that software fault attacks on SGX can also be mounted by pivoting to management chips on Supermicro server systems, and can even permanently damage the host CPUs.
-Apart from voltage scaling-based attacks, SGX-Bomb (Jang et al., 2017) leverages the Rowhammer bug to induce bit flips in enclave memory, resulting in failed data integrity checks and subsequent system lock-down.
-In addition to these fault injection attacks, Lipp et al.
-(Lipp et al., 2021) exploits a side-channel attack by leveraging unprivileged access to the Intel Running Average Power Limits interface.
-This interface reveals values linked to power consumption, creating a power side channel exploitable across Intel server, desktop, and laptop CPUs.
-Wang et al.
-(Wang et al., 2022b) show that power leakage can be measured through indirect channels such as frequency scaling, and Liu et al.
-(Liu et al., 2022) use the related feature of throttling to among others leak secrets from SGX enclaves.
-6.1.6. Hardware-based Attacks
-Apart from software-based fault injection attacks, Chen et al.
-(Chen et al., 2021; Chen and Oswald, 2023) explore hardware-level fault injection attacks on SGX enclaves.
-“VoltPillager” is a low-cost tool that injects messages on the Serial Voltage Identification (SVID) bus between the CPU and the voltage regulator, breaching the confidentiality and integrity of Intel SGX enclaves (Chen et al., 2021).
-Chen and Oswald later demonstrates that undervolting via the Power Management Bus (PMBus) can similarly break SGX enclave integrity, bypassing Intel’s countermeasures against previous software undervolting attacks (Murdock et al., 2020; Kenjar et al., 2020; Qiu et al., 2020).
-“MEMBUSTER” (Lee et al., 2020) introduces an off-chip attack that breaks SGX enclave confidentiality by snooping the memory bus (with high-spec, expensive lab equipment), extracting memory access patterns and increasing cache misses using techniques like critical page whitelisting and cache squeezing.
-6.1.7. Interface-based Attacks
-SGX enclave security can be compromised if they fail to strictly adhere to secure interfaces between trusted and untrusted code.
-Despite well-designed open-source Software Development Kits, Van Bulck et al.
-(Van Bulck et al., 2019) identify several sanitization vulnerabilities in Application Binary Interface (ABI) and API, leading to memory safety and side-channel vulnerabilities.
-Alder et al.
-(Alder et al., 2020) explores the attack surface affecting floating-point computations in SGX enclaves via ABI, showing that control and state registers are not always properly sanitized.
-Recently, there has been a focus on symbolic execution tools for SGX (Cloosters et al., 2020; Khandaker et al., 2020; Alder et al., 2024).
-Pandora (Alder et al., 2024) enables runtime-agnostic symbolic execution of the exact attested enclave binary and validates the enclave shielding runtime.
-Pandora can autonomously discover 200 new and 69 known vulnerable code locations across 11 different SGX shielding runtimes.
-6.1.8. Countermeasures
-In this section, we review and categorize the most relevant countermeasures based on both prior surveys (Nilsson et al., 2020; Fei et al., 2021) and potential solutions proposed in our summarized attack papers.
-(i) System/Microcode-level countermeasures: A CPU’s complex instructions are often managed by low-level software called microcode, rather than being entirely implemented in hardware.
-Thus, Intel can issue microcode updates, which are typically the most effective method.
-For instance, in response to DVFS-based attacks like Plundervolt (Murdock et al., 2020), Intel released microcode updates that fully disable the software voltage scaling interface via MSR $\mathtt{0x150}$.
-(ii) Compiler-based countermeasures: Beyond Intel’s solutions, compiler-based approaches are also effective.
-Deterministic multiplexing (Shinde et al., 2016) ensures that page-fault access patterns remain consistent regardless of input values by proactively accessing all data and code in a predetermined sequence.
-“T-SGX” (Shih et al., 2017) use Intel’s Transactional Synchronization Extensions (TSX) to redirect exceptions and interrupts to a specific page, terminating the program if a page fault is detected.
-Chen et al.
-(Chen et al., 2017) propose shielded execution with reliable time measurement to check the execution time of target programs.
-“Cloak” (Gruss et al., 2017) uses Hardware Transactional Memory (HTM) to protect against malicious cache observations.
-Similarly, glitch detection (Spensky et al., 2021) can help mitigate DVFS-based fault injection attacks.
-(iii) Randomization: “SGX-Shield” (Seo et al., 2017) leverages ASLR to create a secure in-enclave loader that secretly randomizes memory space layout, hiding sensitive operations from attackers.
-Hosseinzadeh et al.
-(Hosseinzadeh et al., 2018) propose runtime control flow randomization, implemented as compiler extensions on llvm.
-“DR.SGX” (Brasser et al., 2019a) disrupts memory observations by permuting data locations.
-(iv) Application/Source code design: Finally, developers can design enclave applications in a side-channel-resistant manner.
-Early works like “OBFUSCURO” (Ahmad et al., 2019) enforce operations with Oblivious RAM (ORAM) protection to prevent attackers from inferring access patterns.
-Several schemes (Ohrimenko et al., 2016; Poddar et al., 2020; Wang et al., 2022a) utilize oblivious primitives, implemented in x86_64 assembly, to design data-oblivious applications.
-These primitives operate exclusively on registers, which are private to the processor, ensuring that the contents hidden from code outside the enclave and any register-to-register data manipulation is data-oblivious.
-6.2. Security of Arm TrustZone
-6.2.1. Implementation Bugs
-Cerdeira et al.
-(Cerdeira et al., 2020) examine bug reports obtained from public Common Vulnerabilities and Exposures (CVE) databases and vendor bulletin reports, categorizing them into validation bugs, functional bugs, and extrinsic bugs.
-These bugs are pervasive and often exploited to escalate privileges, enabling attackers to fully hijack kernels in devices from Qualcomm (Beniamini, 2016) and Huawei (Shen, 2015) or compromise client applications like Samsung Pay (Komaromy, 2018).
-6.2.2. Cache-based Attacks
-Given that the cache is shared between NW and SW in TrustZone-enabled processors, several studies capitalize on the contention arising from this cache coherency to extract sensitive information (Zhang et al., 2016c; Guanciale et al., 2016; Lipp et al., 2016; Zhang et al., 2016b; Cho et al., 2018).
-Several authors (Lipp et al., 2016; Zhang et al., 2016b; Cho et al., 2018) use the Prime+Probe method to monitor the cache activities within TrustZone.
-6.2.3. Branch Prediction Attacks
-Similar to the attacks on Intel SGX mentioned earlier, the branch predictor can also be exploited to target TrustZone.
-Since the BTB is shared between NW and SW, techniques like Prime+Probe can be used to disclose secure information to NW.
-For example, Ryan (Ryan, 2019) successfully recovered a 256-bit private key from Qualcomm’s hardware-backed keystore.
-6.2.4. Software-based Fault Injection Attacks
-Several studies have also expanded software-based fault injection attacks to TrustZone or TrustZone-based systems.
-“CLKSCREW” (Tang et al., 2017) leverages manipulation on DVFS to induce faulty computations to breach TrustZone hardware-enforced boundaries, extract secret keys, and bypass code signing operations.
-Rowhammer-based attacks have also been adapted to undermine TrustZone security.
-Given that overclocking typically involves running the processor at high frequencies, which can be relatively easy to detect and prevent (e.g., through hardware frequency locking), “VoltJockey” (Qiu et al., 2020) adopts a different strategy by manipulating voltages rather than frequencies to generate hardware faults on the victim cores.
-Carru (Carru, 2017) use bitflips to bypass security measures and access secure memory, enabling the extraction of RSA keys and undermining the secure storage provided by TrustZone.
-6.2.5. Countermeasures
-In this section, we review and categorize countermeasures based on both previous survey (Cerdeira et al., 2020) and potential solutions from the summarized attack papers.
-In addressing the architectural defences of Arm TrustZone, some research efforts (Sun et al., 2015; Brasser et al., 2019b) enhance isolation granularity between TEE components and reduce the amount of code running in SW, thereby lowering the risk of severe privilege escalation attacks within SW.
-Countermeasures (Jang et al., 2015; Jang and Kang, 2018) tackle the weaknesses of inadequate/weak authentication when accessing TEE resources from NW, and potentially insecure shared memory used for data exchange across the NW-SW boundary.
-Unlike Intel SGX, TrustZone lacks built-in on-chip memory encryption, prompting some works (Zhang et al., 2016a; Yun and Zhong, 2019) to design encryption strategies to bridge this gap.
-To further strengthen the verification of both TEE and TA binaries’ integrity and identity, several studies (Ferraiuolo et al., 2017; Lentz et al., 2018) introduced additional trusted computing primitives, such as enhanced remote attestation and sealed storage.
-To address implementation bugs, “RustZone” (Evenchick, 2018) extends TrustZone by implementing TAs using Rust, which inherently provides memory safety and thread safety.
-Cache side-channel attacks can be mitigated by carefully implementing cryptographic algorithms in software (Guanciale et al., 2016) or using dedicated hardware (Lipp et al., 2016).
-Cache maintenance techniques can also be effective in mitigating leakage.
-For example, in terms of BTB attacks, flushing the shared microarchitectural structures when transferring between NW and SW can prevent attacks from targeting keys through the BTB.
-However, as observed in the Nexus 5X firmware by (Ryan, 2019), the L1 cache is not flushed, making Prime+Probe-based attacks still feasible.
-Countermeasures against DVFS-based fault injection attacks are similar to those used for SGX.
-In comparison to other attacks on TrustZone, Rowhammer attacks are relatively straightforward to mitigate.
-They rely on inducing bit flips in adjacent memory rows, making them feasible only if non-secure memory is situated next to a memory row within SW.
-If secure memory resides on a separate memory storage device, Rowhammer attacks become impractical.
-6.3. Security of AMD SEV
-6.3.1. Memory Encryption Issues
-Since the release of AMD SEV (AMD, 2023), numerous attacks have targeted its early memory security mechanisms (Hetzelt and Buhren, 2017; Buhren et al., 2017; Du et al., 2017; Morbitzer et al., 2018).
-Hetzelt and Buhren (Hetzelt and Buhren, 2017) utilize a return-oriented programming-like attack to exploit an unencrypted VM control block, enabling arbitrary reading and writing of encrypted memory in the guest VM.
-Such vulnerabilities have been addressed in SEV-ES (Kaplan, 2017).
-Du et al.
-(Du et al., 2017) leverage Electronic Code Book (ECB) mode in memory encryption to conduct a chosen-plaintext attack via an HTTP server installed on the guest VM.
-Sevurity (Wilke et al., 2020) exploits vulnerabilities in the Xor-Encrypt-Xor (XEX) mode to insert arbitrary 2-byte instructions into encrypted memory, while Morbitzer et al.
-(Morbitzer et al., 2018) breach confidentiality by manipulating nested page tables, altering the virtual memory of the guest VMs.
-Most of these attacks arise from the lack of integrity protection, but SEV-SNP (Sev-Snp, 2020) addressed them through the Reverse Map Table.
-Conversely, “CIPHERLEAKS” (Li et al., 2021b) allows a privileged adversary to infer guest VM execution states or recover certain plaintext by monitoring changes in the ciphertext of the victim VM.
-The authors highlight that AMD has confirmed that SEV-SNP is also vulnerable to the CIPHERLEAKS attack (Li, 2022).
-Additionally, “CROSSLINE” and its variants by Li et al.
-(Li et al., 2021a; Li, 2022) extract the memory content of victim VMs by exploiting improper use of the Address Space Identifier (ASID) for controlling VM accesses to encrypted memory pages, cache lines, and TLB entries.
-They also compromise the integrity and confidentiality of SEV VMs by poisoning TLB entries (Li et al., 2021c).
-It is noted that SEV-SNP is expected to address these issues, including TLB misuse.
-6.3.2. Software-based Fault Injection Attacks
-Morbitzer et al.
-(Morbitzer et al., 2021) exploit the system’s lack of memory integrity protection to inject faults and execute arbitrary code within SEV VMs through I/O channels, using the hypervisor to locate and trigger the execution of encrypted payloads.
-However, SEV-SNP (Sev-Snp, 2020) can mitigate this issue through Reverse Map Table (RMP).
-Buhren et al.
-(Buhren et al., 2021) present a voltage fault injection attack enabling the execution of custom payloads on AMD secure processors, subsequently decrypting a VM’s memory by deploying custom SEV firmware.
-“CacheWarp” (Zhang et al., 2024) presents a new software-based fault attack targeting AMD SEV-ES and SEV-SNP, exploiting the possibility to revert modified cache lines of guest VMs to their previous (stale) state at the architectural level.
-This attack only requires interrupting the VM at a point chosen by the attacker to invalidate modified cache lines without writing them back to memory.
-More recently, Schlüter et al.
-(Schlüter et al., 2024a) introduce a novel attack where the hypervisor injects well-crafted, malicious #$\mathtt{\#VC}$ (a new exception by AMD facilitating communication between the VM and untrusted hypervisor) into a victim VM’s CPU, enabling the attacker to induce arbitrary behaviour in the VM.
-In another related attack, “HECKLER” (Schlüter et al., 2024b) shows that the hypervisor can inject malicious non-timer interrupts to compromise VMs.
-They leverage interrupt handlers with global effects, allowing manipulation of a VM’s register states to alter data and control flow.
-In general, SEV-SNP remains vulnerable to such software-based fault injection attacks.
-6.3.3. Interrupt-driven Attacks
-Similar to SGX-Step framework (Van Bulck et al., 2017a), SEV-Step (Wilke et al., 2024) introduces reliable single-stepping against SEV VMs.
-In addition to this functionality, SEV-Step facilitates access to common attack techniques, such as page fault tracking and cache attacks targeting SEV.
-6.3.4. Countermeasures
-Most memory-related attacks have been addressed in recent AMD product updates.
-Here, we focus on countermeasures against fault injection attacks.
-For malicious voltage drops or glitches, similar to countermeasures used for SGX and TrustZone, hardware-level voltage monitoring circuits and software-based detection mechanisms can be implemented.
-To counter fault attacks that enable fine-grained memory write suppression through specific hypervisor instructions, restricting the use of those instructions at the hardware level or implementing compiler-level solutions can ensure correct memory operations.
-For attacks that inject malicious interrupts, one can block or detect the injections by monitoring external interrupts or disabling the interrupt handler.
-7. Conclusions on future directions in GPU TEE
-In this section, based on the relevant work discussed before, we sketch potential attacks that might be relevant to GPU TEEs in the future.
-We note that as GPU TEEs are not widely commercially available, it is not yet possible to practically evaluate these attack vectors, but believe that according vulnerabilities might be discovered in the future.
+6. Discussions on future directions in GPU TEE
+In Section 4, we provided a taxonomy of existing GPU TEE designs, highlighting their architectural differences, security assumptions, and threat models.
+Building on this foundation, Sections 5 and the supplemental material (due to the limited space of this article) reviewed and compared a broad range of attacks targeting standalone GPUs and established CPU TEEs such as Intel SGX, Arm TrustZone, and AMD SEV.
+In this section, we extrapolate from these findings to identify potential attack vectors that may arise in the context of CPU TEEs, mapping them to the specific components and trust boundaries involved in these systems.
+We note that as GPU TEEs are not widely commercially available, it is not yet possible to practically evaluate these attack vectors, but we believe that according vulnerabilities might be discovered in the future.
 6.1. Potential Attacks against x86-based GPU TEE Prototype
@@
 6.1.1. Existing Attacks against Intel SGX
-Given that the essential GPU resource management functions within the GPU driver reside within the TEE enclave, the majority of attacks outlined in Section 6.1 could potentially target this enclave to glean secrets (cf.
+Given that the essential GPU resource management functions within the GPU driver reside within the TEE enclave, the majority of attacks outlined in the supplemental material could potentially target this enclave to glean secrets (cf.
 attack surface ➀ in Figure 6).
@@
 Due to the absence of protection on GPU hardware, attackers with physical access to GPU hardware could directly manipulate GPU voltage, injecting faults into running GPU kernels.
-For instance, “VoltPillager” (Chen et al., 2021) can inject messages on the SVID bus between the CPU and the voltage regulator on the motherboard, enabling precise control over CPU core voltage.
+For instance, “VoltPillager” (Chen et al., 2021) can inject messages on the Serial Voltage Identification (SVID) bus between the CPU and the voltage regulator on the motherboard, enabling precise control over CPU core voltage.
 Ultimately, this jeopardizes the confidentiality and integrity of Intel SGX enclaves.
@@
 Consequently, direct physical access to GPU memory can compromise sensitive data.
-Further, cold boot attacks (Halderman et al., [n.
-d.]) could be relevant, if an adversary can ensure (e.g., by cooling) that TEE-protected memory contents are retained and can be readout after a reset of the GPU.
+Further, cold boot attacks (Halderman et al., ) could be relevant, if an adversary can ensure (e.g., by cooling) that TEE-protected memory contents are retained and can be readout after a reset of the GPU.
 Additionally, as PCIe interconnects are exposed, injecting malicious PCIe packets via specialized hardware is feasible.
@@
 attack surface ➀ in Figure 7).
-For example, the “WeSee” attack (Schlüter et al., 2024a) targets AMD SEV-SNP used in NVIDIA H100-based GPU TEEs, allowing an attacker to perform arbitrary memory writes by faking a #$\mathtt{\#VC}$ for MMIO reads.
+For example, the “WeSee” attack (Schlüter et al., 2024a) targets AMD SEV-SNP used in NVIDIA H100-based GPU TEEs, allowing an attacker to perform arbitrary memory writes by faking a $\mathtt{\#VC}$ for MMIO reads.
 This capability would enabls the attacker to modify commands and data sent to the GPU.
@@
 On the GPU side.
-The H100 whitepaper (NVIDIA, 2022b) lacks detail about voltage and frequency management.
+The H100 whitepaper (NVIDIA, 2022c) lacks detail about voltage and frequency management.
 Assuming an attacker has physical access, they could potentially execute attacks like “VoltPillager” (Chen et al., 2021) to inject malicious SVID packets or equivalent into the relevant GPU card’s bus, thereby altering the GPU voltage.
@@
 6.2.3. Microarchitectural Side-Channel and Covert-Channel Attacks on GPU
-The H100 whitepaper (NVIDIA, 2022b) emphasizes the enhanced security features of the Hopper architecture for confidential computing, particularly in collaborative multi-party computing scenarios, namely single tenant with multiple GPUs and multiple tenants with a single GPU.
+The H100 whitepaper (NVIDIA, 2022c) emphasizes the enhanced security features of the Hopper architecture for confidential computing, particularly in collaborative multi-party computing scenarios, namely single tenant with multiple GPUs and multiple tenants with a single GPU.
 However, vulnerabilities discussed in Section 5.2 may potentially lead to security breaches on the NVIDIA H100 GPU.
-[…]
+Single tenant with multiple GPUs.
+In the scenario of single tenant with multiple GPUs, a single VM is concurrently assigned multiple physically independent GPUs, each connected via NVLink.
+Recall that Dutta et al.
+(Dutta et al., 2023) demonstrate that an attacker on one GPU could induce contention on the L2 cache of another GPU via NVLink, enabling a covert channel and Prime+Probe attacks across GPUs.
+However, such attacks may face challenges in the NVIDIA H100 GPU model due to two key factors (cf.
+attack surface ➁ in Figure 7).
+Firstly, the attacker’s process needs access to a remote GPU’s memory through NVLink, which may be restricted in the threat model of the NVIDIA H100 GPU TEE.
+Thus, the challenge lies in how to enable the spy process to run within one of the assigned GPUs and have access to another GPU’s memory.
+Secondly, the H100 GPU features a 50 MB L2 cache, $1.25{\times}$ larger than the A100 GPU.
+Nevertheless, a larger L2 cache with more sets and ways can make Prime+Probe attacks more difficult, because the attacker needs to fill and monitor a larger number of cache lines.
 Multiple tenants with a single GPU.
@@
 StrongBox serves as a suitable foundation as it was recently introduced and does not rely on specific Arm endpoint features or necessitate hardware modifications to GPU or CPU chips.
-Notably, while secure GPU computing on Arm or Intel platforms may not align with Arm’s CCA realm-style architecture (Wang et al., 2024b), we omit insights regarding this, as it closely resembles that of NVIDIA Hopper H100-based GPU TEE.
+Notably, while secure GPU computing on Arm or Intel platforms may not align with Arm’s CCA realm-style architecture (Wang et al., 2024a), we omit insights regarding this, as it closely resembles that of NVIDIA Hopper H100-based GPU TEE.
 In Figure 8, the attacker gains control over the kernel and the entire GPU software stack, including the GPU driver and runtime.
@@
 6.3.1. Existing Attacks against Arm TrustZone
-The attacks summarized in Section 6.2 that target TrustZone are also applicable to this Arm-based GPU TEE prototype.
+The attacks summarized in the supplemental material that target TrustZone are also applicable to this Arm-based GPU TEE prototype.
 The impact of these side-channel attacks on TrustZone indirectly affects Arm-based GPU TEE (cf.
@@
 However, establishing covert channels between malicious and victim tasks is unlikely, as GPU tasks are executed sequentially on Arm endpoint GPUs (Deng et al., 2022).
-(ii) An attacker may monitor cache activity to extract AES keys from TrustZone (Guanciale et al., 2016; Zhang et al., 2016b; Lipp et al., 2016), thus compromising the confidentiality of data transferred between the CPU and GPU.
+(ii) An attacker may monitor cache activity to extract AES keys from TrustZone (Guanciale et al., 2016; Zhang et al., 2016c; Lipp et al., 2016), thus compromising the confidentiality of data transferred between the CPU and GPU.
 6.3.2. Microarchitectural Side-Channel and Covert-Channel Attacks on Integrated GPU
@@
 However, challenges arise due to undocumented LLC/system-level cache behavior and the asymmetric view of cache utilization between the CPU and GPU in the Arm architecture.
-(ii) The existing threat model of Arm-based GPU TEEs may not effectively mitigate attacks exploiting data-dependent DRAM traffic and cache utilization induced by compression to reveal user pixels in the normal world (Wang et al., 2024a).
+(ii) The existing threat model of Arm-based GPU TEEs may not effectively mitigate attacks exploiting data-dependent DRAM traffic and cache utilization induced by compression to reveal user pixels in the normal world (Wang et al., 2024b).
 Further testing is needed to assess the success rate of such attacks on Arm-based GPU TEEs.
@@
 Most existing Arm-based GPU TEE prototypes primarily focus on delivering hardware-enforced isolation to protect sensitive code and data, rather than employing memory encryption mechanisms like Intel SGX’s MEE.
-Consequently, direct physical access to GPU memory (again also possibly via cold boot attacks (Halderman et al., [n.
-d.])) could potentially compromise sensitive data (cf.
+Consequently, direct physical access to GPU memory (again also possibly via cold boot attacks (Halderman et al., )) could potentially compromise sensitive data (cf.
 attack surface ➃ in Figure 8).
@@
 Additionally, assuming an attacker gains physical access to the voltage regulator or clock control module within the System on Chip (SoC), this prompts further investigation into whether hardware-based attacks targeting these components (Chen et al., 2021) are also applicable to Arm GPUs.
+6.4. Comparison of Attacks on Different GPU TEEs
+x86-based GPU TEE prototypes, NVIDIA Hopper H100, and Arm-based GPU TEEs adopt distinct trust models, hardware assumptions, and protection strategies.
+x86-based designs (e.g., Graviton, HIX, HETEE) typically build on Intel SGX, placing only essential GPU driver components inside the enclave to minimize the TCB.
+These prototypes often lack GPU-side memory encryption and leave out DVFS and power management from protection, making them susceptible to software-based fault injection and physical attacks (e.g., cold boot, SVID manipulation).
+In contrast, NVIDIA Hopper H100 integrates vendor-supported confidential computing features, including memory encryption and compatibility with VM-based TEEs.
+While this offers system-wide isolation, it inherits known vulnerabilities of VM-based TEEs, such as software-based fault injection (e.g., WeSee) and cache manipulation (e.g., CacheWarp).
+Furthermore, the limited public documentation presents challenges for third-party validation and systematic defense design.
+Arm-based GPU TEE designs, such as StrongBox, use TrustZone to provide software-based isolation without requiring hardware modifications.
+This approach improves portability and ease of deployment but sacrifices granularity.
+These systems typically lack hardware-enforced memory encryption and rely on unified memory and shared caches, which can expose sensitive data to physical and side-channel attacks.
+Operating in resource-constrained environments (e.g., edge devices), Arm-based TEEs tend to have a smaller TCB but may not adequately defend against DVFS abuse or microarchitectural leakage.
+In summary, existing x86-based prototypes prioritize minimal hardware changes but depend on SGX’s integrity; NVIDIA H100 adopts holistic, hardware-enhanced isolation but inherits the complexity and weaknesses of VM-based TEEs; and Arm-based TEEs favor lightweight integration at the cost of fine-grained protection.
+From an attack perspective, x86 prototypes are vulnerable at the interface between CPU TEEs and unprotected GPUs, NVIDIA’s model is exposed through VM-level fault channels and microarchitectural sharing, and Arm designs are weakest at the hardware and cache/memory system level.
+These architectural differences reflect distinct trade-offs and highlight the need for architecture-aware defenses tailored to each GPU TEE design.
 7. Conclusion
@@
 This research is partially funded by the Engineering and Physical Sciences Research Council (EPSRC) under grants EP/X03738X/1, EP/V000454/1, and EP/R012598/1. The results feed into DsbDtech.
+Appendix A Security of CPU TEEs
+In this online appendix, we introduce typical attacks against CPU TEEs.
+As there already exist survey papers focusing on reviewing existing attacks on CPU TEEs, and the aim of this survey is to explore potential attacks on GPU TEEs, we simply categorize the most common attacks based on their technical details and existing surveys (Fei et al., 2021; Nilsson et al., 2020; Cerdeira et al., 2020; Li, 2022) here.
+We provide a brief summary of these attacks regarding attack surfaces and countermeasures in table 4.
+Table 4. Summary of attacks on CPU TEEs
+CPU TEE Attacks Ref Attack Surfaces Countermeasures
+Arch. Threat Model PT SE $C$ DRAM BTB PHT RSB IFC HC Bug INT
+Trusted Untrusted
+SGX CPU package, EPC, MEE OS/Kernel AT (Xu et al., 2015) Obfuscation, randomization (Brasser et al., 2019a; Ahmad and Bates, 2019)
+(Van Bulck et al., 2017b; Kim et al., 2019)
+(Gyselinck et al., 2018)
+Cache (Brasser et al., 2017b; Dall et al., 2018; Götzfried et al., 2017; Moghimi et al., 2017) Obfuscation, randomization, anomaly detection (Brasser et al., 2019a; Ahmad and Bates, 2019)
+(Schwarz et al., 2017)
+BP (Lee et al., 2017) Obfuscation (Ahmad and Bates, 2019)
+(Chen et al., 2019)
+(Evtyushkin et al., 2018; Huo et al., 2020)
+(Koruyeh et al., 2018)
+(Van Bulck et al., 2018)
+TE (Van Bulck et al., 2018) Microcode updates from Intel
+(Van Schaik et al., 2020; Ragab et al., 2021)
+SwFIA (Murdock et al., 2020; Kenjar et al., 2020; Qiu et al., 2020) Microcode updates from Intel, restricted access to registers, ECC
+(Lipp et al., 2021; Wang et al., 2022b; Liu et al., 2022)
+(Jang et al., 2017)
+Hw (Chen et al., 2021; Chen and Oswald, 2023; Lee et al., 2020) Fuse-level lockouts, access pattern obfuscation, harden BMC/PMU firmware
+Interface (Van Bulck et al., 2019; Alder et al., 2020; Cloosters et al., 2020; Khandaker et al., 2020; Alder et al., 2024) State and register sanitization, minimize the ABI surface, automated runtime analysis
+TrustZone SW, Secure monitor NW OS IB (Beniamini, 2016; Shen, 2015; Komaromy, 2018) Software-only verification (Ferraiuolo et al., 2017; Jin et al., 2022)
+Cache (Zhang et al., 2016a; Guanciale et al., 2016; Lipp et al., 2016; Zhang et al., 2016c; Cho et al., 2018) Cache isolation (Dessouky et al., 2021)
+BP (Ryan, 2019) Flush BTB
+SwFIA (Tang et al., 2017; Qiu et al., 2020; Carru, 2017) Restrict access to DVFS interfaces
+SEV PSP, MEE OS/ Hypervisor ME (Hetzelt and Buhren, 2017; Buhren et al., 2017; Du et al., 2017; Morbitzer et al., 2018; Wilke et al., 2020; Li et al., 2021b) Mitigated in AMD SEV-SNP
+(Morbitzer et al., 2018)
+(Li et al., 2021a)
+SwFIA (Morbitzer et al., 2021; Zhang et al., 2024) Microcode update from AMD
+ID (Schlüter et al., 2024a, b; Wilke et al., 2024) Isolate interrupt handlers, detect single-stepping
+Attacks contains Address Translation (AT), Cache, Branch Prediction (BP), Transient Execution (TE), Software-based Fault Injection (SwFIA), Interface, Hardware (Hw), Implementation Bug (IB), Memory Encryption (ME), Interrupt-driven (ID).
+Attack Surfaces includes Page Table (PT), Segmentation (SE), CPU Cache (CC), Branch Target Buffer (BTB), Page History Table (PHT), Return Stack Buffer (RSB), Interface (IFC, containing DVFS interfaces, enclave interfaces, etc.), Interrupt (INT, interrupt controller/interrupt handling logic).
+A.1. Security of Intel SGX
+A.1.1. Controlled Channel and Address Translation-based Attacks
+Controlled channels attack, a concept introduced by Xu et al.
+(Xu et al., 2015) in 2015, exploits the privileges of the OS to monitor and manipulate page tables.
+This allows the attacker to uncover the memory access patterns of the enclave code, cf.
+also (Shinde et al., 2016).
+This attack is especially effective because, while Intel SGX enclaves protect sensitive computations from a compromised OS, the OS still manages memory and can observe page faults, thereby inferring details about the enclave’s operations.
+SGX-Step (Van Bulck et al., 2017a) serves as a practical tool for, among others, executing such controlled channel attacks by configuring Advanced Programmable Interrupt Controller (APIC) timers, issuing interrupts, and tracking page-table entries.
+SGX-Step enables deterministic single-stepping of enclave execution, allowing for precise monitoring of memory access patterns over time.
+In fault-less page table attacks (Kim et al., 2019; Van Bulck et al., 2017b; Wang et al., 2017), attackers exploit the page table itself rather than relying on page faults.
+This allows them to infer the target’s memory access patterns by examining page table attributes and/or observing the caching behaviour of unprotected page table memory.
+However, these attacks are limited to page-level accuracy.
+Segmentation attacks (Gyselinck et al., 2018) allow attackers to deduce finer memory access patterns by manipulating the segmentation unit, but this only works on 32-bit enclaves, as segmentation is disabled in 64-bit enclaves (Fei et al., 2021).
+A.1.2. Cache-based Attacks
+Typical cache timing-channel attacks (in the non-SGX scenario) comprise four primary variants, including Evict+Time (Tromer et al., 2010), Evict+Reload (Gruss et al., 2015), Prime+Probe (Osvik et al., 2006), and Flush+Reload (Yarom and Falkner, 2014), all relying on the time difference between cache hits and misses.
+Recent studies (Götzfried et al., 2017; Moghimi et al., 2017; Schwarz et al., 2017; Brasser et al., 2017b; Dall et al., 2018; Moghimi et al., 2019) show that SGX enclaves are susceptible to the same cache attacks as any other software application.
+In fact, they may be even more vulnerable due to the enhanced attackers within SGX’s threat model.
+Several papers (Götzfried et al., 2017; Moghimi et al., 2017; Schwarz et al., 2017; Brasser et al., 2017b) exploit L1 cache and the shared LLC as the attack surface by employing Prime+Probe method.
+Dall et al.
+(Dall et al., 2018) take a slightly different approach by using Prime+Probe on Intel’s provisioning enclave, allowing Intel themselves to compromise the Enhanced Privacy ID (EPID) unlinkability property.
+Memjam (Moghimi et al., 2019) utilizes read-after-write false dependencies caused by the 4K aliasing of the L1 cache to conduct an Evict+Time-style attack.
+A.1.3. Branch Prediction Attacks
+Branch prediction attacks offer a more detailed level of utilization compared to address translation-based and cache-based attacks.
+They target the branch prediction unit, including the Branch Target Buffer (BTB(Lee et al., 2017; Chen et al., 2019), the Page History Table (PHT(Evtyushkin et al., 2018; Huo et al., 2020), and the Return Stack Buffer (RSB(Koruyeh et al., 2018).
+Attackers exploit the BTB to ascertain whether a particular branch instruction is taken by the victim.
+The PHT attack seeks to provoke collisions between branches of victim and attacker processes and subsequently infer the direction of the victim process branch.
+RSB attacks rely on contaminating the RSB to incite mis-speculation, consequently resulting in the leakage.
+A.1.4. Transient Execution Attacks
+Transient execution attacks, which exploit out-of-order and speculative execution of instructions, have also been studied in the context of SGX.
+The aforementioned BTB and RSB attacks (Chen et al., 2019; Koruyeh et al., 2018) can also fall into this class of attacks, stemming from Spectre (Kocher et al., 2020).
+Additionally, the Foreshadow attack (Van Bulck et al., 2018), which targets SGX by employing Meltdown-like techniques, breaches enclave execution confidentiality.
+While MDS attacks (Canella et al., 2019; Van Schaik et al., 2019; Schwarz et al., 2019) also leverage speculative and out-of-order execution, they primarily exploit information leakage from various implementation-specific and undocumented intermediary buffers of the targeted microarchitecture.
+SGAxe (Van Schaik et al., 2020) and Crosstalk (Ragab et al., 2021) extend MDS to extract private keys from SGX enclaves.
+A.1.5. Software-based Fault Injection and Side-Channel Attacks
+In fault injection attacks, the attacker disrupts the normal operation of a system by introducing malicious faults and then derives secret information from the faulty execution.
+Software-based fault attacks expand the threat model from a local attacker, who requires physical access to the target device, to a potentially remote attacker with only local code execution capabilities.
+Several studies (Murdock et al., 2020; Kenjar et al., 2020; Qiu et al., 2020) exploit privileged DVFS on x86 CPUs to reliably corrupt enclave computations, thereby compromising the confidentiality and integrity of enclaves.
+Chen and Oswald (Chen and Oswald, 2023) show that software fault attacks on SGX can also be mounted by pivoting to management chips on Supermicro server systems, and can even permanently damage the host CPUs.
+Apart from voltage scaling-based attacks, SGX-Bomb (Jang et al., 2017) leverages the Rowhammer bug to induce bit flips in enclave memory, resulting in failed data integrity checks and subsequent system lock-down.
+In addition to these fault injection attacks, Lipp et al.
+(Lipp et al., 2021) exploits a side-channel attack by leveraging unprivileged access to the Intel Running Average Power Limits interface.
+This interface reveals values linked to power consumption, creating a power side channel exploitable across Intel server, desktop, and laptop CPUs.
+Wang et al.
+(Wang et al., 2022b) show that power leakage can be measured through indirect channels such as frequency scaling, and Liu et al.
+(Liu et al., 2022) use the related feature of throttling to among others leak secrets from SGX enclaves.
+A.1.6. Hardware-based Attacks
+Apart from software-based fault injection attacks, Chen et al.
+(Chen et al., 2021; Chen and Oswald, 2023) explore hardware-level fault injection attacks on SGX enclaves.
+“VoltPillager” is a low-cost tool that injects messages on the SVID bus between the CPU and the voltage regulator, breaching the confidentiality and integrity of Intel SGX enclaves (Chen et al., 2021).
+Chen and Oswald later demonstrates that undervolting via the Power Management Bus (PMBus) can similarly break SGX enclave integrity, bypassing Intel’s countermeasures against previous software undervolting attacks (Murdock et al., 2020; Kenjar et al., 2020; Qiu et al., 2020).
+“MEMBUSTER” (Lee et al., 2020) introduces an off-chip attack that breaks SGX enclave confidentiality by snooping the memory bus (with high-spec, expensive lab equipment), extracting memory access patterns and increasing cache misses using techniques like critical page whitelisting and cache squeezing.
+A.1.7. Interface-based Attacks
+SGX enclave security can be compromised if they fail to strictly adhere to secure interfaces between trusted and untrusted code.
+Despite well-designed open-source Software Development Kits, Van Bulck et al.
+(Van Bulck et al., 2019) identify several sanitization vulnerabilities in Application Binary Interface (ABI) and API, leading to memory safety and side-channel vulnerabilities.
+Alder et al.
+(Alder et al., 2020) explores the attack surface affecting floating-point computations in SGX enclaves via ABI, showing that control and state registers are not always properly sanitized.
+Recently, there has been a focus on symbolic execution tools for SGX (Cloosters et al., 2020; Khandaker et al., 2020; Alder et al., 2024).
+Pandora (Alder et al., 2024) enables runtime-agnostic symbolic execution of the exact attested enclave binary and validates the enclave shielding runtime.
+Pandora can autonomously discover 200 new and 69 known vulnerable code locations across 11 different SGX shielding runtimes.
+A.1.8. Countermeasures
+In this section, we review and categorize the most relevant countermeasures based on both prior surveys (Nilsson et al., 2020; Fei et al., 2021) and potential solutions proposed in our summarized attack papers.
+(i) System/Microcode-level countermeasures: A CPU’s complex instructions are often managed by low-level software called microcode, rather than being entirely implemented in hardware.
+Thus, Intel can issue microcode updates, which are typically the most effective method.
+For instance, in response to DVFS-based attacks like Plundervolt (Murdock et al., 2020), Intel released microcode updates that fully disable the software voltage scaling interface via MSR $\mathtt{0x150}$.
+(ii) Compiler-based countermeasures: Beyond Intel’s solutions, compiler-based approaches are also effective.
+Deterministic multiplexing (Shinde et al., 2016) ensures that page-fault access patterns remain consistent regardless of input values by proactively accessing all data and code in a predetermined sequence.
+“T-SGX” (Shih et al., 2017) use Intel’s Transactional Synchronization Extensions (TSX) to redirect exceptions and interrupts to a specific page, terminating the program if a page fault is detected.
+Chen et al.
+(Chen et al., 2017) propose shielded execution with reliable time measurement to check the execution time of target programs.
+“Cloak” (Gruss et al., 2017) uses Hardware Transactional Memory (HTM) to protect against malicious cache observations.
+Similarly, glitch detection (Spensky et al., 2021) can help mitigate DVFS-based fault injection attacks.
+(iii) Randomization: “SGX-Shield” (Seo et al., 2017) leverages ASLR to create a secure in-enclave loader that secretly randomizes memory space layout.
+Hosseinzadeh et al.
+(Hosseinzadeh et al., 2018) propose runtime control flow randomization, implemented as compiler extensions on llvm.
+“DR.SGX” (Brasser et al., 2019a) disrupts memory observations by permuting data locations.
+(iv) Application/Source code design: Early works like “OBFUSCURO” (Ahmad and Bates, 2019) enforce operations with Oblivious RAM (ORAM) protection to prevent attackers from inferring access patterns.
+Several schemes (Ohrimenko et al., 2016; Poddar et al., 2020; Wang et al., 2022a) utilize oblivious primitives, implemented in x86_64 assembly, to design data-oblivious applications.
+A.2. Security of Arm TrustZone
+A.2.1. Implementation Bugs
+Cerdeira et al.
+(Cerdeira et al., 2020) examine bug reports obtained from public Common Vulnerabilities and Exposures (CVE) databases and vendor bulletin reports, categorising them into validation bugs, functional bugs, and extrinsic bugs.
+These bugs are pervasive and often exploited to escalate privileges, enabling attackers to fully hijack kernels in devices from Qualcomm (Beniamini, 2016) and Huawei (Shen, 2015) or compromise client applications like Samsung Pay (Komaromy, 2018).
+A.2.2. Cache-based Attacks
+Given that the cache is shared between NW and SW in TrustZone-enabled processors, several studies capitalize on the contention arising from this cache coherency to extract sensitive information (Zhang et al., 2016a; Guanciale et al., 2016; Lipp et al., 2016; Zhang et al., 2016c; Cho et al., 2018).
+(Lipp et al., 2016; Zhang et al., 2016c; Cho et al., 2018) use the Prime+Probe method to monitor the cache activities within TrustZone.
+A.2.3. Branch Prediction Attacks
+Similar to the attacks on Intel SGX mentioned earlier, the branch predictor can also be exploited to target TrustZone.
+Since the BTB is shared between NW and SW, techniques like Prime+Probe can be used to disclose secure information to NW.
+For example, Ryan (Ryan, 2019) successfully recovered a 256-bit private key from Qualcomm’s hardware-backed keystore.
+A.2.4. Software-based Fault Injection Attacks
+Several studies have also expanded software-based fault injection attacks to TrustZone or TrustZone-based systems.
+“CLKSCREW” (Tang et al., 2017) leverages manipulation on DVFS to induce faulty computations to breach TrustZone hardware-enforced boundaries, extract secret keys, and bypass code signing operations.
+Rowhammer-based attacks have also been adapted to undermine TrustZone security.
+Given that overclocking typically involves running the processor at high frequencies, which can be relatively easy to detect and prevent (e.g., through hardware frequency locking), “VoltJockey” (Qiu et al., 2020) adopts a different strategy by manipulating voltages rather than frequencies to generate hardware faults on the victim cores.
+Carru (Carru, 2017) use bitflips to bypass security measures and access secure memory, enabling the extraction of RSA keys and undermining the secure storage provided by TrustZone.
+A.2.5. Countermeasures
+In this section, we review and categorize countermeasures based on both previous survey (Cerdeira et al., 2020) and potential solutions from the summarized attack papers.
+In addressing the architectural defences, some research efforts (Sun et al., 2015; Brasser et al., 2019b) enhance isolation granularity between TEE components and reduce the amount of code running in SW, thereby limiting severe privilege escalation attacks within SW.
+Countermeasures (Jang et al., 2015; Jang and Kang, 2018) tackle the weaknesses of inadequate/weak authentication when accessing TEE resources from NW, and potentially insecure shared memory used for data exchange across the NW-SW boundary.
+Unlike Intel SGX, TrustZone lacks built-in on-chip memory encryption, prompting some works (Zhang et al., 2016b; Yun and Zhong, 2019) to design encryption strategies to bridge this gap.
+To further strengthen the verification of both TEE and TA binaries’ integrity and identity, several studies (Ferraiuolo et al., 2017; Lentz et al., 2018) introduced additional primitives, such as enhanced remote attestation and sealed storage.
+To address implementation bugs, “RustZone” (Evenchick, 2018) extends TrustZone by implementing TAs using Rust, which inherently provides memory safety and thread safety.
+Cache side-channel attacks can be mitigated by carefully implementing cryptographic algorithms in software (Guanciale et al., 2016) or using dedicated hardware (Lipp et al., 2016).
+In terms of BTB attacks, flushing the shared microarchitectural structures when transferring between NW and SW can prevent attacks from targeting keys through the BTB.
+However, as observed in the Nexus 5X firmware by (Ryan, 2019), the L1 cache is not flushed, making Prime+Probe-based attacks still feasible.
+Countermeasures against DVFS-based fault injection attacks are similar to those used for SGX.
+In comparison to other attacks on TrustZone, Rowhammer attacks are relatively straightforward to mitigate.
+They rely on inducing bit flips in adjacent memory rows, making them feasible only if non-secure memory is situated next to a memory row within SW.
+If secure memory resides on a separate memory storage device, Rowhammer attacks become impractical.
+A.3. Security of AMD SEV
+A.3.1. Memory Encryption Issues
+Since the release of AMD SEV (AMD, 2023), numerous attacks have targeted its early memory security mechanisms (Hetzelt and Buhren, 2017; Buhren et al., 2017; Du et al., 2017; Morbitzer et al., 2018).
+Hetzelt and Buhren (Hetzelt and Buhren, 2017) utilize a return-oriented programming-like attack to exploit an unencrypted VM control block, enabling arbitrary reading and writing of encrypted memory in the guest VM.
+Such vulnerabilities have been addressed in SEV-ES (Kaplan, 2017).
+Du et al.
+(Du et al., 2017) leverage Electronic Code Book (ECB) mode in memory encryption to conduct a chosen-plaintext attack via an HTTP server installed on the guest VM.
+Sevurity (Wilke et al., 2020) exploits vulnerabilities in the Xor-Encrypt-Xor (XEX) mode to insert arbitrary 2-byte instructions into encrypted memory, while Morbitzer et al.
+(Morbitzer et al., 2018) breach confidentiality by manipulating nested page tables, altering the virtual memory of the guest VMs.
+Most of these attacks arise from the lack of integrity protection, but SEV-SNP (Sev-Snp, 2020) addressed them through the Reverse Map Table.
+Conversely, “CIPHERLEAKS” (Li et al., 2021b) allows a privileged adversary to infer guest VM execution states or recover certain plaintext by monitoring changes in the ciphertext of the victim VM.
+The authors highlight that AMD has confirmed that SEV-SNP is also vulnerable to the CIPHERLEAKS attack (Li, 2022).
+Additionally, “CROSSLINE” and its variants by Li et al.
+(Li et al., 2021a; Li, 2022) extract the memory content of victim VMs by exploiting improper use of the Address Space Identifier (ASID) for controlling VM accesses to encrypted memory pages, cache lines, and TLB entries.
+They also compromise the integrity and confidentiality of SEV VMs by poisoning TLB entries (Li et al., 2021c).
+It is noted that SEV-SNP is expected to address these issues, including TLB misuse.
+A.3.2. Software-based Fault Injection Attacks
+Morbitzer et al.
+(Morbitzer et al., 2021) exploit the system’s lack of memory integrity protection to inject faults and execute arbitrary code within SEV VMs through I/O channels, using the hypervisor to locate and trigger the execution of encrypted payloads.
+However, SEV-SNP (Sev-Snp, 2020) can mitigate this issue through Reverse Map Table (RMP).
+Buhren et al.
+(Buhren et al., 2021) present a voltage fault injection attack enabling the execution of custom payloads on AMD secure processors, subsequently decrypting a VM’s memory by deploying custom SEV firmware.
+“CacheWarp” (Zhang et al., 2024) presents a new software-based fault attack targeting AMD SEV-ES and SEV-SNP, exploiting the possibility to revert modified cache lines of guest VMs to their previous (stale) state at the architectural level.
+This attack only requires interrupting the VM at a point chosen by the attacker to invalidate modified cache lines without writing them back to memory.
+A.3.3. Interrupt-driven Attacks
+Similar to SGX-Step framework (Van Bulck et al., 2017a), SEV-Step (Wilke et al., 2024) introduces reliable single-stepping against SEV VMs.
+In addition to this functionality, SEV-Step facilitates access to common attack techniques, such as page fault tracking and cache attacks targeting SEV.
+More recently, Schlüter et al.
+(Schlüter et al., 2024a) introduce a novel attack where the hypervisor injects well-crafted, malicious $\mathtt{\#VC}$ (a new exception by AMD facilitating communication between the VM and untrusted hypervisor) into a victim VM’s CPU, enabling the attacker to induce arbitrary behaviour in the VM.
+In another related attack, “HECKLER” (Schlüter et al., 2024b) shows that the hypervisor can inject malicious non-timer interrupts to compromise VMs.
+They leverage interrupt handlers with global effects, allowing manipulation of a VM’s register states to alter data and control flow.
+In general, SEV-SNP remains vulnerable to these attacks.
+A.3.4. Countermeasures
+Most memory-related attacks have been addressed in recent AMD product updates.
+Here, we focus on countermeasures against fault injection attacks.
+For malicious voltage drops or glitches, similar to countermeasures used for SGX and TrustZone, hardware-level voltage monitoring circuits and software-based detection mechanisms can be implemented.
+To counter fault attacks that enable fine-grained memory write suppression through specific hypervisor instructions, restricting the use of those instructions at the hardware level or implementing compiler-level solutions can ensure correct memory operations.
+For attacks that inject malicious interrupts, one can block or detect the injections by monitoring external interrupts or disabling the interrupt handler.